### v2

# The Dial-a-Ride Problem with Limited Pickups per Trip

Abstract: The Dial-a-Ride Problem (DARP) is an optimization problem that involves determining optimal routes and schedules for several vehicles to pick up and deliver items at minimum cost. Motivated by real-world carpooling and crowdshipping scenarios, we introduce an additional constraint imposing a maximum number on the number of pickups per trip. This results in the Dial-a-Ride Problem with Limited Pickups per Trip (DARP-LPT). We apply a fragment-based method for DARP-LPT, where a fragment is a partial path. Specifically, we extend two formulations from Rist & Forbes (2021): the Fragment Flow Formulation (FFF) and the Fragment Assignment Formulation (FAF). We establish FFF’s superiority over FAF, both from a theoretical as well as from a computational perspective. Furthermore, our results show that FFF and FAF significantly outperform traditional arc-based formulations in terms of solution quality and time. Additionally, compared to the two existing fragment sets, one with longer partial paths and another with shorter ones, our newly generated fragment sets perform better in terms of solution quality and time when fed into FFF.
Keywords: routing, dial-a-ride problem, limited number of pickups per trip, fragments
1 Background
The Dial-a-Ride Problem (DARP) is an optimization problem that involves finding optimal routes and schedules for a fleet of vehicles to pick up and deliver a set of items (passengers or parcels). Each item has a specific pickup and delivery location and time window, and the objective is to schedule the vehicles in a way that minimizes the total travel time or distance, while satisfying various operational constraints such as vehicle capacity, time windows, and pickup and delivery locations.
DARP has found extensive applications in the sharing mobility (Mourad et al., 2019; Tafreshian et al., 2020) and food delivery sectors. Within the sharing mobility domain, the ride-sharing routing problem represents a variant of DARP. This problem focuses on optimizing routes for drivers, allowing them to efficiently reach their destinations while accommodating passenger pickups and drop-offs, all within the minimum possible time or cost (Dessouky & Mahtab, 2022). Similarly, in the context of online food delivery, the central system needs to make decisions on how to serve submitted requests by efficiently picking up and delivering orders with the goal of minimizing delivery time, which can be seen as an online variant of DARP (Guo et al., 2022).
A frequently disregarded yet vital aspect in this type of routing problems is the limit on the number of pickups per trip. Many studies tend to focus on constraints such as time windows and maximum ride duration, always assuming these can substitute for pickup limits. Nevertheless, explicitly incorporating restrictions on the number of pickups can be crucial for optimizing delivery operations, especially within the shared mobility sector. For instance, in the context of carpooling, each driver typically accommodates a maximum of four passengers per trip, considering vehicle seat capacity and passenger convenience. Service providers such as UberPool [^1] https://www.uberguide.net/how-many-people-can-ride-in-an-uber/, DiDi [^2] https://web.didiglobal.com/au/help-center/how-many-people-are-allowed-on-a-didi-share-ride/, and BlaBlaCar [^3] https://support.blablacar.com/hc/en-gb/articles/360014470360 enforce such restrictions, allowing drivers to pick up no more than two to four passengers on a single trip. Indeed, optimizing a route by initially picking up one passenger and dynamically delivering other passengers, ultimately reaching the first passenger’s destination, is technically feasible. However, private mobility companies often avoid this strategy due to considerations related to passenger comfort. Another influential factor in this consideration is driver preference. In food delivery platforms like Meituan [^4] https://finance.sina.com.cn/tech/2022-05-10/doc-imcwipii9054944.shtml, drivers possess the autonomy to determine the highest number of orders they undertake at one time. Similarly, in the crowdshipping domain, considering drivers’ preferences is crucial as they have the right to set their preferred maximum number for parcel pickups (Arslan et al., 2019). In particular, a company should not anticipate an occasional driver, who might only intend to deliver a few parcels on a sporadic basis or is simply looking to earn some extra income, to handle a high volume of parcels like a full-time staff driver. A similar assumption is found in the concept of limited delivery stops as discussed in Simoni et al. (2020) and Fessler et al. (2022). In this paper, a “customer” can entail more than one person and/or item, but gives rise to only one pickup and two stops (one for pickup and one for delivery).
In summary, although most previous research has ignored explicit limitations on the number of pickups per trip, in reality such limits have a significant impact on improving customer satisfaction and accommodating driver’s preference. To address such a carpooling scenario where a fleet of vehicles needs to pick up and deliver a group of passengers with the goal to minimize cost, and each vehicle has its unique origin and destination and adheres to a maximum customer pickup limit, this paper introduces an innovative problem termed the Multi-depot Dial-a-Ride Problem with Limited Pickups per Trip (MDARP-LPT). We also introduce a variant called the Dial-a-Ride Problem with Limited Pickups per Trip (DARP-LPT), where all vehicles start and end their routes at the same depot. DARP-LPT is applicable in crowdshipping scenarios, where several occasional riders are used to handle pickup and delivery orders, such as food or parcel deliveries. These riders (e.g. students or part-time riders) do not have specific origins and destinations for travel; their goal is to complete a limited number of orders as a part-time job for additional income. Unlike carpooling, food and parcel deliveries typically occur within a specific physical area, which is why the problem assumes the use of a central depot.
Since MDARP-LPT and DARP-LPT share the same problem structure, we collectively refer to them as (M)DARP-LPT. To efficiently address (M)DARP-LPT, we propose a unified method with two innovative components: fragment-based formulations and novel fragment generation procedures. Firstly, we tailor the fragment-based formulation of DARP proposed by Rist & Forbes (2021) to accommodate (M)DARP-LPT by modifying the decision variables from the fragment-based format to the fragment-vehicle-based format. This change is also necessary for DARP-LPT, where all homogeneous vehicles start and end at the same depot. Secondly, considering the limited number of pickups, we introduce two novel fragment generation procedures. These procedures create improved fragment sets, enhancing computational efficiency. Traditionally, there are two fragment sets: one with longer partial paths and the other with shorter ones. Our two proposed sets diverge: the first refines the fragment set with longer partial paths through path enumeration, while the second blends longer and shorter partial paths from both traditional fragment sets.
From a practical perspective, this research addresses a commonly overlooked aspect in the dial-a-ride problem: the limited number of pickups per trip. Our developed method offers more tailored and efficient solutions for the problem, thereby contributing to the improvement of carpooling and crowdshipping services. From an academic perspective, this study compares and evaluates various fragment sets for the fragment-based formulation in routing problems, providing essential insights for tackling similar pickup and delivery challenges. Furthermore, our innovative fragment generation method accelerates computations through enumeration; this logic can offer methodological insights for addressing similar routing problems.
The remainder of this article is organized as follows: Section 2 provides a comprehensive literature review. Following that, Section 3 formally defines (M)DARP-LPT, presenting an arc-based formulation and two fragment-based formulations. Section 4 introduces traditional and two novel methods for fragment generation. Then, in Section 5, we analyze the results of our computational experiments. Finally, we draw conclusions in Section 6.
2 Literature review
Since (M)DARP-LPT is a variant of DARP, we investigate existing research methodologies on related optimization problems, such as the Pickup and Delivery Problem (PDP) and the Dial-a-Ride Problem. PDP involves transporting a set of items from their pickup locations to their respective delivery locations (where multiple items from a single pickup node may be destined for different delivery nodes), with the objective of finding optimal routes while minimizing transportation costs (Savelsbergh & Sol, 1995). DARP is a special case of PDP where each location corresponds to a single item, and every item is associated with a unique pickup and delivery location pair.
Traditional methods, like branch and cut (Cordeau, 2006) and branch and price (Ropke & Cordeau, 2009; Baldacci et al., 2011), have been utilized for PDP and DARP, but recent studies suggest that fragment-based methods exhibit better performance. Here, a fragment is a partial path where only the start and end node have an empty load.
A notable example of the fragment-based approach is the work of Alyasiry et al. (2019), who apply a fragment-based strategy to address the Pickup and Delivery Problem with Time Windows, following the Last-In-First-Out loading strategy. They employ the symbol sequence ``$ppppdddd"$ to denote a full fragment, with “$p$” representing a pickup node and “$d$” indicating a delivery node (this representation is also adopted below). This notation implies that, within the fragment, the last pickup must precede the first delivery. Alyasiry et al. (2019) enumerate all possible paths by combining fragments, thus leading to a fragment-based formulation on the time-space network.
The fragments’ applicability extends to various related problems, retaining the characteristic of empty-loaded start and end nodes but not limited to the specific “$ppppdddd$” structure. Su et al. (2023) effectively compute the minimum excess user ride time in an electric autonomous DARP using the fragment-based method. Additionally, Zhang et al. (2023) tackle a vehicle-customer coordination problem that shares similarities with DARP but incorporates the flexibility for passengers to move in order to meet their drivers faster. To tackle this problem, they design an algorithm based on the fragment-based time-space network. Their procedure not only achieves effective results, but also demonstrates its value for online DARP.
Building on Alyasiry et al. (2019), Rist & Forbes (2021) propose the use of restricted fragments to address the dial-a-ride problem. Since the full fragments in this problem may have the form “$ppdpdpdd$” (as an illustration that the last pickup does not precede the first delivery, this does not imply a universal pattern), their number would be quite high. Nevertheless, Rist & Forbes find that a full fragment with the form “$ppdpdpdd$” can be decomposed into the restricted fragments “$ppd$” +“$pd$”+ “$pdd$”, which start at a pickup node whose predecessor is a delivery node (or depot) and ends at a delivery node whose successor is a pickup node (or depot). Obtaining these restricted fragments involves initially deriving them from fragments conforming to the form “$ppppdddd$” (where the last pickup must precede the first delivery) and then applying a dominance rule to minimize their count. By incorporating temporal information into these restricted fragments, they enable a formulation based on a physical network instead of the time-space network.
While recent literature has made significant contributions to fragment-based methods, other methods may be further improved for (M)DARP-LPT. In particular, alternative methods for fragment generation may result in a higher-quality set of fragments, further enhancing computational performance. Moreover, to the best of our knowledge, no study has conducted a comparison of various fragment types or explored the possibility of combining different fragment categories.
3 Problem statement and linear formulations
Both MDARP-LPT and DARP-LPT are extensions of DARP, and their major difference from DARP lies in the limitations on the number of pickups per trip. In this section, we begin by providing a formal problem description of DARP-LPT and a NP-hardness proof in Section 3.1. Subsequently, we present several linear formulations for DARP-LPT: an arc-based formulation in Section 3.2 and two fragment-based formulations in Section 3.3. The arc-based formulation closely resembles the one outlined in Cordeau (2006), except for the constraints on the limited number of pickups. The two fragment-based formulations expand on the framework of Rist & Forbes (2021), transitioning from a formulation using fragment-based decision variables to one employing fragment-vehicle-based variables. This adaptation is tailored to our specific requirements for limited number of pickups. Finally, in Section 3.4, we introduce MDARP-LPT, illustrating the differences in its definition and formulations as compared to DARP-LPT.
3.1 Problem description
DARP-LPT involves $n$ customers, with each customer $i$ having a unique pickup and delivery location pair represented as $(i,i+n)$, where $n$ is the total number of customers. The pickup and delivery locations can be represented by the sets $P=\{1,2,\dots,n\}$ and $D=\{n+1,n+2,\dots,2n\}$, respectively. The location set $N=P\cup D\cup\{0,2n+1\}$ consists of all pickup and delivery locations, along with the origin depot location 00 and the destination depot location $12n+$ 1. Each location $i\in P\cup D$ has a specific time window $(e_{i},l_{i})$, defining the earliest and latest time a vehicle can arrive at that location. The arc set is denoted by $A$, including all arcs $(i,j)\in N\times N$ with $i\neq j$, and the cost and time associated with traversing the arc $(i,j)\in A$ are represented by $C_{ij}$ and $T_{ij}$, respectively. The set of vehicles is represented by $V$. The load quantity change at location $i\in N$ is expressed as $q_{i}$, with $q_{0}=q_{2n+1}=0$ and $q_{i}=-q_{n+i}$ ($q_{i}\in\mathbb{N}^{+}$) for $i\in P$. This quantity represents the customer size, such as the size of an order or the number of passengers. At pickup location $i\in P$, the vehicle adds load $q_{i}$, while at delivery location $i+n\in D$, the vehicle reduces load by $q_{i}$.
The objective of DARP-LPT is to minimize travel costs while ensuring that at most $|V|$ vehicles are utilized to visit all pickup and delivery locations, starting from the origin depot location 00 and ending at the destination depot location $12n+$ 1. To achieve this, a constraint is imposed where each pickup location $i\in P$ must be visited by the same vehicle as its corresponding delivery location $i+n\in D$ (pairing constraints), with the pickup location visited before the delivery location (precedence constraints). Moreover, each customer $i\in P$ has a maximum ride time $R_{i}$, each location $i\in N$ has its time window, and each vehicle is limited to a capacity of at most $Q$ with its maximum ride time $R_{0}$.
Up to this point, all the statements adhere to the standard DARP framework. However, an important and unique feature in DARP-LPT is that each vehicle is allowed to have at most $L$ number of pickups. Notably, $Q$ and $L$ are distinct and represent different constraints. $Q$ restricts the maximum size of concurrent customer loads within the vehicle, and the concurrent load can both increase and decrease throughout the trip. In contrast, $L$ limits the total number of pickups, which can only increase monotonically.
DARP-LPT is NP-hard. This can be proven by a reduction from the Capacitated Vehicle Routing Problem (CVRP). CVRP involves arranging a route plan using a fleet of vehicles to visit a set of customers at minimum cost. Each vehicle has a capacity denoted by $C$ (Cornuejols & Harche, 1993; Achuthan et al., 2003). To perform the reduction, in a DARP-LPT instance, we set the maximum number of pickups per trip $L$ equal to the vehicle capacity $C$ and merge each customer’s pickup and delivery nodes into a single node. This creates a problem instance equivalent to CVRP, where each merged node represents a customer with unit size and each vehicle has the capacity $C$ equal to $L$. Since CVRP is known to be NP-hard when $C\geq 3$ (Chen, 2023), this reduction demonstrates that DARP-LPT is also NP-hard. MDARP-LPT can also be proven to be NP-hard using a reduction similar to DARP-LPT.
3.2 Arc-based formulation of DARP-LPT
The arc-based formulation (ABF), as presented below, is the same as that of Cordeau (2006), except for the constraints (13).
The decision variables are as follows:
Table 1: Decision variables of arc-based formulation
Decision variables Definition
$f_{ijv}$ = 1, if the arc $(i,j)\in A$ is traversed by vehicle $v\in V$; = 0, otherwise
$t_{iv}$ the departure time of vehicle $v$ at location $i\in N$
$Q_{iv}$ the load of vehicle $v$ after visiting $i\in N$
$\min\sum_{v\in V}\sum_{(i,j)\in A}C_{ij}f_{ijv}$ (1)
$s.t.$ $\sum_{v\in V}\sum_{j\in N}f_{ijv}$ $=1$ $i\in P$ (2)
$\sum_{j\in N}f_{ijv}$ $=\sum_{j\in N}f_{(i+n)jv}$ $i\in P,v\in V$ (3)
$\sum_{j\in N}f_{0jv}$ $=1$ $v\in V$ (4)
$\sum_{j\in N}f_{jiv}$ $=\sum_{j\in N}f_{ijv}$ $i\in P\cup D,v\in V$ (5)
$\sum_{i\in N}f_{i(2n+1)v}$ $=1$ $v\in V$ (6)
$t_{iv}+T_{ij}$ $\leq t_{jv}+M_{ij}(1-f_{ijv})$ $(i,j)\in A,v\in V$ (7)
$e_{i}\leq t_{iv}$ $\leq l_{i}$ $i\in N,v\in V$ (8)
$Q_{iv}+q_{j}$ $\leq Q_{jv}+W_{ij}(1-f_{ijv})$ $(i,j)\in A,v\in V$ (9)
$\max(0,q_{i})\leq Q_{iv}$ $\leq\min(Q,Q+q_{i})$ $i\in N,v\in V$ (10)
$t_{(n+i)v}-t_{iv}$ $\leq R_{i}$ $i\in P,v\in V$ (11)
$t_{(2n+1)v}-t_{0v}$ $\leq R_{0}$ $v\in V$ (12)
$\sum_{(i,j)\in A}f_{ijv}$ $\leq 2L+1$ $v\in V$ (13)
$f_{ijv}$$\in\{0,1\}$ $(i,j)\in A,v\in V$ (14)
$t_{iv}$ $\geq 0$ $i\in N,v\in V$ (15)
$Q_{iv}$ $\in\mathbb{N}$ $i\in N,v\in V$ (16)
The objective function (1) aims to minimize the total cost of traversing arcs. Constraints (2) and (3) ensure that each pickup location is visited exactly once and that the paired locations are traversed by the same vehicle. Constraints (4) to (6) specify the network flow requirements. Constraints (7) to (10) ensure time and load consistency. Here, $M_{ij}=\max(0,l_{i}+T_{ij}-e_{j})$ and $W_{ij}=\min(Q,Q+q_{i})$. Constraints (11) and constraints (12) set the maximum ride time for customers and vehicles. Constraints (13) restrict the maximum number of pickup locations per trip. If the number of pickups on a route is at most $L$, then the route has at most $2L$ pickup and delivery locations and $12L+$ 1 arcs. Constraints (14) to (16) specify the domain of the variables $f$, $t$, and $Q$.
In order to reduce part of the symmetry that is inherent in the formulation, we define $V_{f}=\{1,2,\ldots,\lceil n/L\rceil\}$ as the fixed set of vehicles that must be used. This definition arises from the observation that, if a vehicle can deliver at most $L$ customers, a minimum of $\lceil n/L\rceil$ vehicles are needed. Therefore, we add $f_{0(2n+1)v}=0$ for each fixed vehicle $v\in V_{f}$.
3.3 Fragment-based formulations of DARP-LPT
The framework presented by Rist & Forbes (2021) defines a fragment-based formulation for DARP, which necessitates the inclusion of callbacks to handle rarely violated constraints related to subtour elimination, time windows, and maximum ride time. In the case of DARP, even when multiple vehicles are used, this fragment-based format remains feasible. In this section, our fragment-based formulations build upon the work of Rist & Forbes (2021). However, since DARP-LPT involves frequently violated constraints, in particular the limitations on the number of pickups per trip (per vehicle), we introduce different decision variables, shifting from a fragment-based format to a fragment-vehicle-based format. Consequently, we propose two distinct fragment-based formulations tailored for multi-vehicle settings: a fragment flow formulation and a fragment assignment formulation. The former assigns a vehicle to each fragment and node arc, while the latter exclusively links a vehicle with each fragment. In particular, the latter formulation draws inspiration from the client-depot assignment formulation in multi-depot routing problems (Bektaş et al., 2020). In our context, the “client” corresponds to the fragment, while the “depot” corresponds to the vehicle.
3.3.1 Preliminaries and notation
In this subsection, we largely follow the definitions of Rist & Forbes (2021).
A DARP route is a route $(0,i_{1},i_{2},\ldots,i_{K},2n+1)$ for which at least one feasible schedule exists, which respects the pairing, precedence, capacity, and time window constraints of DARP. The sequence of locations $(i_{1},i_{2},\ldots,i_{K})$, excluding depot locations, is called a DARP route path. Any segment of this sequence is referred to as a route path.
Rist & Forbes (2021) introduce the concept of restricted fragments for DARP. These fragments can be combined to enumerate all DARP routes, and a network is constructed to include all these restricted fragments. Prior to introducing restricted fragments, we describe a full fragment in more detail.
A full fragment (FF) refers to a partial DARP route path where only the start and end nodes have an empty load. For instance, a valid fragment could have a sequence of locations $(p1,p2,d2,d1)$ or $(p1,p2,d1,p3,d2,d3)$, where $pj$ and $dj$ respectively denote the pickup and delivery locations for customer $j$. An invalid case is $(p1,d1,p2,p3,d2,d3)$ as this violates the empty load condition after location $d1$ and should be decomposed into $(p1,d1)$ and $(p2,p3,d2,d3)$.
A restricted fragment (RF) is a part of a full fragment, starting at a pickup node whose predecessor is a delivery node or depot and ending at a delivery node whose successor is a pickup node or depot. An RF contains at least one pickup-to-delivery movement (not necessarily corresponding to a customer’s pickup and delivery location pair) and cannot be divided into smaller RFs (meaning each RF is as small as possible), e.g., $(p1,p2,d1)$, $(p1,d2,d1)$, $(p1,p2,p3,d4,d5)$, $(p5,d8)$, $(p5,p7,d1,d2,d3)$. The following two examples are not RFs: $(p1,p2,p3,p4)$ and $(p2,d1,p3,d3,d2)$, where the former lacks a pickup-to-delivery movement and the latter can be split after $d1$. A special type of RF is a FF with the format “$ppppdddd$”, where all pickups precede deliveries, and which cannot be divided into smaller RFs, for instance $(p1,p2,d1,d2)$ or $(p1,p3,p2,d2,d1,d3)$. Below are two further examples, where the left-hand side of the equality is the route path of an FF, and the right-hand side consists of RFs. This decomposition is entirely unique.
$(p2,p1,d2,p4,d4,p3,d1,d3)=(p2,p1,d2)+(p4,d4)+(p3,d1,d3)$;
$(p2,p5,d2,p4,d4,p3,d5,d3)=(p2,p5,d2)+(p4,d4)+(p3,d5,d3)$.
When combining RFs into routes, it will be crucial to track loads on board. In the first example, the load after visiting locations $d2$ and $d4$ is ${l_{1}}$, while in the second example, the load after visiting these locations is ${l_{5}}$, where $l_{j}$ denotes the load for customer $j$. This difference highlights that both the route path and the load information are essential for accurately representing an RF. A more precise definition for fragments (RF and FF) will be provided after the introduction of the following concepts.
A node is a pair consisting of a location and a set of loads (with customer identification and not just the number of customers) on board, denoted as $(loc(i),loadset(i))$ (or just $loc(i)$ if $loadset(i)=\emptyset$). The $loadset(i)$ includes the unserved customers along a path after departing from location $loc(i)$ but excludes those from $loc(i)$. For instance, the nodes from the route path $(p1,p2,d1,p3,d2,p4,d3,d4)$ are expressed as $p1$, $(p2,\{l_{1}\})$, $(d1,\{l_{2}\})$, $(p3,\{l_{2}\})$, $(d2,\{l_{3}\})$, $(p4,\{l_{3}\})$, $(d3,\{l_{4}\})$, and $d4$. Notably, nodes with the same location can be distinct, as illustrated by examples such as $(p2,\{l_{1}\})$, $(p2,\{l_{3}\})$, $(p2,\{l_{2},l_{3}\})$, and $p2$. We define a node with a pickup location as a “pickup node” and a node with a delivery location as a “delivery node”. Specifically, we denote the sets of pickup and delivery nodes as $P_{N}$ and $D_{N}$, respectively, distinguishing them from the sets of pickup and delivery locations denoted as $P$ and $D$.
A fragment (FF and RF) is defined as a partial DARP route path with distinct start and end nodes. The route path records the sequence of locations traversed by a fragment, while the start and end nodes include information regarding the loads onboard. For a FF, merely the route path suffices for reference, while for an RF, alongside the route path, inclusion of both the start and end nodes is necessary for identification. For illustrative purposes, we present several example fragments: $p4-(p4,d4)-d4$, $(p4,\{l_{1}\})-(p4,d4)-(d4,\{l_{1}\})$, $(p4,\{l_{2}\})-(p4,d4)-(d4,\{l_{2}\})$, $(p4,\{l_{1},l_{2}\})-(p4,d4)-(d4,\{l_{1},l_{2}\})$. In this context and below, for such a format, the first and last element represent the start and end nodes of the fragment respectively, while the part between two hyphens denotes the route path. Despite following the same route path, these fragments are not the same.
3.3.2 Fragment-based network
Building upon the introductory concepts, we now introduce the fragment-based network. To facilitate comprehension, we present several figures (Fig. 1, Fig. 2, and Fig. 3). In these figures, ellipses represent nodes, solid lines (with ellipses on both sides) indicate fragments (with their route paths written above or to the right), and dashed lines with arrows are node arcs (further explained below). $O^{+}$ and $O^{-}$ denote the origin depot and destination depot, respectively.
[…]
Figure 1: Several feasible routes for DARP and their decomposition into fragments (RF and FF)
Before explaining the fragment-based network, we will introduce the connections between fragments and the connections between nodes (referred to later as node arcs), both of which can be seen as the dashed lines in Fig. 1. In the former category, the connections between FFs are straightforward. For example: $p1-(p1,d1)-d1$ can connect to $p2-(p2,p3,d2,d3)-d3$ to form a route path $(p1,d1,p2,p3,d2,d3)$. However, the connections between RFs require ensuring consistency between the last node’s load of the preceding fragment and the first node’s load of the succeeding one. Take the following three RFs as an example.
$p2-(p2,p1,d2)-(d2,\{l_{1}\})$,
$(p4,\{l_{1}\})-(p4,d1)-(d1,\{l_{4}\})$,
$(p3,\{l_{4}\})-(p3,d4,d3)-d3$.
These can be combined to a route path $(p2,p1,d2,p4,d1,p3,d4,d3)$. However, the first RF cannot connect to the third one since the load of the end node of the first RF is not the same as that of the start node of the third one.
The connection between two fragments can be indicated by the connection between the preceding fragment’s end node and the succeeding fragment’s start node. For instance, in the third row of Fig. 1, the connection between $p1-(p1,d1)-d1$ and $p2-(p2,d2)-d2$ can be regarded as the connection between $d1$ and $p2$. Furthermore, the connections between distinct fragments sharing identical end nodes $h^{-}$ and different fragments with identical start nodes $h^{+}$, can be consistently depicted by the connection $(h^{-},h^{+})$. For instance, in the third and fourth row of Fig. 1, the connections between $p1-(p1,d1)-d1$ and $p2-(p2,d2)-d2$, as well as between $p1-(p1,d1)-d1$ and $p2-(p2,p3,d2,d3)-d3$, can both be represented by the connection $(d1,p2)$.
Therefore, we introduce the concept of node arc to describe such a connection between nodes. In the node arc $(h^{-},h^{+})$, $h^{-}$ and $h^{+}$ then represent the start and end nodes of the node arc, respectively. This node arc concept differs from the arc concept defined in Section 3.2, as node arcs connect one node to another, with each node encompassing both location and load information (unless the node has empty loads). An additional note is that a node arc can only represent a connection from a delivery node or the origin depot to a pickup node or from a delivery node to the destination depot. This is because all fragments’ end nodes must be delivery nodes, while their start nodes must be pickup nodes. Crucially, the connected nodes must share the same load.
Decomposing feasible DARP routes into fragments yields a significantly smaller number of fragments compared to DARP routes. Conversely, if we combine these fragments, we can obtain all feasible DARP routes, albeit with the possibility of some redundant routes. Fig. 2 illustrates how feasible DARP routes can be gathered in a transitional network (as an intermediate step towards the fragment-based network). Fig. 1 presents several DARP routes and their corresponding fragments. The network in Fig. 2 will use these fragments and allow to enumerate all the DARP routes shown in Fig. 1. Identical fragments across different routes in Fig. 1 can be consolidated into a single fragment while maintaining the connections to other fragments. Subsequently, the fragment-based network of Fig. 3 is derived from the network of Fig. 2. When different fragments share the same start and/or end nodes, then those nodes can also be merged. For instance, two fragments with the same start and end nodes but different route paths $(p1,d1)$ and $(p1,p4,d4,d1)$ become two route paths originating from node $p1$ and terminating at node $d1$. Such consolidations help reduce connections (node arcs). Node arcs with the same start node and end node can also be merged; the three connections from $(p1,d1)$ to $(p2,d2)$, from $(p1,d1)$ to $(p2,p3,d2,d3)$, and from $(p1,p4,d4,d1)$ to $(p2,d2)$ in Fig. 2, for instance, can be merged into one node arc $(d1,p2)$ in Fig. 3.
$O^{+}$$p3$($d2$,{$l_{3}$})($p1$,{$l_{3}$})$d3$$O^{-}$($p3$,$p2$,$d2$)($p1$,$d1$,$d3$)$p1$$d1$$p2$$d2$$p3$$d3$($p1$,$d1$)($p2$,$d2$)($p3$,$d3$)$p1$$d1$$p2$$d3$($p1$,$p4$,$d4$,$d1$)($p2$,$p3$,$d2$,$d3$)
Figure 2: A transitional network for DARP by merging all identical fragments from Fig. 1
[…]
Figure 3: Fragment-based network for DARP by merging all identical nodes from Fig. 2
The practical process of network generation proceeds as follows. The fragment-based network comprises nodes, fragments, and node arcs. The initial step involves generating fragments (outlined in Section 4). Following this, the start and end nodes of all fragments are identified and added to the network. Subsequently, all fragments are integrated into the network, connecting from their start nodes to end nodes. Different fragments can share the same start and/or end nodes. In the final phase, for the node arcs, all delivery nodes and the origin depot connect with all pickup nodes, and all delivery nodes connect with the destination depot. If none of the fragments ending at $h^{-}$ can connect to any fragment starting at $h^{+}$ due to time window or any other constraints, the node arc $(h^{-},h^{+})$ can be eliminated. When the network exclusively includes FFs, the number of node arcs (excluding those involving depots) is at most $|n|\times|n|$. For the network with RFs, the potential number of node arcs could be higher due to the possibility of different nodes with the same location having different loads. Moreover, a feasible route typically starts with a node arc from the depot, followed by a fragment leading to a delivery node. Subsequently, there must be a node arc from the current delivery node to another pickup node or the destination depot. After that, if the current node is not the depot yet, it must be succeeded by another fragment and then a node arc again, continuing in this pattern. Therefore, any feasible route in Fig. 3 must adhere to the alternating pattern of dashed lines and solid lines.
We now formalize the foregoing concepts. The set of nodes is expressed as: $N_{N}=P_{N}\cup D_{N}\cup\{O^{+},O^{-}\}$, where the origin depot $O^{+}$ and destination depot $O^{-}$ correspond to location 00 and location $12n+$ 1, respectively. The set of fragments, denoted as $F$, is derived from customer locations, as detailed in Section 4. $F$ may include both FFs and RFs or exclusively one type. We use the notation $F_{h}^{+}$ for the sets of fragments starting at pickup node $h\in P_{N}$, $F_{h}^{-}$ for those ending at delivery node $h\in D_{N}$, and $F_{i}$ for the set of fragments traversing location $i\in P\cup D$. Additionally, $A_{N}$, $A^{+}_{n}$, and $A^{-}_{n}$ represent the set of node arcs, the set of node arcs starting from delivery node (or the origin depot) $h\in D_{N}\cup\{O^{+}\}$, and the set of node arcs ending at pickup nodes (or the destination depot) $h\in P_{N}\cup\{O^{-}\}$, respectively. Notably, $A_{N}$ exclusively includes node arcs originating from delivery nodes (or the origin depot) and terminating at pickup nodes (or the destination depot).
For each fragment $f\in F$, we designate the number of included pickup nodes as $L_{f}$, and its cost $c_{f}$ is calculated as the sum of travel costs over all locations visited by the fragment, i.e., $c_{f}=\sum_{(i,j)\in f}C_{loc(i)loc(j)}$, where $loc(i)$ represents the location of node $i\in N_{N}$. Furthermore, the cost of node arc $a$, $c_{a}$, is determined as follows: $c_{a}=C_{loc(i)loc(j)}$, $a=(i,j)$, $a\in A_{N}$.
3.3.3 Fragment flow formulation
The first fragment-based formulation is the fragment flow formulation (FFF), which expands on the framework of Rist & Forbes (2021). It transitions from a formulation using fragment-based and node arc-based decision variables to one employing fragment-vehicle-based and node arc-vehicle-based variables. Please note that the current formulation below is not complete. Lazy constraints, specifically designed for the elimination of infeasible paths, are applied to address subtour elimination, time windows, and maximum ride time. The details are described in Section 3.3.6.
Table 2: Decision variables of FFF
Decision variables Definition
$x_{fv}$ = 1, if fragment $f\in F$ is traversed by vehicle $v\in V$; = 0, otherwise
$y_{av}$ = 1, if node arc $a\in A_{N}$ is traversed by vehicle $v\in V$; = 0, otherwise
$\min\sum_{v\in V}\sum_{f\in F}c_{f}x_{fv}+\sum_{v\in V}\sum_{a\in A_{N}}c_{a}y_{av}$ (17)
$s.t.$ $\sum_{f\in F_{h}^{-}}x_{fv}$ $=\sum_{a\in A_{h}^{+}}y_{av}$ $h\in P_{N},v\in V$ (18)
$\sum_{f\in F_{h}^{+}}x_{fv}$ $=\sum_{a\in A_{h}^{-}}y_{av}$ $h\in D_{N},v\in V$ (19)
$\sum_{v\in V}\sum_{f\in F_{i}}x_{fv}$ $=1$ $i\in P$ (20)
$\sum_{f\in F}L_{f}x_{fv}$ $\leq L$ $v\in V$ (21)
$x_{fv},y_{av}$$\in\{0,1\}$ $f{\in}F,a{\in}A_{N},v{\in}V$ (22)
The objective function (17) minimizes the weighted cost of used fragments and node arcs. Constraints (18) and (19) describe the network flow for fragments and node arcs. For each pickup node at which a fragment ends, there must be a node arc starting from it. Similarly, for each delivery node at which a fragment starts, a node arc must end at that node. These relationships are applicable to each vehicle. Constraints (20) guarantee that each pickup location is visited exactly once. Constraints (21) set the maximal number of pickups. Constraints (22) specify the domains of $x$ and $y$.
The additional constraints (23) below specify whether a vehicle is used or unused. They are not necessary but can assist the solver in branching, and thus, improve computational performance. Here, we define $w_{v}$ as 1 if vehicle $v\in V$ is utilized, and 0 otherwise.
$\sum_{a\in A_{O^{+}}^{+}}y_{av}$ $=w_{v}$ $v\in V$ (23)
3.3.4 Fragment assignment formulation
Table 3: Decision variables of FAF
Decision variables Definition
$X_{f}$ = 1, if fragment $f\in F$ is traversed; = 0, otherwise
$Y_{a}$ = 1, if node arc $a\in A_{N}$ is traversed; = 0, otherwise
$x_{fv}$ = 1, if fragment $f\in F$ is traversed by vehicle $v\in V$; = 0, otherwise
An alternative formulation involves the assignment of fragments to vehicles and is called fragment assignment formulation (FAF). FAF extends the framework of Rist & Forbes (2021) by transitioning from a formulation using fragment-based decision variables to one that incorporates fragment-vehicle-based variables. Unlike FFF, however, which introduces node arc-vehicle-based decision variables, FAF keeps the node arc-based variables by means of a separate set of constraints (which hinges upon constraint set (31)). This model is inspired by the client-depot assignment formulation used in multi-depot routing problems (Bektaş et al., 2020). In the same way as for FFF, the current formulation below is not complete. Lazy constraints to address subtour elimination, time windows, and maximum ride time are discussed in Section 3.3.6.
$\min$ $\sum_{f\in F}c_{f}X_{f}+\sum_{a\in A_{N}}c_{a}Y_{a}$ (24)
$s.t.$ $\sum_{f\in F_{h}^{-}}X_{f}$ $=\sum_{a\in A_{h}^{+}}Y_{a}$ $h\in P_{N}$ (25)
$\sum_{f\in F_{h}^{+}}X_{f}$ $=\sum_{a\in A_{h}^{-}}Y_{a}$ $h\in D_{N}$ (26)
$\sum_{f\in F_{i}}X_{f}$ $=1$ $i\in P$ (27)
$\sum_{a\in A_{O^{+}}^{+}}Y_{a}$ $\leq|V|$ (28)
$x_{fv}$ $\leq X_{f}$ $f\in F,v\in V$ (29)
$\sum_{v\in V}x_{fv}$ $=X_{f}$ $f\in F$ (30)
$\sum_{f\in F^{-}_{i}}x_{fv}+Y_{a}$ $\leq\sum_{f\in F^{+}_{j}}x_{fv}+1$ $a=(i,j)\in A_{N}:i\neq O^{+},j\neq O^{-},v\in V$ (31)
$\sum_{f\in F}L_{f}x_{fv}$ $\leq L$ $v\in V$ (32)
$X_{f},Y_{a}$$\in\{0,1\}$ $f{\in}F,a{\in}A_{N}$ (33)
$x_{fv}$$\in\{0,1\}$ $f{\in}F,v{\in}V$ (34)
The objective function (24) minimizes the weighted cost of used fragments and node arcs. Constraints (25) and (26) describe the network flow for fragments and node arcs. Constraints (27) guarantee that each pickup location is visited exactly once. Constraints (28) state that at most $|V|$ vehicles are used. Constraints (29)-(31) assign the fragments to vehicles. Specifically, constraints (29) impose that if a fragment is traversed by a vehicle, the fragment is used; constraints (30) state that if a fragment is chosen, it must be assigned to a vehicle; and constraints (31) express that if a node arc is used, the fragments preceding the arc and the fragments succeeding the arc must be traversed by the same vehicle. This explains why it does not matter which vehicle traverses the node arc. Constraints (32) set the maximal number of pickups. Constraints (33) and constraints (34) specify the domains of $X$, $Y$, and $x$.
Although the variables $X$ and $x$ in this formulation can be expressed solely in terms of $x$, keeping both $X$ and $x$ yields better computational results.
3.3.5 Comparison of FFF and FAF
In this subsection, we will establish that FFF is a stronger formulation than FAF. We will show this in two steps. Firstly, any feasible solution to the linear relaxation of FFF also qualifies as a feasible solution to the linear relaxation of FAF. Secondly, we will provide an example solution from the polytope of FAF that does not appear in the polytope of FFF.
Regarding the first step, it is easy to observe that constraints (18)-(20) and constraints (21) in FFF can imply constraints (25)-(27) and constraints (32) in FAF. Furthermore, constraints (23) in FFF can indicate constraints (28) in FAF. Therefore, our only task is to demonstrate how a solution ($x^{\prime}$, $y^{\prime}$) that is feasible under FFF constraints (18) and (19) can meet the requirements of FAF constraints (31). Let ($x^{\prime}$, $y^{\prime}$) represent a solution from the linear relaxation of FFF. We introduce two auxiliary variables: $X^{\prime}$ and $Y^{\prime}$. Here, $X^{\prime}_{f}=\sum_{v\in V}x_{fv}$ for all $f\in F$ and $v\in V$; and $Y^{\prime}_{f}=\sum_{v\in V}y^{\prime}_{av}$ for all $a\in A_{N}$ and $v\in V$.
Proposition 1.
Constraints (18) and constraints (19) in FFF imply constraints (31) in FAF.
Proof.
Consider an arc $a=(i,j)$. Our objective is to demonstrate that any solution $(x^{\prime},y^{\prime},X^{\prime},Y^{\prime})$ satisfying the FFF constraints will also be feasible for the FAF network flow constraints (31).
According to the FFF network flow constraints (18) and (19), we have:
$\sum_{f\in F_{i}^{-}}x^{\prime}_{fv}=\sum_{a\in A_{i}^{+}}y^{\prime}_{av},\quad\text{and}\quad\sum_{f\in F_{j}^{+}}x^{\prime}_{fv}=\sum_{a\in A_{j}^{-}}y^{\prime}_{av}$
Thus, we can transform the FAF constraints (31) as follows:
$\sum_{f\in F^{-}_{i}}x^{\prime}_{fv}+Y^{\prime}_{a}\leq\sum_{f\in F^{+}_{j}}x^{\prime}_{fv}+1\iff\sum_{a^{\prime}\in A_{i}^{+}}y^{\prime}_{a^{\prime}v}+Y^{\prime}_{a}\leq\sum_{a^{\prime}\in A_{j}^{-}}y^{\prime}_{a^{\prime}v}+1\iff\sum_{a^{\prime}\in A_{i}^{+}\setminus a}y^{\prime}_{a^{\prime}v}+Y^{\prime}_{a}\leq\sum_{a^{\prime}\in A_{j}^{-}\setminus a}y^{\prime}_{a^{\prime}v}+1$
We now examine the left side of the last inequality, $\sum_{a^{\prime}\in A_{i}^{+}\setminus a}y^{\prime}_{a^{\prime}v}+Y^{\prime}_{a}$, using FFF constraints (18). Then, we obtain:
$\sum_{a^{\prime}\in A_{i}^{+}\setminus a}y^{\prime}_{a^{\prime}v}+Y^{\prime}_{a}\leq\sum_{a\in A_{i}^{+}}Y_{a}=\sum_{f\in F^{-}_{i}}X_{f}\leq 1\leq\sum_{a^{\prime}\in A_{j}^{-}\setminus a}y^{\prime}_{a^{\prime}v}+1$
Thus, we see that constraints (31) are automatically fulfilled if constraints (18) and (19) hold. ∎
The second step is to show that the polyhedra of the two formulations do not coincide. To illustrate this point, we provide an example solution of FAF that is cut away by FFF. For example: if $V=\{v1,v2\}$, $F^{-}_{i}=\{f\}$, $F^{+}_{j}=\{g\}$, and $a=(i,j)$, constraints (31) of FAF become: $x_{fv1}+x_{fv2}+Y_{a}\leq x_{gv1}+x_{gv2}+1$. A potential solution under the given constraints is $x_{fv1}=0.5$, $x_{fv2}=0.5$, $Y_{a}=1$, $x_{gv1}=1$, $x_{gv2}=0$. However, this solution does not satisfy constraints (18) and (19) of FFF, where $x_{fv2}=y_{av2}$ and $y_{av2}=x_{gv2}$.
Based on the foregoing, we conclude that FFF dominates FAF. Despite this dominance result, we should still verify empirically how the two formulations behave in terms of runtimes. FAF, for instance, has less decision variables since FAF utilizes arc-based variables $Y$, whereas FFF uses arc-vehicle-based variables $y$.
3.3.6 Inequalities for fragment formulations
The above fragment-based formulations FFF and FAF may not be sufficient to lead to feasible solutions. Subtour elimination, time windows, and maximum ride time constraints will be implemented by the callbacks of infeasible path constraints, in line with Alyasiry et al. (2019); Rist & Forbes (2021). The details are as follows. As long as we encounter an incumbent solution in the branch-and-bound tree, we create chains or cycles. A chain is a sequence of fragments $(f_{1},\dots,f_{c})$ with a fragment count of $c>1$, along with $c-1$ node arcs $(a_{1},\dots,a_{c-1})$ connecting them. The load of the end node of fragment $f_{i}$ must be the same as that of the start node of $f_{i+1}$ for $11\leq i\leq c-$ 1; the start node and the end node of $a_{i}$ correspond to the end node of fragment $f_{i}$ and the start node of $f_{i+1}$, respectively, for $11\leq i\leq c-$ 1. Similarly, a cycle is a sequence of fragments $(f_{1},\dots,f_{c})$ with a fragment count of $c>1$, along with $c$ node arcs $(a_{1},\dots,a_{c})$ connecting them. A cycle has the same connection requirement as a chain but includes an additional node arc $a_{c}$ that connects $f_{c}$ back to $f_{1}$. If we obtain a chain, we connect its initial and final fragments to the origin and destination depots, respectively, to construct a path. After that, we establish a valid time schedule for the path, following the constraints of time windows and maximum ride time. If no valid schedule can be found, the chain is considered infeasible and then needs to be eliminated by constraints (35), which ensures that the total count of fragments and node arcs in such a chain does not exceed $22c-$ 2. If a cycle is detected, we then remove it using constraints (36) to ensure that the total count of its elements does not exceed $12c-$ 1, thereby achieving subtour elimination. While callbacks are a useful approach for rarely violated constraints, like the constraints mentioned above, they are not ideal for constraints that are easily violated, such as the limit on the number of pickups.
$\sum_{k=1}^{c}x_{f_{k}v}+\sum_{k=1}^{c-1}y_{a_{k}v}$ $\leq 2c-2$ $v\in V$ (35)
$\sum_{k=1}^{c}x_{f_{k}v}+\sum_{k=1}^{c}y_{a_{k}v}$ $\leq 2c-1$ $v\in V$ (36)
In addition to the necessary constraints mentioned earlier, we incorporate three types of inequalities proposed by Rist & Forbes (2021), namely fragment-fragment, arc-fragment, and fragment-arc inequalities, and add the subset-row inequalities proposed by Jepsen et al. (2008). Fragment-fragment inequalities stipulate that when a fragment is traversed, at least one fragment from its preceding and one fragment from its succeeding fragment set must also be traversed. The preceding and succeeding fragment sets include all fragments except the ones that, when combined with the current fragment, would result in infeasible paths. Arc-fragment inequalities state that if a node arc is traversed, at least one fragment from its succeeding and one fragment from its preceding fragment set must be traversed to maintain path feasibility. Fragment-arc inequalities specify that if a fragment is traversed, at least one succeeding node arc and one preceding node arc in the corresponding sets must also be traversed. In addition, for subset-row inequalities, we follow the same approach as Rist & Forbes (2021) and Jepsen et al. (2008).
Furthermore, to enhance the formulations (both FFF and FAF), we introduce additional constraints based on the minimum number of used vehicles, as shown in constraints (37). While these constraints may be redundant, they significantly speed up the computation process. Additionally, for constraints (23), we set $w_{v}$ equal to 1 for $v\in V_{f}$.
$\sum_{f\in F}x_{vf}$ $\geq 1$ $v\in V_{f}$ (37)
3.4 MDARP-LPT: problem description and formulations
The primary distinction between MDARP-LPT and DARP-LPT lies in the fact that in DARP-LPT all drivers must initiate their journeys from the origin depot $O^{+}$ and return to the destination depot $O^{-}$. Conversely, in MDARP-LPT, each vehicle $v\in V$ possesses its own unique origin location $o_{v}$ and destination location $d_{v}$ (denoted as $O_{v}$ and $D_{v}$ for origin and destination nodes, respectively). In all other aspects, MDARP-LPT adheres to the same principles as DARP-LPT.
The formulations developed supra (ABF, FFF, and FAF) for DARP-LPT can also be used for MDARP-LPT, with some modifications as explained below.
For ABF, only the network flow constraints (4)-(6) need to be updated to reflect the unique origin and destination for each vehicle. The complete formulation can be found in A.
For FFF, all constraints (18)-(22) remain applicable, but the network configuration assigns unique origin and destination nodes, fragment sets, node arc sets, and node sets to each vehicle.
For FAF, in addition to the network adjustments, two further constraints, constraints (38) and (39), must be incorporated. These constraints stipulate that when a node arc starting from an origin is utilized, the subsequent fragments must also be employed, and similarly, when a node arc ending at a destination is employed, the preceding fragments must also be utilized. Specifically, for each vehicle, it possesses a unique set of node arcs that originate from its origin and a unique set of node arcs that end at its destination node.
$Y_{a}$ $\leq\sum_{f\in F^{+}_{j}}x_{fv}$ $a=(O_{v},j)\in A_{N},v\in V$ (38)
$Y_{a}$ $\leq\sum_{f\in F^{-}_{j}}x_{fv}$ $a=(j,D_{v})\in A_{N},v\in V$ (39)
Additionally, for both FFF and FAF, the model for MDARP-LPT needs to incorporate infeasible path constraints (constraints (35)) through callback procedures to handle the constraints of subtour elimination, time windows, and maximum ride time. Given that MDARP-LPT involves distinct vehicles with unique origins and destinations, constraints such as (37) based on the minimum number of vehicles are not used for MDARP-LPT.
4 Fragment generation
In this section we describe how the fragments that constitute the basis for the formulations FFF and FAF can be acquired. Section 4.1 describes the FF generation procedure of Alyasiry et al. (2019) and the RF generation method of Rist & Forbes (2021), both of which involve enumeration. The resulting fragment sets for (M)DARP-LPT are denoted as $S_{FF}$ and $S_{RF}$, respectively. From these sets, all fragments whose length exceeds $L$ are already eliminated. The set $S_{FF}$ with longer fragments tends to be larger than $S_{RF}$ with shorter fragments, yet all combinations of RFs from $S_{RF}$ can generate more (redundant) FFs than those in $S_{FF}$.
In DARP, solution paths can vary in length, potentially exceeding $L$. However, in (M)DARP-LPT, solution paths must not exceed this limit. Since $S_{FF}$ and $S_{RF}$ are primarily designed for DARP, some of their fragments (even if not exceeding $L$ in length) might only appear on solution paths that exceed $L$ and not on those that adhere to the limit (because some higher-cost paths are eliminated from consideration; see Section 4.2). Therefore, in Section 4.2, we introduce a novel path-enumerated full fragment set $S_{PFF}$ by enumerating all (M)DARP-LPT paths first and then decomposing them into full fragments. Consequently, $S_{PFF}$ must be a subset of $S_{FF}$. Another novel method in Section 4.3 is to create a mixed fragment set $S_{MF}$ by converting certain FFs from $S_{FF}$ into RFs while keeping others, aiming to leverage the strength of both FFs and RFs and potentially reduce the set size.
All these fragment sets ($S_{FF}$, $S_{RF}$, $S_{PFF}$, and $S_{MF}$) are valid for both MDARP-LPT and DARP-LPT. However, in MDARP-LPT, where vehicles have different origins and destinations, each vehicle requires a unique fragment set (though many of the fragments will still be the same for each vehicle). Therefore, the generation procedure for MDARP-LPT must be executed $|V|$ times, where $|V|$ denotes the number of vehicles. In contrast, DARP-LPT, with a fixed depot for all vehicles, only necessitates a single execution of the generation procedure.
4.1 Traditional fragment generation
Alyasiry et al. (2019) introduce FF and $S_{FF}$, while Rist & Forbes (2021) propose RF and $S_{RF}$. $S_{FF}$ exclusively contains FFs, while $S_{RF}$ exclusively contains RFs. Their intersection includes all FFs with the format “$ppppdddd$”, which can also be regarded as RFs. As previously explained, “$ppppdddd$” indicates a sequence where all pickups precede all deliveries.
[…]
4.2 Path-enumerated full fragment generation
Generating $S_{PFF}$ involves three steps: enumerating all paths that adhere to the problem constraints, selecting enumerated paths, and generating fragments from these selected paths. The first step is to enumerate all possible paths. Typically, such a task can be challenging due to the exponential growth in the number of paths. With limited pickups, however, this number is polynomial for $L$. To be more precise, the number of paths is bounded by $\mathcal{O}((2n)^{2L})$. The second step involves path selection. For any pair of enumerated paths that visit an identical set of locations, only the one with lower cost is kept. The third step involves extracting full fragments from the selected paths. Within each path, nodes with empty loads are identified and marked. The route paths delimited by consecutive markers are then extracted to create fragments, which are subsequently recorded in $S_{PFF}$. Notably, $S_{PFF}$ should be a subset of $S_{FF}$, as $S_{PFF}$ is derived from the selected paths, and not all FFs from $S_{FF}$ will appear on these paths. Since $S_{PFF}$ includes all FFs obtained from all non-dominated (M)DARP-LPT paths (paths with higher cost are excluded), these FFs can also be re-combined to reconstitute all non-dominated paths.
We illustrate the foregoing with an example. Consider an instance with only two customers, 1 and 2. The set $S_{FF}$ includes the following fragments: $(p1,p2,d1,d2)$, $(p1,p2,d2,\\ d1)$, $(p2,p1,d1,d2)$, $(p2,p1,d2,d1)$, $(p1,d1)$, and $(p2,d2)$. However, there are at most three useful paths: $(0,p1,d1,2n+1)$, $(0,p2,d2,2n+1)$, and then out of the four paths that include the four locations $p1$, $p2$, $d1$, and $d2$ and serve both customers, only one of the four (a shortest path) needs to be retained. As a result, $S_{PFF}$ can contain at most three fragments from the three paths and is thus clearly a subset of $S_{FF}$.
4.3 Mixed fragment generation
Path enumeration can become significantly more time-consuming as the limit on the number of pickups $L$ increases. Hence, we present an alternative procedure for creating fragments as $L$ grows. In this procedure, we initially create all FFs to get $S_{FF}$. Next, we employ a selection process to identify specific FFs to be decomposed into RFs. Finally, the remaining FFs together with the newly formed RFs constitute the set $S_{MF}$.
The first step generates FFs adhering to the (M)DARP-LPT constraints, according to the procedure described in Section 4.1, resulting in the same fragment set $S_{FF}$.
The second step decomposes all FFs into RFs and records the frequency of occurrence of each RF during this process. In a FF, pickup nodes whose preceding nodes are depots or delivery nodes, and delivery nodes whose subsequent nodes are depots or pickup nodes, are marked. The route paths between consecutive marks can be utilized to create RFs. During the process of obtaining RFs from FFs, we keep track of how frequently each RF appears in all FFs. This work is preparatory for the third step. An important aspect to note is that the resulting full RF set should be a subset of $S_{RF}$. This can be seen as follows: $S_{RF}$ encompasses all RFs obtained by partitioning FFs with the “ppppdddd” format in all possible ways. However, some RFs in $S_{RF}$ may never appear in any FF in $S_{FF}$. Specifically, these RFs might only be combined with specific other RFs to form FFs that are dominated.
The third step involves selecting FFs for decomposition, aiming to create a mixed fragment set comprising both FFs and RFs. The main objective of working with such a mixed set is to reduce the fragment count. An initial assumption is that $S_{FF}$ typically contains more fragments than $S_{RF}$. Intuitively, decomposing all FFs into RFs might seem beneficial. However, we find that if an FF includes several RFs and these RFs do not appear in any other non-dominated FFs in $S_{FF}$, decomposing it can increase the total fragment count. Furthermore, our computational results in Section 5.2.1 indicate that when RFs are used in the formulations, they may lead to weaker lower bounds compared to FFs. Based on these two observations, we only decompose a portion of FFs in $S_{FF}$. Specifically, we maintain a record of how frequently each RF appears in the FFs from the previous step. Only the FFs for which all the constituent RFs appear multiple times (at least three times, in our implementation) within all FFs in $S_{FF}$ are candidates for decomposition. Other FFs that do not meet this criterion are retained in their original form. As a result, the undecomposed FFs together with the RFs obtained through this process constitute $S_{MF}$. Thus, it can be seen that $S_{MF}$ can serve the same role as $S_{FF}$ in enumerating all (M)DARP-LPT paths and act as the input for formulations without compromising optimality.
5 Computational results
5.1 Data generation and implementation details
To assess computational efficiency of different formulations and fragment sets, we utilize benchmark instances from Cordeau (2006), which are widely adopted in studies such as Ropke et al. (2007); Gschwind & Irnich (2015); Rist & Forbes (2021). These instances comprise two types, A and B, with type A having a tighter maximum ride time. In type A, customer demand is unit with a vehicle capacity of three, while type B involves demands from one to six customers with a vehicle capacity of six.
For DARP-LPT, we generate both basic and extended instances. The basic instances adjust the vehicle numbers $|V|$ in Cordeau’s instances. Given that DARP-LPT incorporates an additional parameter $L$ for the maximal number of pickups per trip, ensuring coverage for $n$ customers requires a minimum of $\lceil n/L\rceil$ vehicles. Here, we set $|V|$ as $\lceil n/(L-1)\rceil$, which is slightly higher than the minimum required. We also create an extra instance set to which we refer as “extended” instances, following Gschwind & Irnich, in which we increase time windows and capacity by 1/3 relative to the basic instances, thereby enhancing the diversity of the problem instances.
For MDARP-LPT, which assumes different drivers have distinct origins and destinations, we generate driver information based on Cordeau’s instances. Specifically, the first $\lceil n/\alpha\rceil$ customers are designated as drivers, whose pickup and delivery nodes becoming the origins and destinations, while the remaining customers are passengers. The value $n$ represents the number of customers in each Cordeau instance, and $\alpha$ is an auxiliary parameter. The choice $L=4$ and $\alpha=4$ implies that one driver serves three customers on average, and $\alpha=3$ has one driver per two customers on average. For drivers’ time windows, we set the earliest departure to 0, maintaining latest arrival times. Their maximum ride time is halved from the original driver’s maximum ride time (which is between six and 12 hours), as the latter is excessive for realistic scenarios such as carpooling. Unit customer demand and capacity 4 are chosen.
Table 4 summarizes our experimental design. We use $S_{PFF}$ for MDARP-LPT and $S_{MF}$ for DARP-LPT. This decision is based on the appropriateness of $S_{PFF}$ for low $L$ values, which aligns well with real carpooling scenarios in MDARP-LPT. On the other hand, $S_{MF}$ is used for comparison with $S_{RF}$ and $S_{FF}$ across diverse $L$ values, a context more suitable for DARP-LPT. We do not apply $S_{MF}$ for MDARP-LPT because each vehicle has distinct fragment sets, and mixing the FFs and RFs per vehicle would become overly complex.
Table 4: Experimental design
[b] Problem Instance $L$ $|V|$ Customer number Formulation Fragment set DARP-LPT basic $4,6$ $\lceil n/(L-1)\rceil$ $n$ ABF, FAF, FFF $S_{FF}$, $S_{RF}$, $S_{MF}$ extended MDARP-LPT basic 4 $\lceil n/\alpha\rceil$ $n-\lceil n/\alpha\rceil$ FFF $S_{FF}$, $S_{RF}$, $S_{PFF}$
  • 1
    Value $n$ is the number of customers in the original instance of Cordeau (2006).
All our algorithms are implemented using the Python programming language. All computational experiments are run on a system with an Intel Core i7-7820HQ processor with 2.90 GHz CPU speed and 32 GB of RAM under a Windows 10 64-bit OS. All linear formulations are solved with the commercial solver Gurobi 9.0.3 with a single thread; all other Gurobi parameters are set to their default values. The generation of fragments is implemented using the Python programming language. For DARP-LPT, fragment generation occurs in a single-processing environment, whereas for the MDARP-LPT, fragment generation is accomplished via multi-processing.
Since this study involves two problems (DARP-LPT and MDARP-LPT), two different fragment-based formulations, four types of fragment sets, and two types of instances (basic and extended), we need to test multiple combinations. Therefore, from the existing benchmark instances, we only choose a subset, but we ensure a variety of sizes as indicated by the instance names in all following tables.
5.2 Discussion of the results
In our evaluation of computational performance, we incorporate several performance indicators: time, OBJ, LB, and gap. Time (s) indicates the total time, including the preprocessing time, network time, and CPU time for solver execution. Here, preprocessing time refers to the duration spent on time window tightening and arc elimination, as detailed by Cordeau (2006). Network time refers to the time required for fragment generation and the construction of the fragment-based network. OBJ and LB are the best-found integer solution value and the best lower bound within a 30-minute timeframe, respectively, and gap represents the relative percentage difference between OBJ and LB. When comparing different fragment sets, we also introduce two additional indicators: fragment count and network time. Fragment count is the total number of fragments in the fragment set.
5.2.1 Results for DARP-LPT
Table 5: Computational performance of different formulations for DARP-LPT within 30 minutes
[b] ABF FAF ($S_{FF}$) FFF ($S_{FF}$) Name Time (s) OBJ LB Gap Time (s) OBJ LB Gap Time (s) OBJ LB Gap a3-18 24.3 302.5 302.5 0.00% 0.2 302.5 302.5 0.00% 0.2 302.5 302.5 0.00% a3-30 1800.0 496.7 451.2 9.16% 1800.0 494.8 491.6 0.64% 1800.0 494.8 491.1 0.74% a3-36 1800.0 574.4 541.2 5.78% 1800.0 574.4 565.9 1.49% 1800.0 574.4 565.0 1.63% a4-48 1800.0 712.0 578.3 18.78% 1800.0 698.6 685.4 1.90% 1800.0 697.9 687.4 1.50% a6-60 1800.0 920.2 635.4 30.95% 1800.0 879.2 843.2 4.09% 1800.0 864.9 843.4 2.48% a7-70 1800.0 1284.3 664.6 48.25% 1800.0 1004.0 958.2 4.56% 1800.0 986.6 956.5 3.05% a8-80 1800.0 1359.6 670.4 50.69% 1800.0 1100.7 1017.9 7.53% 1800.0 1071.1 1019.3 4.84% Avg-A 1546.3 23.37% 1542.9 2.89% 1542.9 2.03% b3-36 1800.0 576.5 548.4 4.88% 1800.0 576.5 566.8 1.69% 1800.0 576.5 567.7 1.53% b4-40 1800.0 637.5 612.5 3.92% 1800.0 635.2 630.2 0.79% 1800.0 635.2 631.1 0.64% b5-50 1800.0 796.5 659.3 17.22% 1800.0 798.4 765.7 4.10% 1800.0 789.1 768.3 2.63% b6-60 1800.0 937.9 749.9 20.05% 1800.0 940.6 896.5 4.68% 1800.0 931.7 900.7 3.33% b7-70 1800.0 1162.9 799.9 31.21% 1800.0 965.3 919.4 4.75% 1800.0 950.4 920.7 3.13% Avg-B 1800.0 15.46% 1800.0 3.20% 1800.0 2.25%
  • 1
    The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
Table 5 compares the three formulations ABF, FAF, and FFF for DARP-LPT for a runtime limit of 30 minutes, using the basic instances with the parameters $|V|=\lceil n/3\rceil$ and $L=4$. Both FAF and FFF utilize $S_{FF}$ as the input fragment set. The table indicates that, as observed from the time and gap indicators, both FFF and FAF largely outperform ABF, with FFF exhibiting superior performance compared to FAF.
$S_{RF}$ $S_{FF}$ $S_{MF}$ Type F Net Time Gap F Net Time Gap F Net Time Gap Basic, $L=4$ A 475.7 0.4 1542.9 2.17% 567.3 1.4 1542.9 2.03% 452.0 1.7 1543.0 1.92% B 579.6 0.4 1800.0 2.42% 808.0 1.5 1800.0 2.25% 506.0 1.9 1800.0 2.32% Basic, $L=6$ A 475.7 0.4 1107.8 0.86% 937.3 5.8 1034.5 0.68% 454.6 5.8 1057.9 0.65% B 579.6 0.4 1444.6 0.57% 2449.2 10.2 1422.5 0.47% 511.8 11.1 1443.4 0.45% Extended, $L=4$ A 661.0 0.5 1542.9 2.44% 875.3 2.2 1542.9 2.15% 612.7 3.0 1543.0 2.17% B 1914.0 1.7 1800.0 3.67% 3070.0 8.7 1800.0 2.87% 1376.0 8.5 1800.0 2.94% Extended, $L=6$ A 661.0 0.5 1031.5 0.99% 1807.3 15.0 1031.7 0.91% 620.6 14.6 1030.8 0.94% B 1914.0 1.9 1228.9 1.08% 20382.6 204.3 1104.4 1.25% 1495.6 202.8 1191.7 0.96% 1 Type: instance type; F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver; basic: basic instances; extended: extended instances.
Table 6: Average computational performance of FFF across all instances using various fragment sets for DARP-LPT within a 30-minute timeframe
Table 6 presents a comparative analysis of three fragment set types $S_{RF}$, $S_{FF}$, and $S_{MF}$. These fragment sets are utilized as input for formulation FFF for solving DARP-LPT. The table covers both basic and extended instances, with $|V|=\lceil n/3\rceil$ when $L=4$ and $|V|=\lceil n/5\rceil$ when $L=6$. Due to the extensive amount of results, we present only the average metrics (fragment count, network time, total time, and gap) for all instances here. Detailed results for basic and extended instances with $L=4$ and $L=6$ are available in Tables 1, 2, 3, and 4 in the Appendix.
We begin by comparing the two traditional fragment sets $S_{FF}$ and $S_{RF}$. The set $S_{FF}$ is better than $S_{RF}$ in most scenarios, particularly excelling in scenarios with a limited number of pickups per trip (e.g., $L=4$), flexible time windows (extended instances), and increased maximum ride time (instance type B), as reflected in the columns Gap. However, for extended type B instances with $L=6$, $S_{RF}$ may outperform $S_{FF}$. This variation can be attributed to fragment count and the quality of lower bound. When $L$ is lower, or when time windows or maximum ride time are tight, the fragment counts for $S_{FF}$ and $S_{RF}$ are similar or at least of the same magnitude. Conversely, when $L$ is higher and both time windows and maximum ride time are more flexible, the fragment count for $S_{FF}$ is substantially higher than that for $S_{RF}$. This is important because employing a large fragment set as input can slow down the computation process. Additionally, the fragments from $S_{FF}$ are longer than those from $S_{RF}$, in a similar way as paths are longer than individual arcs. This similarity helps to explain why $S_{FF}$ generally achieves better lower bounds than $S_{RF}$, because path-based (extended) formulations for routing problems typically yield superior lower bounds compared to arc-based formulations (Desrosiers et al., 2024).
The mixed fragment set $S_{MF}$ demonstrates superior solution quality in scenarios with more flexible time windows and maximum ride time and higher values of $L$ (extended type B instances with $L=6$), as indicated by the gap metrics. Notably, the fragment count of $S_{MF}$ remains relatively low, especially when compared to $S_{FF}$. Since a larger number of fragments can slow down fragment-based formulations, $S_{MF}$ presents itself as an efficient alternative to $S_{FF}$ in this regard.
5.2.2 Results for MDARP-LPT
$S_{RF}$ $S_{FF}$ $S_{MF}$ $|V|$ Type F Net Time Gap F Net Time Gap F Net Time Gap $\lceil n/4\rceil$ A 269.1 8.2 813.9 1.50% 274.4 9.5 806.4 1.19% 162.7 69.7 777.4 0.82% B 1489.2 14.8 1471.1 3.02% 1650.6 20.5 1210.7 1.91% 674.4 96.6 797.5 0.78% $\lceil n/3\rceil$ A 201.9 10.0 636.5 0.56% 206.4 10.9 464.3 0.59% 124.3 52.7 423.5 0.31% B 1195.0 16.7 1107.3 1.80% 1227.6 20.0 872.0 0.55% 518.8 90.9 552.6 0.36% 1 Type: instance type; F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
Table 7: Average computational performance of FFF across all instances using various fragment sets for MDARP-LPT within a 30-minute timeframe (basic instances with $L=4$)
Table 7 compares three types of fragment sets for MDARP-LPT using FFF under different parameters. Formulation FAF for DARP-LPT can be extended with additional constraints to handle also MDARP-LPT, but its performance significantly lags behind that of FFF, which is why we only present results for FFF and not FAF. The reason for not altering $L$ is that $L=4$ represents a commonly used parameter in carpooling scenarios (the realistic background of MDARP-LPT). Here, we present only the average metrics for all instances. Detailed results can be found in Table 1 and 2 in the Appendix.
Table 7 demonstrates the superiority of $S_{PFF}$ over $S_{RF}$ and $S_{FF}$ in computational performance, particularly for handling large instances. This is evident from the consistently lower time and gap values achieved with $S_{PFF}$ compared to $S_{FF}$. Notably, $S_{PFF}$’s superiority can likely be attributed to its lower fragment count. Our findings also indicate that across most instances, $S_{FF}$ generally outperforms $S_{RF}$ in terms of both time efficiency and solution quality.
6 Summary and conclusions
In this paper we have introduced two variants of the dial-a-ride problem, specifically MDARP-LPT and DARP-LPT. MDARP-LPT addresses scenarios where vehicles have unique origins and destinations, while DARP-LPT assumes that all vehicles begin and end their routes at a fixed depot. These problem variants are motivated by real-world carpooling and crowdshipping scenarios, where each trip can pick up at most a certain number $L$ of customers. (M)DARP-LPT has been proven to be NP-hard through a reduction from the capacitated vehicle routing problem.
To efficiently solve both MDARP-LPT and DARP-LPT, we have applied a unified fragment-based method, where a fragment is a partial path. Our main contributions entail the extensions of a fragment-based formulation, the creation of improved fragment sets, and the insights gained from comparing fragment sets. More specifically, we have modeled (M)DARP-LPT using two fragment-based formulations, FFF and FAF, both of which are extended from the DARP fragment-based formulation of Rist & Forbes (2021), transitioning from fragment-based decision variables to fragment-vehicle-based variables. We find that FFF achieves better results than FAF in terms of solving time and optimality gap, which confirms our theoretical finding that FFF will also yield better lower bounds. We have also introduced improved fragment sets as formulation inputs. The two benchmark fragment sets, the full fragments set $S_{FF}$ and the restricted fragments set $S_{RF}$, are generated using the methods described by Alyasiry et al. (2019) and Rist & Forbes (2021), respectively. Our two novel sets are the path-enumerated full fragment set $S_{PFF}$ and the mixed fragment set $S_{MF}$. The generation of $S_{PFF}$ involves exhaustively enumerating paths, selectively choosing those meeting specific criteria, and then decomposing them into full fragments. This approach draws inspiration from the observation that carpooling and crowdshipping DARP routes typically involve a limited number of customers. We construct the set $S_{MF}$ by decomposing some of the full fragments in $S_{FF}$ into restricted fragments while preserving the remainder. This approach is designed for scenarios when $L$ is higher. All these fragment sets can be employed as input for fragment-based formulations to reach optimality. Our computational results demonstrate that for MDARP-LPT with $L=4$, a realistic carpooling setting, $S_{PFF}$ outperforms $S_{FF}$, and $S_{FF}$ performs better than $S_{RF}$ in terms of computational performance. For DARP-LPT, $S_{FF}$ generally outperforms $S_{RF}$ except when the instances have more flexible time windows, increased maximum ride times, and higher $L$ values. In such cases, $S_{MF}$ can replace $S_{FF}$ to achieve the best performance.
We conclude with two supplementary observations. First, in DARP-LPT, FAF may outperform FFF for a number of small instances, as shown in Table 5. This advantage becomes more prevalent when a fixed number of vehicles is utilized. The underlying reason for this, intuitively, is that FAF, having fewer variables than FFF, solves its linear relaxations more quickly, but produces worse lower bounds. Thus, in small instances with fairly simple branch-and-bound trees, FAF tends to perform better. Conversely, in large instances, FAF’s performance tends to be worse than FFF. Additionally, when the number of vehicles is fixed, the complexity of the problem is reduced, allowing FAF to perform better than FFF in more scenarios. Second, the technique of variable fixing could be a viable option to reduce the number of fragments by removing those whose reduced cost is higher than the gap between a feasible solution’s objective value and the linear relaxation bound (Desaulniers et al., 2020). In the instances that we have tested, however, even if we set the objective value as the optimal value, this method typically removes only a small number of fragments, which hardly affects the final computational results.
Appendix A Arc-based formulation of MDARP-LPT
For MDARP-LPT, the full arc-based formulation (following Cordeau (2006)) is as follows:
Table 1: Sets, parameters, and decision variables of arc-based formulation for MDARP-LPT
Sets Definition
$N_{v}$ Set of locations, $N=P_{v}\cup D_{v}\cup\{o_{v},d_{v}\}$, where $o_{v}$ and $d_{v}$ are respectively
origin and destination of vehicle $v\in V$
$P_{v}$, $D_{v}$ Pickup and delivery locations of vehicle $v\in V$
$A_{v}$ Set of location arcs of vehicle $v\in V$
Parameters Definition
$e_{iv}$, $l_{iv}$ The earliest and latest time of vehicle $v$ at location $i\in N$
Decision variables Definition
$f_{ijv}$ =1, if the arc $(i,j)\in A_{v}$ is traversed by vehicle $v\in V$; =0, otherwise
$t_{iv}$ the departure time of vehicle $v$ on location $i\in N_{v}$
$Q_{iv}$ the load of vehicle $v$ after visiting location $i\in N_{v}$
$\min\sum_{v\in V}\sum_{(i,j)\in A_{v}}C_{ij}f_{ijv}$ (40)
$s.t.$ $\sum_{v\in V}\sum_{j\in N_{v}}f_{ijv}$ $=1$ $i\in P$ (41)
$\sum_{j\in N_{v}}f_{ijv}$ $=\sum_{j\in N_{v}}f_{(i+n)jv}$ $i\in P_{v},v\in V$ (42)
$\sum_{j\in N_{v}}f_{o_{v}jv}$ $=1$ $v\in V$ (43)
$\sum_{j\in N_{v}}f_{jiv}$ $=\sum_{j\in N_{v}}f_{ijv}$ $i\in P_{v}\cup D_{v},v\in V$ (44)
$\sum_{i\in N_{v}}f_{id_{v}v}$ $=1$ $v\in V$ (45)
$t_{iv}+T_{ij}$ $\leq t_{jv}+M_{ijv}(1-f_{ijv})$ $(i,j)\in A_{v},v\in V$ (46)
$e_{iv}\leq t_{iv}$ $\leq l_{iv}$ $i\in N_{v},v\in V$ (47)
$Q_{iv}+q_{i}$ $\leq Q_{jv}+W_{ij}(1-f_{ijv})$ $(i,j)\in A_{v},v\in V$ (48)
$\max(0,q_{i})\leq Q_{iv}$ $\leq\min(Q,Q+q_{i})$ $i\in N_{v},v\in V$ (49)
$T_{i(n+i)}\leq t_{(n+i)v}-t_{iv}$ $\leq R_{i}$ $i\in P_{v},v\in V$ (50)
$t_{(2n+1)v}-t_{0v}$ $\leq R_{0}$ $v\in V$ (51)
$\sum_{(i,j)\in A_{v}}f_{ijv}$ $\leq 2L+1$ $v\in V$ (52)
$f_{ijv}$$\in\{0,1\}$ $(i,j)\in A_{v},v\in V$ (53)
To be concise, we outline the differences between the formulations of MDARP-LPT and DARP-LPT. Constraints (4) to (6) in DARP-LPT are modified to constraints (43) to (45) in MDARP-LPT, accommodating the network flow constraints for distinct vehicles with unique origins and destinations. Constraints (52) can be modified to $\sum_{(i,j)\in A_{v}}\max(q_{i},0)f_{ijv}\\ \leq 2\lambda+1$ if the constraint on the maximum number of pickups per trip is replaced by the constraint on the maximum quantity of items (or passengers) $\lambda$ per trip, where $q_{i}$ represents the customer load quantity. Importantly, the constraints on $\lambda$ remain applicable within our fragment-based approach. The necessary modification involves updating constraint (21) in FFF and constraint (32) in FAF from $\sum_{f\in F}L_{f}x_{fv}\leq L$ to $\sum_{f\in F}\lambda_{f}x_{fv}\leq\lambda$, where $\lambda_{f}$ represents the total quantity of items in fragment $f$.
Appendix B Full computational results for DARP-LPT
All four tables below serve as supplements to Table 6.
$S_{RF}$ $S_{FF}$ $S_{MF}$ Name F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 73 0.3 0.0 0.4 302.5 302.5 0.00% 64 0.1 0.0 0.2 302.5 302.5 0.00% 64 0.3 0.2 1.0 302.5 302.5 0.00% a3-30 93 0.1 1799.9 1800.0 494.8 491.0 0.75% 89 0.1 1799.9 1800.0 494.8 491.1 0.74% 89 0.5 1798.8 1800.0 494.8 489.7 1.01% a3-36 121 0.1 1800.5 1800.0 574.4 560.6 2.40% 112 0.2 1797.6 1800.0 574.4 565.0 1.63% 113 0.1 1799.8 1800.0 574.4 566.7 1.35% a4-48 318 0.2 1798.9 1800.0 697.9 686.1 1.69% 312 0.5 1799.4 1800.0 697.9 687.4 1.50% 303 0.5 1798.8 1800.0 697.9 687.9 1.43% a6-60 601 0.4 1793.4 1800.0 869.0 843.2 2.97% 648 1.4 1795.5 1800.0 864.9 843.4 2.48% 579 1.8 1795.7 1800.0 864.9 844.8 2.32% a7-70 1036 0.8 1796.3 1800.0 986.3 958.5 2.82% 1273 2.6 1792.3 1800.0 986.6 956.5 3.05% 932 2.9 1795.1 1800.0 984.9 956.5 2.88% a8-80 1088 0.8 1789.1 1800.0 1064.6 1016.4 4.53% 1473 5.1 1785.7 1800.0 1071.1 1019.3 4.84% 1084 6.2 1782.4 1800.0 1063.4 1016.3 4.43% Avg-A 475.7 0.4 1542.9 2.17% 567.3 1.4 1542.9 2.03% 452.0 1.7 1543.0 1.92% b3-36 169 0.1 1799.9 1800.0 576.5 566.2 1.79% 161 0.1 1799.7 1800.0 576.5 567.7 1.53% 146 0.1 1799.8 1800.0 576.5 566.1 1.81% b4-40 311 0.2 1799.5 1800.0 635.2 630.4 0.76% 203 0.4 1799.1 1800.0 635.2 631.1 0.64% 202 0.4 1799.5 1800.0 635.2 631.8 0.54% b5-50 518 0.4 1798.6 1800.0 789.1 766.2 2.91% 803 1.2 1797.6 1800.0 789.1 768.3 2.63% 496 2.2 1796.2 1800.0 789.5 766.7 2.90% b6-60 940 0.6 1793.3 1800.0 931.4 895.5 3.86% 1218 2.6 1792.3 1800.0 931.7 900.7 3.33% 799 2.0 1794.2 1800.0 931.7 900.4 3.36% b7-70 960 0.7 1796.4 1800.0 948.0 921.6 2.79% 1655 3.2 1789.5 1800.0 950.4 920.7 3.13% 887 4.6 1789.1 1800.0 950.2 921.7 3.00% Avg-B 579.6 0.4 1800.0 2.42% 808.0 1.5 1800.0 2.25% 506.0 1.9 1800.0 2.32% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
Table 1: Computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes (Basic instances, $L=4$)
$S_{RF}$ $S_{FF}$ $S_{MF}$ Name F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 73 0.2 0.3 1.0 295.8 295.8 0.00% 64 0.1 0.0 0.2 295.8 295.8 0.00% 64 0.1 0.0 0.2 295.8 295.8 0.00% a3-30 93 0.3 17.5 18.9 477.0 477.0 0.00% 89 0.1 4.5 4.7 477.0 477.0 0.00% 89 0.2 4.4 4.7 477.0 477.0 0.00% a3-36 121 0.1 534.4 535.0 549.8 549.8 0.00% 115 0.2 36.0 36.4 549.8 549.8 0.00% 120 0.2 200.2 200.6 549.8 549.8 0.00% a4-48 318 0.3 1798.6 1800.0 669.4 662.9 0.97% 342 0.7 1798.9 1800.0 669.4 664.4 0.74% 305 0.7 1798.3 1800.0 669.4 665.5 0.58% a6-60 601 0.4 1798.0 1800.0 825.0 819.0 0.72% 817 4.5 1796.3 1800.0 825.0 820.3 0.57% 589 3.6 1794.6 1800.0 825.0 820.6 0.54% a7-70 1036 0.7 1793.6 1800.0 926.7 912.4 1.54% 2114 10.8 1786.0 1800.0 927.6 912.8 1.60% 944 10.7 1786.6 1800.0 927.6 913.1 1.56% a8-80 1088 0.7 1790.5 1800.0 1000.9 973.1 2.77% 3020 24.2 1770.1 1800.0 991.3 973.1 1.83% 1071 25.3 1767.5 1800.0 992.8 974.0 1.89% Avg-A 475.7 0.4 1107.8 0.86% 937.3 5.8 1034.5 0.68% 454.6 5.8 1057.9 0.65% b3-36 169 0.1 22.7 23.1 557.9 557.9 0.00% 210 0.3 13.5 14.0 557.9 557.9 0.00% 154 0.3 16.2 16.8 557.9 557.9 0.00% b4-40 311 0.2 1799.3 1800.0 627.4 623.9 0.56% 228 1.5 1696.5 1698.5 627.4 627.4 0.00% 216 1.7 1797.9 1800.0 627.4 626.2 0.20% b5-50 518 0.3 1798.7 1800.0 751.9 745.3 0.87% 1868 5.9 1791.6 1800.0 751.9 744.2 1.03% 492 7.1 1793.4 1800.0 751.9 745.7 0.82% b6-60 940 0.6 1795.9 1800.0 877.0 868.0 1.03% 3769 14.7 1782.2 1800.0 877.0 867.1 1.13% 806 14.7 1783.4 1800.0 877.0 867.8 1.05% b7-70 960 0.7 1798.2 1800.0 896.4 892.9 0.39% 6171 28.5 1766.2 1800.0 896.4 894.6 0.21% 891 31.7 1765.5 1800.0 896.4 894.8 0.19% Avg-B 579.6 0.4 1444.6 0.57% 2449.2 10.2 1422.5 0.47% 511.8 11.1 1443.4 0.45% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
Table 2: Computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes (Basic instances, $L=6$)
$S_{RF}$ $S_{FF}$ $S_{MF}$ Name F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 96 0.1 0.3 0.4 288.5 288.5 0.00% 81 0.1 0.5 0.6 288.5 288.5 0.00% 81 0.1 0.5 0.7 288.5 288.5 0.00% a3-30 106 0.1 1799.8 1800.0 491.4 486.6 0.97% 101 0.1 1799.7 1800.0 491.4 485.9 1.12% 101 0.1 1799.7 1800.0 491.4 486.7 0.96% a3-36 161 0.1 1799.5 1800.0 541.9 532.2 1.79% 159 0.2 1798.7 1800.0 541.9 535.4 1.20% 155 0.2 1799.7 1800.0 541.9 536.4 1.01% a4-48 420 0.2 1798.2 1800.0 697.8 680.2 2.52% 431 0.6 1798.0 1800.0 697.8 680.9 2.42% 407 0.7 1798.6 1800.0 697.9 679.9 2.58% a6-60 859 0.6 1793.8 1800.0 846.1 828.9 2.04% 1026 2.3 1794.7 1800.0 847.5 830.6 1.99% 830 2.2 1793.9 1800.0 846.1 826.6 2.31% a7-70 1502 1.2 1792.9 1800.0 971.5 939.8 3.26% 2055 5.0 1789.1 1800.0 968.3 941.4 2.78% 1285 6.8 1786.7 1800.0 968.9 940.5 2.93% a8-80 1483 1.1 1788.7 1800.0 1073.8 1004.4 6.46% 2274 7.0 1782.7 1800.0 1065.2 1006.0 5.56% 1430 10.7 1771.7 1800.0 1060.3 1003.0 5.41% Avg-A 661.0 0.5 1542.9 2.44% 875.3 2.2 1542.9 2.15% 612.7 3.0 1543.0 2.17% b3-36 460 0.3 1799.4 1800.0 556.5 540.9 2.80% 567 0.6 1799.1 1800.0 556.5 542.9 2.45% 435 0.7 1797.7 1800.0 556.5 544.0 2.25% b4-40 910 0.6 1798.1 1800.0 582.8 571.0 2.02% 906 2.1 1797.4 1800.0 582.8 573.4 1.61% 614 1.8 1797.2 1800.0 582.8 573.8 1.55% b5-50 1686 1.4 1792.0 1800.0 755.1 730.2 3.30% 3196 12.4 1785.3 1800.0 754.7 732.3 2.97% 1348 7.7 1789.4 1800.0 754.7 731.5 3.07% b6-60 3036 2.9 1790.3 1800.0 873.8 837.0 4.21% 3890 9.6 1781.1 1800.0 868.9 841.7 3.13% 1929 10.5 1782.0 1800.0 874.6 842.0 3.73% b7-70 3478 3.4 1753.1 1800.0 911.6 856.7 6.02% 6791 18.8 1762.9 1800.0 900.4 862.7 4.19% 2554 21.8 1755.4 1800.0 897.1 860.5 4.09% Avg-B 1914.0 1.7 1800.0 3.67% 3070.0 8.7 1800.0 2.87% 1376.0 8.5 1800.0 2.94% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
Table 3: Computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes (Extended instances, $L=4$)
$S_{RF}$ $S_{FF}$ $S_{MF}$ Name F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 96 0.0 0.9 1.0 282.4 282.4 0.00% 88 0.1 0.6 0.8 282.4 282.4 0.00% 90 0.3 1.9 2.5 282.4 282.4 0.00% a3-30 106 0.1 11.9 12.2 472.5 472.5 0.00% 101 0.1 9.4 9.7 472.5 472.5 0.00% 101 0.3 8.2 8.7 472.5 472.5 0.00% a3-36 161 0.1 6.9 7.5 510.6 510.6 0.00% 179 0.4 10.6 11.6 510.6 510.6 0.00% 158 0.3 3.6 4.2 510.6 510.6 0.00% a4-48 420 0.3 1798.6 1800.0 661.2 656.5 0.70% 501 1.3 1797.6 1800.0 661.2 657.7 0.52% 404 1.3 1797.6 1800.0 661.2 657.4 0.57% a6-60 859 0.6 1797.8 1800.0 807.1 796.7 1.29% 1623 9.6 1790.0 1800.0 807.1 797.7 1.17% 833 9.8 1788.7 1800.0 807.1 797.6 1.18% a7-70 1502 1.1 1794.6 1800.0 908.5 891.3 1.89% 4266 26.8 1766.3 1800.0 907.9 891.7 1.79% 1341 28.1 1767.3 1800.0 908.5 892.6 1.75% a8-80 1483 1.1 1782.0 1800.0 988.3 958.3 3.03% 5893 66.6 1720.0 1800.0 986.9 958.2 2.90% 1417 62.1 1727.7 1800.0 989.3 959.1 3.06% Avg-A 661.0 0.5 1031.5 0.99% 1807.3 15.0 1031.7 0.91% 620.6 14.6 1030.8 0.94% b3-36 460 0.2 9.0 9.7 530.6 530.6 0.00% 1199 2.5 18.6 22.1 530.6 530.6 0.00% 436 2.2 12.0 14.9 530.6 530.6 0.00% b4-40 910 0.7 733.4 734.9 564.6 564.6 0.00% 2247 18.0 80.7 99.9 564.6 564.6 0.00% 676 18.5 524.5 543.8 564.6 564.6 0.00% b5-50 1686 1.4 1795.4 1800.0 710.8 701.4 1.33% 17112 90.7 1699.1 1800.0 710.8 703.0 1.10% 1420 94.8 1701.4 1800.0 710.8 702.8 1.14% b6-60 3036 3.1 1793.6 1800.0 817.8 799.3 2.26% 22909 177.8 1600.9 1800.0 818.7 799.3 2.38% 2135 185.2 1609.2 1800.0 818.3 804.5 1.69% b7-70 3478 4.1 1778.4 1800.0 837.7 822.7 1.79% 58446 732.6 957.9 1800.0 844.8 821.3 2.78% 2811 713.3 1075.2 1800.0 841.0 824.5 1.97% Avg-B 1914.0 1.9 1228.9 1.08% 20382.6 204.3 1104.4 1.25% 1495.6 202.8 1191.7 0.96% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
Table 4: Computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes (Extended instances, $L=6$)
Appendix C Full computational results for MDARP-LPT
Both tables below serve as supplements to Table 7. One point worth noting is that while the instances used in both tables are derived from Cordeau, they are not identical since we modify the number of vehicles and the number of customers (see Table 4 for details).
$S_{RF}$ $S_{FF}$ $S_{PFF}$ F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 45 8.9 0.5 9.4 239.6 239.6 0.00% 40 6.2 0.4 6.6 239.6 239.6 0.00% 35 17.9 0.7 18.7 239.6 239.6 0.00% a3-30 42 7.9 1.4 9.4 357.9 357.9 0.00% 42 9.6 2.1 11.9 357.9 357.9 0.00% 40 10.2 1.2 11.4 357.9 357.9 0.00% a3-36 51 6.5 47.1 53.7 474.8 474.8 0.00% 48 8.0 31.3 39.5 474.8 474.8 0.00% 37 11.1 38.1 49.2 474.8 474.8 0.00% a4-48 182 6.8 217.5 224.7 499.1 499.1 0.00% 166 6.9 179.6 186.8 499.1 499.1 0.00% 124 21.0 131.0 152.2 499.1 499.1 0.00% a6-60 236 7.2 1791.4 1800.0 632.1 624.9 1.14% 210 8.3 1791.2 1800.0 632.1 625.7 1.00% 150 52.3 1557.4 1610.2 632.1 632.1 0.00% a7-70 715 10.3 1788.5 1800.0 646.9 619.6 4.23% 687 13.0 1786.1 1800.0 644.6 622.7 3.40% 364 183.5 1615.6 1800.0 639.9 625.9 2.19% a8-80 613 10.1 1788.7 1800.0 753.6 714.7 5.16% 728 14.2 1784.6 1800.0 747.4 718.0 3.93% 389 192.2 1606.8 1800.0 746.7 720.1 3.56% Avg-A 269.1 8.2 813.9 1.50% 274.4 9.5 806.4 1.19% 162.7 69.7 777.4 0.82% b3-36 203 6.6 148.7 155.5 425.6 425.6 0.00% 227 8.3 479.7 488.2 425.6 425.6 0.00% 133 12.7 181.1 193.8 425.6 425.6 0.00% b4-40 518 9.7 1789.8 1800.0 405.9 402.4 0.85% 502 8.8 156.0 165.1 405.9 405.9 0.00% 260 24.2 132.4 156.8 405.9 405.9 0.00% b5-50 1878 20.9 1777.6 1800.0 497.5 480.1 3.51% 1903 15.6 1783.3 1800.0 494.3 488.4 1.20% 865 44.8 432.3 477.6 494.3 494.3 0.00% b6-60 2537 17.4 1779.9 1800.0 603.1 574.4 4.77% 2334 23.5 1774.6 1800.0 597.8 589.9 1.33% 969 127.6 1230.6 1359.3 596.6 596.5 0.00% b7-70 2310 19.5 1777.3 1800.0 610.8 574.4 5.96% 3287 46.4 1750.2 1800.0 625.7 581.8 7.02% 1145 273.7 1525.2 1800.0 609.0 585.3 3.89% Avg-B 1489.2 14.8 1471.1 3.02% 1650.6 20.5 1210.7 1.91% 674.4 96.6 797.5 0.78% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
Table 1: Computational performance of FFF with different fragment sets for MDARP-LPT within 30 minutes (basic instances: $|V|=\lceil n/4\rceil$, $L=4$)
$S_{RF}$ $S_{FF}$ $S_{PFF}$ F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 37 7.5 0.7 8.2 229.2 229.2 0.00% 33 15.1 1.3 16.5 229.2 229.2 0.00% 29 8.0 0.3 8.4 229.2 229.2 0.00% a3-30 36 6.9 0.6 7.6 325.8 325.8 0.00% 36 7.5 0.6 8.2 325.8 325.8 0.00% 34 7.0 2.0 9.2 325.8 325.8 0.00% a3-36 47 9.1 5.9 15.1 397.2 397.2 0.00% 44 6.3 6.5 12.9 397.2 397.2 0.00% 34 15.2 4.0 19.3 397.2 397.2 0.00% a4-48 116 8.3 40.0 48.6 450.8 450.8 0.01% 109 10.0 27.5 37.9 450.8 450.8 0.00% 92 21.1 23.7 45.1 450.8 450.8 0.00% a6-60 175 9.3 766.0 775.9 556.1 556.1 0.00% 155 7.5 490.1 498.0 556.1 556.1 0.00% 113 40.3 396.6 437.4 556.1 556.1 0.00% a7-70 507 12.0 1786.8 1800.0 552.5 549.8 0.48% 482 14.2 861.7 876.9 552.5 552.5 0.00% 250 139.2 504.8 645.1 552.5 552.5 0.00% a8-80 495 16.9 1781.1 1800.0 663.9 641.2 3.41% 586 15.9 1783.0 1800.0 670.7 642.9 4.15% 318 137.9 1661.4 1800.0 661.8 647.6 2.14% Aver-A 201.9 10.0 636.5 0.56% 206.4 10.9 464.3 0.59% 124.3 52.7 423.5 0.31% b3-36 146 8.3 72.9 81.9 389.3 389.3 0.00% 148 9.5 36.9 46.9 389.3 389.3 0.00% 100 15.2 27.6 42.9 389.3 389.3 0.00% b4-40 370 8.7 45.6 54.7 355.7 355.7 0.00% 314 8.6 23.5 32.3 355.7 355.7 0.00% 178 21.0 16.3 37.5 355.7 355.7 0.00% b5-50 1649 16.8 1781.2 1800.0 436.3 427.0 2.12% 1635 20.4 659.2 680.9 436.3 436.3 0.00% 753 45.4 298.9 344.9 436.3 436.3 0.00% b6-60 2269 31.3 1763.9 1800.0 527.9 512.3 2.95% 2041 28.8 1768.7 1800.0 527.7 522.6 0.98% 816 139.1 397.4 537.9 527.7 527.7 0.00% b7-70 1541 18.6 1778.4 1800.0 530.5 509.6 3.93% 2000 32.5 1764.2 1800.0 528.4 519.1 1.76% 747 233.9 1565.0 1800.0 528.4 518.8 1.81% Aver-B 1195.0 16.7 1107.3 1.80% 1227.6 20.0 872.0 0.55% 518.8 90.9 552.6 0.36% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
Table 2: Computational performance of FFF with different fragment sets for MDARP-LPT within 30 minutes (basic instances: $|V|=\lceil n/3\rceil$, $L=4$)

### v3

@@
 The Dial-a-Ride Problem with Limited Pickups per Trip
-Abstract: The Dial-a-Ride Problem (DARP) is an optimization problem that involves determining optimal routes and schedules for several vehicles to pick up and deliver items at minimum cost.
+Abstract
+The Dial-a-Ride Problem (DARP) is an optimization problem that involves determining optimal routes and schedules for several vehicles to pick up and deliver items at minimum cost.
 Motivated by real-world carpooling and crowdshipping scenarios, we introduce an additional constraint imposing a maximum number on the number of pickups per trip.
@@
 We apply a fragment-based method for DARP-LPT, where a fragment is a partial path.
-Specifically, we extend two formulations from Rist & Forbes (2021): the Fragment Flow Formulation (FFF) and the Fragment Assignment Formulation (FAF).
-We establish FFF’s superiority over FAF, both from a theoretical as well as from a computational perspective.
-Furthermore, our results show that FFF and FAF significantly outperform traditional arc-based formulations in terms of solution quality and time.
-Additionally, compared to the two existing fragment sets, one with longer partial paths and another with shorter ones, our newly generated fragment sets perform better in terms of solution quality and time when fed into FFF.
-Keywords: routing, dial-a-ride problem, limited number of pickups per trip, fragments
+Specifically, we extend two formulations from Rist & Forbes [18]: the Fragment Flow Formulation (FFF) and the Pickup-Space Fragment Formulation (PSFF).
+Furthermore, our results show that PSFF outperforms FFF, which in turn surpasses traditional arc-based formulations in both solution quality and computational efficiency.
+Additionally, we compare several existing fragment sets that differ in the length of their partial paths and find that the sets with shorter partial paths yield the best solution times when used with PSFF.
+In addition, we propose a new mixed fragment set, which is useful when the sets with longer partial paths become too large.
+In such cases, it yields the lowest CPU time.
+keywords:
+routing, dial-a-ride problem, limited number of pickups per trip, fragments
+journal: European Journal of Operational Research
 1 Background
@@
 Each item has a specific pickup and delivery location and time window, and the objective is to schedule the vehicles in a way that minimizes the total travel time or distance, while satisfying various operational constraints such as vehicle capacity, time windows, and pickup and delivery locations.
-DARP has found extensive applications in the sharing mobility (Mourad et al., 2019; Tafreshian et al., 2020) and food delivery sectors.
+DARP has found extensive applications in the sharing mobility [15, 25] and food delivery sectors.
 Within the sharing mobility domain, the ride-sharing routing problem represents a variant of DARP.
-This problem focuses on optimizing routes for drivers, allowing them to efficiently reach their destinations while accommodating passenger pickups and drop-offs, all within the minimum possible time or cost (Dessouky & Mahtab, 2022).
-Similarly, in the context of online food delivery, the central system needs to make decisions on how to serve submitted requests by efficiently picking up and delivering orders with the goal of minimizing delivery time, which can be seen as an online variant of DARP (Guo et al., 2022).
+This problem focuses on optimizing routes for drivers, allowing them to efficiently reach their destinations while accommodating passenger pickups and drop-offs, all within the minimum possible time or cost [9].
+Similarly, in the context of online food delivery, the central system needs to make decisions on how to serve submitted requests by efficiently picking up and delivering orders with the goal of minimizing delivery time, which can be seen as an online variant of DARP [14].
 A frequently disregarded yet vital aspect in this type of routing problems is the limit on the number of pickups per trip.
@@
 In food delivery platforms like Meituan [^4] https://finance.sina.com.cn/tech/2022-05-10/doc-imcwipii9054944.shtml, drivers possess the autonomy to determine the highest number of orders they undertake at one time.
-Similarly, in the crowdshipping domain, considering drivers’ preferences is crucial as they have the right to set their preferred maximum number for parcel pickups (Arslan et al., 2019).
+Similarly, in the crowdshipping domain, considering drivers’ preferences is crucial as they have the right to set their preferred maximum number for parcel pickups [3].
 In particular, a company should not anticipate an occasional driver, who might only intend to deliver a few parcels on a sporadic basis or is simply looking to earn some extra income, to handle a high volume of parcels like a full-time staff driver.
 A similar assumption is found in the concept of limited delivery stops as discussed in Simoni et al.
-(2020) and Fessler et al.
-(2022).
+[22] and Fessler et al.
+[10].
 In this paper, a “customer” can entail more than one person and/or item, but gives rise to only one pickup and two stops (one for pickup and one for delivery).
@@
 Since MDARP-LPT and DARP-LPT share the same problem structure, we collectively refer to them as (M)DARP-LPT.
-To efficiently address (M)DARP-LPT, we propose a unified method with two innovative components: fragment-based formulations and novel fragment generation procedures.
-Firstly, we tailor the fragment-based formulation of DARP proposed by Rist & Forbes (2021) to accommodate (M)DARP-LPT by modifying the decision variables from the fragment-based format to the fragment-vehicle-based format.
-This change is also necessary for DARP-LPT, where all homogeneous vehicles start and end at the same depot.
-Secondly, considering the limited number of pickups, we introduce two novel fragment generation procedures.
-These procedures create improved fragment sets, enhancing computational efficiency.
-Traditionally, there are two fragment sets: one with longer partial paths and the other with shorter ones.
-Our two proposed sets diverge: the first refines the fragment set with longer partial paths through path enumeration, while the second blends longer and shorter partial paths from both traditional fragment sets.
+To efficiently address (M)DARP-LPT, we propose a unified approach featuring two distinct fragment-based formulations and evaluate different types of fragment sets as input.
+Both formulations are adapted from the DARP framework of Rist & Forbes [18].
+The first converts fragment-based decision variables into a fragment–vehicle–based form, a necessary adjustment even for (single-depot) DARP-LPT.
+The second introduces a novel pickup-space fragment formulation, extending the network with a pickup count dimension, similar to time in the time-space formulation, and which uses pickup-space decision variables.
+As for fragment sets: traditional methodologies employ several distinct sets, each with fragments (partial paths) of different levels of length.
+We also introduce a novel mixed fragment set, combining fragments of different lengths to reduce the total number while preserving useful ones as much as possible.
 From a practical perspective, this research addresses a commonly overlooked aspect in the dial-a-ride problem: the limited number of pickups per trip.
 Our developed method offers more tailored and efficient solutions for the problem, thereby contributing to the improvement of carpooling and crowdshipping services.
-From an academic perspective, this study compares and evaluates various fragment sets for the fragment-based formulation in routing problems, providing essential insights for tackling similar pickup and delivery challenges.
-Furthermore, our innovative fragment generation method accelerates computations through enumeration; this logic can offer methodological insights for addressing similar routing problems.
+From an academic perspective, this study proposes novel extended fragment-based formulations and conducts a comprehensive comparison and evaluation of multiple fragment sets for routing problems.
+Our findings offer valuable insights for addressing various problems that incorporate pickup and delivery.
+Additionally, this study is the first to explore the combination of different fragment types for constructing the fragment set used in the formulations, providing valuable insights for routing problems.
 The remainder of this article is organized as follows: Section 2 provides a comprehensive literature review.
-Following that, Section 3 formally defines (M)DARP-LPT, presenting an arc-based formulation and two fragment-based formulations.
-Section 4 introduces traditional and two novel methods for fragment generation.
+Following that, Section 3 formally defines (M)DARP-LPT, presenting an arc-based formulation, two fragment-based formulations, and a path-based formulation.
+Section 4 introduces traditional and novel methods for fragment generation.
 Then, in Section 5, we analyze the results of our computational experiments.
@@
 Since (M)DARP-LPT is a variant of DARP, we investigate existing research methodologies on related optimization problems, such as the Pickup and Delivery Problem (PDP) and the Dial-a-Ride Problem.
-PDP involves transporting a set of items from their pickup locations to their respective delivery locations (where multiple items from a single pickup node may be destined for different delivery nodes), with the objective of finding optimal routes while minimizing transportation costs (Savelsbergh & Sol, 1995).
+PDP involves transporting a set of items from their pickup locations to their respective delivery locations (where multiple items from a single pickup node may be destined for different delivery nodes), with the objective of finding optimal routes while minimizing transportation costs [21].
 DARP is a special case of PDP where each location corresponds to a single item, and every item is associated with a unique pickup and delivery location pair.
-Traditional methods, like branch and cut (Cordeau, 2006) and branch and price (Ropke & Cordeau, 2009; Baldacci et al., 2011), have been utilized for PDP and DARP, but recent studies suggest that fragment-based methods exhibit better performance.
+Traditional methods, like branch and cut [6] and branch and price [19, 4], have been utilized for PDP and DARP, but recent studies suggest that fragment-based methods exhibit better performance.
 Here, a fragment is a partial path where only the start and end node have an empty load.
 A notable example of the fragment-based approach is the work of Alyasiry et al.
-(2019), who apply a fragment-based strategy to address the Pickup and Delivery Problem with Time Windows, following the Last-In-First-Out loading strategy.
-They employ the symbol sequence ``$ppppdddd"$ to denote a full fragment, with “$p$” representing a pickup node and “$d$” indicating a delivery node (this representation is also adopted below).
-This notation implies that, within the fragment, the last pickup must precede the first delivery.
-Alyasiry et al.
-(2019) enumerate all possible paths by combining fragments, thus leading to a fragment-based formulation on the time-space network.
-The fragments’ applicability extends to various related problems, retaining the characteristic of empty-loaded start and end nodes but not limited to the specific “$ppppdddd$” structure.
-Su et al.
-(2023) effectively compute the minimum excess user ride time in an electric autonomous DARP using the fragment-based method.
-Additionally, Zhang et al.
-(2023) tackle a vehicle-customer coordination problem that shares similarities with DARP but incorporates the flexibility for passengers to move in order to meet their drivers faster.
-To tackle this problem, they design an algorithm based on the fragment-based time-space network.
+[2], who apply a fragment-based strategy to address the Pickup and Delivery Problem with Time Windows, following the Last-In-First-Out loading strategy.
+Examples can be represented as $p1p2p3d3d2d1$ or $p1p2d2p3d3d1$, where “$p$” denotes a pickup node, “$d$” denotes a delivery node, and the number indicates the customer index.
+This representation will also be used throughout the remainder of the paper.
+By combining fragments, Alyasiry et al.
+[2] are able to enumerate all feasible paths; based on this, they construct a fragment-based formulation on a time-space network.
+The fragments’ applicability extends to various related problems, retaining the characteristic of empty-loaded start and end nodes but not limited to the specific Last-In-First-Out setting.
+Subsequently, these fragments are referred to as “full fragments”.
+In this context, Su et al.
+[24] effectively compute the minimum excess user ride time in an electric autonomous DARP using the fragment-based method.
+Building on Alyasiry et al.
+[2], Rist & Forbes [18] propose the use of restricted fragments to address the dial-a-ride problem.
+Since the full fragments in this problem don’t necessarily adhere to the Last-In-First-Out property, their number would be quite high.
+However, Rist & Forbes find that a full fragment with, for instance, the form “$ppdpdpdd$” can be decomposed into the restricted fragments “$ppd$” +“$pd$”+ “$pdd$”, which start at a pickup node whose predecessor is a delivery node (or depot) and ends at a delivery node whose successor is a pickup node (or depot).
+Note that “$ppdpdpdd$” is merely an illustration where the last pickup does not necessarily precede the first delivery, this does not imply a universal pattern.
+Rather, it is intended to contrast with the form “$ppppdddd$”, where the last pickup must precede the first delivery.
+Obtaining these restricted fragments involves initially deriving them from fragments conforming to the form “$ppppdddd$” and then applying a dominance rule to minimize their count.
+Unlike Alyasiry et al.
+[2], Rist & Forbes [18] formulate the model on a physical network instead of a time-space network.
+Time-related constraints are handled purely via callbacks, yet the computational burden remains modest.
+Subsequently, Rist & Forbes [17] introduced the concept of extended full fragments for the Selective Dial-a-Ride Problem (SDARP), a variant of DARP that aims to maximize passenger service.
+An extended full fragment is defined as the concatenation of a full fragment followed by an empty arc.
+Notably, the instances used in SDARP feature stricter time windows and shorter ride times than those in standard DARP.
+These tighter constraints motivate the use of a time-space network and help ensure a manageable number of extended full fragments.
+Rist & Forbes’s [2022] computational results demonstrate that extended full fragments outperform existing fragment types for such tightly constrained instances.
+Similarly, Zhang et al.
+[26] adopt a similar definition of extended full fragments, where each fragment is formed by concatenating an empty arc followed by a full fragment.
+They tackle a vehicle-customer coordination problem that shares similarities with DARP but incorporates the flexibility for passengers to move in order to meet their drivers faster.
+In line with Rist & Forbes [17], the test instances in Zhang et al.
+[26] feature even tighter time windows and ride time constraints (e.g., 1.2 to 1.5 times the delivery time limit of an ideal detour-free trip), which further contributes to keeping the number of fragments manageable.
+To tackle this problem, they design an algorithm based on a time-space fragment-based network.
 Their procedure not only achieves effective results, but also demonstrates its value for online DARP.
-Building on Alyasiry et al.
-(2019), Rist & Forbes (2021) propose the use of restricted fragments to address the dial-a-ride problem.
-Since the full fragments in this problem may have the form “$ppdpdpdd$” (as an illustration that the last pickup does not precede the first delivery, this does not imply a universal pattern), their number would be quite high.
-Nevertheless, Rist & Forbes find that a full fragment with the form “$ppdpdpdd$” can be decomposed into the restricted fragments “$ppd$” +“$pd$”+ “$pdd$”, which start at a pickup node whose predecessor is a delivery node (or depot) and ends at a delivery node whose successor is a pickup node (or depot).
-Obtaining these restricted fragments involves initially deriving them from fragments conforming to the form “$ppppdddd$” (where the last pickup must precede the first delivery) and then applying a dominance rule to minimize their count.
-By incorporating temporal information into these restricted fragments, they enable a formulation based on a physical network instead of the time-space network.
-While recent literature has made significant contributions to fragment-based methods, other methods may be further improved for (M)DARP-LPT.
-In particular, alternative methods for fragment generation may result in a higher-quality set of fragments, further enhancing computational performance.
-Moreover, to the best of our knowledge, no study has conducted a comparison of various fragment types or explored the possibility of combining different fragment categories.
+While recent literature has made significant contributions to fragment-based methods, none can be directly applied to handle the limited number of pickups in our problem statement.
+Additionally, different fragment types may result in varying computational performance.
+To the best of our knowledge, no existing study has systematically compared different fragment types or explored the potential of combining multiple fragment categories.
+Another recent development is the event-based formulation for DARP [12, 11], where an event represents a vehicle’s state as defined by its location and onboard passengers.
+The event-based formulation is quite similar to the traditional arc-based formulation of Cordeau [6], but in the event-based model the graph is constructed with events as nodes and feasible transitions as arcs.
+This design implicitly enforces capacity, precedence, and pairing constraints, thereby reducing the number of explicit constraints and improving computational efficiency compared to traditional arc-based models.
+Nevertheless, events inherently provide less routing information than fragments, as noted by Gaul et al.
+[11].
+Given the fairly superior performance of the fragment-based method on the benchmark instances of [18], we only concentrate on fragment-based strategies in this study.
 3 Problem statement and linear formulations
 Both MDARP-LPT and DARP-LPT are extensions of DARP, and their major difference from DARP lies in the limitations on the number of pickups per trip.
-In this section, we begin by providing a formal problem description of DARP-LPT and a NP-hardness proof in Section 3.1.
-Subsequently, we present several linear formulations for DARP-LPT: an arc-based formulation in Section 3.2 and two fragment-based formulations in Section 3.3.
-The arc-based formulation closely resembles the one outlined in Cordeau (2006), except for the constraints on the limited number of pickups.
-The two fragment-based formulations expand on the framework of Rist & Forbes (2021), transitioning from a formulation using fragment-based decision variables to one employing fragment-vehicle-based variables.
+In this section, we begin by providing a formal problem description of DARP-LPT and an NP-hardness proof in Section 3.1.
+Subsequently, we present several linear formulations for DARP-LPT: an arc-based formulation in Section 3.2, two fragment-based formulations in Section 3.3, and a path-based formulation in Section 3.4.
+The arc-based formulation closely resembles the one outlined in Cordeau [6], except for the constraints on the limited number of pickups.
+The two fragment-based formulations build upon the framework of Rist & Forbes [18].
+The first formulation extends the framework by transitioning from fragment-based decision variables to fragment–vehicle-based ones.
+The second extends the network into a pickup–space representation, introducing pickup–space fragment-based variables to account for pickup constraints.
 This adaptation is tailored to our specific requirements for limited number of pickups.
+The path-based formulation presented in Section 3.4 adopts a set partitioning structure, as in Ropke & Cordeau [19].
 Finally, in Section 3.5, we introduce MDARP-LPT, illustrating the differences in its definition and formulations as compared to DARP-LPT.
@@
 CVRP involves arranging a route plan using a fleet of vehicles to visit a set of customers at minimum cost.
-Each vehicle has a capacity denoted by $C$ (Cornuejols & Harche, 1993; Achuthan et al., 2003).
+Each vehicle has a capacity denoted by $C$ [7, 1].
 To perform the reduction, in a DARP-LPT instance, we set the maximum number of pickups per trip $L$ equal to the vehicle capacity $C$ and merge each customer’s pickup and delivery nodes into a single node.
 This creates a problem instance equivalent to CVRP, where each merged node represents a customer with unit size and each vehicle has the capacity $C$ equal to $L$.
-Since CVRP is known to be NP-hard when $C\geq 3$ (Chen, 2023), this reduction demonstrates that DARP-LPT is also NP-hard.
+Since CVRP is known to be NP-hard when $C\geq 3$ [5], this reduction demonstrates that DARP-LPT is also NP-hard.
 MDARP-LPT can also be proven to be NP-hard using a reduction similar to DARP-LPT.
 3.2 Arc-based formulation of DARP-LPT
-The arc-based formulation (ABF), as presented below, is the same as that of Cordeau (2006), except for the constraints (13).
+The arc-based formulation (ABF), as presented below, is the same as that of Cordeau [6], except for the constraints (13).
 The decision variables are as follows:
@@
 $s.t.$ $\sum_{v\in V}\sum_{j\in N}f_{ijv}$ $=1$ $i\in P$ (2)
-$\sum_{j\in N}f_{ijv}$ $=\sum_{j\in N}f_{(i+n)jv}$ $i\in P,v\in V$ (3)
+$\sum_{j\in N}f_{ijv}$ $=\sum_{j\in N}f_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i+n,j,v}}$ $i\in P,v\in V$ (3)
 $\sum_{j\in N}f_{0jv}$ $=1$ $v\in V$ (4)
 $\sum_{j\in N}f_{jiv}$ $=\sum_{j\in N}f_{ijv}$ $i\in P\cup D,v\in V$ (5)
-$\sum_{i\in N}f_{i(2n+1)v}$ $=1$ $v\in V$ (6)
+$\sum_{i\in N}f_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,2n+1,v}}$ $=1$ $v\in V$ (6)
 $t_{iv}+T_{ij}$ $\leq t_{jv}+M_{ij}(1-f_{ijv})$ $(i,j)\in A,v\in V$ (7)
@@
 $\max(0,q_{i})\leq Q_{iv}$ $\leq\min(Q,Q+q_{i})$ $i\in N,v\in V$ (10)
-$t_{(n+i)v}-t_{iv}$ $\leq R_{i}$ $i\in P,v\in V$ (11)
-$t_{(2n+1)v}-t_{0v}$ $\leq R_{0}$ $v\in V$ (12)
+$t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n+i,v}}-t_{iv}$ $\leq R_{i}$ $i\in P,v\in V$ (11)
+$t_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2n+1,v}}-t_{0v}$ $\leq R_{0}$ $v\in V$ (12)
 $\sum_{(i,j)\in A}f_{ijv}$ $\leq 2L+1$ $v\in V$ (13)
@@
 This definition arises from the observation that, if a vehicle can deliver at most $L$ customers, a minimum of $\lceil n/L\rceil$ vehicles are needed.
-Therefore, we add $f_{0(2n+1)v}=0$ for each fixed vehicle $v\in V_{f}$.
+Therefore, we add $f_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0,2n+1,v}}=0$ for each fixed vehicle $v\in V_{f}$.
 3.3 Fragment-based formulations of DARP-LPT
-The framework presented by Rist & Forbes (2021) defines a fragment-based formulation for DARP, which necessitates the inclusion of callbacks to handle rarely violated constraints related to subtour elimination, time windows, and maximum ride time.
+The framework presented by Rist & Forbes [18] defines a fragment-based formulation for DARP, which necessitates the inclusion of callbacks to handle rarely violated constraints related to subtour elimination, time windows, and maximum ride time.
 In the case of DARP, even when multiple vehicles are used, this fragment-based format remains feasible.
-In this section, our fragment-based formulations build upon the work of Rist & Forbes (2021).
-However, since DARP-LPT involves frequently violated constraints, in particular the limitations on the number of pickups per trip (per vehicle), we introduce different decision variables, shifting from a fragment-based format to a fragment-vehicle-based format.
-Consequently, we propose two distinct fragment-based formulations tailored for multi-vehicle settings: a fragment flow formulation and a fragment assignment formulation.
-The former assigns a vehicle to each fragment and node arc, while the latter exclusively links a vehicle with each fragment.
-In particular, the latter formulation draws inspiration from the client-depot assignment formulation in multi-depot routing problems (Bektaş et al., 2020).
-In our context, the “client” corresponds to the fragment, while the “depot” corresponds to the vehicle.
+This section builds upon the fragment-based formulations introduced by Rist & Forbes [18].
+We propose two extended formulations to suitably manage the frequent violation of specific constraints in DARP-LPT, particularly the limited number of pickups per trip.
+Formulation FFF augments the decision variables by incorporating vehicle indices, while PSFF reformulates the network using a pickup-space structure and omits vehicle details.
+In the case of MDARP-LPT, FFF retains its structure, whereas PSFF must reintroduce vehicle indices to account for the unique origin and destination associated with each vehicle.
 3.3.1 Preliminaries and notation
-In this subsection, we largely follow the definitions of Rist & Forbes (2021).
+In this subsection, we largely follow the definitions of Rist & Forbes [18].
 A DARP route is a route $(0,i_{1},i_{2},\ldots,i_{K},2n+1)$ for which at least one feasible schedule exists, which respects the pairing, precedence, capacity, and time window constraints of DARP.
@@
 Any segment of this sequence is referred to as a route path.
-Rist & Forbes (2021) introduce the concept of restricted fragments for DARP.
+Rist & Forbes [18] introduce the concept of restricted fragments for DARP.
 These fragments can be combined to enumerate all DARP routes, and a network is constructed to include all these restricted fragments.
@@
 A restricted fragment (RF) is a part of a full fragment, starting at a pickup node whose predecessor is a delivery node or depot and ending at a delivery node whose successor is a pickup node or depot.
-An RF contains at least one pickup-to-delivery movement (not necessarily corresponding to a customer’s pickup and delivery location pair) and cannot be divided into smaller RFs (meaning each RF is as small as possible), e.g., $(p1,p2,d1)$, $(p1,d2,d1)$, $(p1,p2,p3,d4,d5)$, $(p5,d8)$, $(p5,p7,d1,d2,d3)$.
+An RF contains exactly one pickup-to-delivery movement (not necessarily corresponding to a customer’s pickup and delivery location pair), e.g., $(p1,p2,d1)$, $(p1,d2,d1)$, $(p1,p2,p3,d4,d5)$, $(p5,d8)$, $(p5,p7,d1,d2,d3)$.
 The following two examples are not RFs: $(p1,p2,p3,p4)$ and $(p2,d1,p3,d3,d2)$, where the former lacks a pickup-to-delivery movement and the latter can be split after $d1$.
@@
 For instance, the nodes from the route path $(p1,p2,d1,p3,d2,p4,d3,d4)$ are expressed as $p1$, $(p2,\{l_{1}\})$, $(d1,\{l_{2}\})$, $(p3,\{l_{2}\})$, $(d2,\{l_{3}\})$, $(p4,\{l_{3}\})$, $(d3,\{l_{4}\})$, and $d4$.
-Notably, nodes with the same location can be distinct, as illustrated by examples such as $(p2,\{l_{1}\})$, $(p2,\{l_{3}\})$, $(p2,\{l_{2},l_{3}\})$, and $p2$.
+Notably, nodes with the same location can be distinct, as illustrated by examples such as $(p2,\{l_{1}\})$, $(p2,\{l_{3}\})$, $(p2,\{l_{1},l_{3}\})$, and $p2$.
 We define a node with a pickup location as a “pickup node” and a node with a delivery location as a “delivery node”.
@@
 Building upon the introductory concepts, we now introduce the fragment-based network.
-To facilitate comprehension, we present several figures (Fig.
-1, Fig.
-2, and Fig.
-3).
+To facilitate comprehension, we provide Fig.
+1 and Fig.
+2.
+Fig.
+1 illustrates several DARP routes alongside their corresponding fragments, whereas Fig.
+2 depicts the corresponding fragment-based network.
+This network can be derived from Fig.
+1 by first merging all identical fragments (with an intermediate transformation shown in Fig.
+A.1 in A), and then merging identical nodes while preserving the original connections.
 In these figures, ellipses represent nodes, solid lines (with ellipses on both sides) indicate fragments (with their route paths written above or to the right), and dashed lines with arrows are node arcs (further explained below).
@@
 Figure 1: Several feasible routes for DARP and their decomposition into fragments (RF and FF)
+[…]
+Figure 2: Fragment-based network for DARP by merging all identical fragments and nodes from Fig. 1
 Before explaining the fragment-based network, we will introduce the connections between fragments and the connections between nodes (referred to later as node arcs), both of which can be seen as the dashed lines in Fig.
@@
 Crucially, the connected nodes must share the same load.
-Decomposing feasible DARP routes into fragments yields a significantly smaller number of fragments compared to DARP routes.
-Conversely, if we combine these fragments, we can obtain all feasible DARP routes, albeit with the possibility of some redundant routes.
-Fig.
-2 illustrates how feasible DARP routes can be gathered in a transitional network (as an intermediate step towards the fragment-based network).
-Fig.
-1 presents several DARP routes and their corresponding fragments.
-The network in Fig.
-2 will use these fragments and allow to enumerate all the DARP routes shown in Fig.
-1.
-Identical fragments across different routes in Fig.
-1 can be consolidated into a single fragment while maintaining the connections to other fragments.
-Subsequently, the fragment-based network of Fig.
-3 is derived from the network of Fig.
-2.
-When different fragments share the same start and/or end nodes, then those nodes can also be merged.
-For instance, two fragments with the same start and end nodes but different route paths $(p1,d1)$ and $(p1,p4,d4,d1)$ become two route paths originating from node $p1$ and terminating at node $d1$.
-Such consolidations help reduce connections (node arcs).
-Node arcs with the same start node and end node can also be merged; the three connections from $(p1,d1)$ to $(p2,d2)$, from $(p1,d1)$ to $(p2,p3,d2,d3)$, and from $(p1,p4,d4,d1)$ to $(p2,d2)$ in Fig.
-2, for instance, can be merged into one node arc $(d1,p2)$ in Fig.
-3.
-$O^{+}$$p3$($d2$,{$l_{3}$})($p1$,{$l_{3}$})$d3$$O^{-}$($p3$,$p2$,$d2$)($p1$,$d1$,$d3$)$p1$$d1$$p2$$d2$$p3$$d3$($p1$,$d1$)($p2$,$d2$)($p3$,$d3$)$p1$$d1$$p2$$d3$($p1$,$p4$,$d4$,$d1$)($p2$,$p3$,$d2$,$d3$)
-Figure 2: A transitional network for DARP by merging all identical fragments from Fig. 1
-[…]
-Figure 3: Fragment-based network for DARP by merging all identical nodes from Fig. 2
 The practical process of network generation proceeds as follows.
@@
 2 must adhere to the alternating pattern of dashed lines and solid lines.
+For details on why Fig.
+2 can be used to enumerate all paths, please refer to A.
 We now formalize the foregoing concepts.
@@
 3.3.3 Fragment flow formulation
-The first fragment-based formulation is the fragment flow formulation (FFF), which expands on the framework of Rist & Forbes (2021).
+The first fragment-based formulation is the fragment flow formulation (FFF), which expands on the framework of Rist & Forbes [18].
 It transitions from a formulation using fragment-based and node arc-based decision variables to one employing fragment-vehicle-based and node arc-vehicle-based variables.
@@
 $\sum_{a\in A_{O^{+}}^{+}}y_{av}$ $=w_{v}$ $v\in V$ (23)
-3.3.4 Fragment assignment formulation
-Table 3: Decision variables of FAF
+Furthermore, to enhance the formulation FFF, we introduce additional constraints based on the minimum number of used vehicles, as shown in constraints (24).
+While these constraints may be redundant, they significantly speed up the computation process.
+Additionally, for constraints (23), we set $w_{v}$ equal to 1 for $v\in V_{f}$.
+$\sum_{f\in F}x_{vf}$ $\geq 1$ $v\in V_{f}$ (24)
+3.3.4 Pickup-space fragment formulation
+We now describe the Pickup-Space Fragment Formulation (PSFF), which extends the fragment-based formulation of Rist & Forbes [18] by introducing the number of pickups per trip as a new dimension, analogous to a time-space representation.
+Unlike FFF, which uses a separate constraint for pickup limits, PSFF integrates this constraint directly into the pickup-space network.
+The pickup-space network is formed by replicating the fragment-based network across $L+1$ levels, where each level represents the remaining number of allowable pickups throughout a trip.
+The index sets $\mathcal{L_{P}}=\{1,\dots,L\}$ and $\mathcal{L_{D}}=\{0,\dots,L-1\}$ define the remaining pickup levels for pickup and delivery nodes, respectively.
+When a vehicle traverses a fragment with $l\in\mathcal{L_{P}}$ remaining allowable pickups before entry, it concludes at level $l-L_{f}\in\mathcal{L_{D}}$.
+The transitions for node arcs adhere to the same level structure, ensuring consistency in tracking the remaining allowable number of pickups.
+The following formulation does not include all constraints. Lazy constraints are incorporated to address subtour elimination, time windows, and ride-time constraints, as detailed in Section 3.3.5.
+Table 3: Decision variables of LSF
 Decision Variables Definition
-$X_{f}$ = 1, if fragment $f\in F$ is traversed; = 0, otherwise
-$Y_{a}$ = 1, if node arc $a\in A_{N}$ is traversed; = 0, otherwise
-$x_{fv}$ = 1, if fragment $f\in F$ is traversed by vehicle $v\in V$; = 0, otherwise
-An alternative formulation involves the assignment of fragments to vehicles and is called fragment assignment formulation (FAF).
-FAF extends the framework of Rist & Forbes (2021) by transitioning from a formulation using fragment-based decision variables to one that incorporates fragment-vehicle-based variables.
-Unlike FFF, however, which introduces node arc-vehicle-based decision variables, FAF keeps the node arc-based variables by means of a separate set of constraints (which hinges upon constraint set (31)).
-This model is inspired by the client-depot assignment formulation used in multi-depot routing problems (Bektaş et al., 2020).
-In the same way as for FFF, the current formulation below is not complete.
-Lazy constraints to address subtour elimination, time windows, and maximum ride time are discussed in Section 3.3.6.
-$\min$ $\sum_{f\in F}c_{f}X_{f}+\sum_{a\in A_{N}}c_{a}Y_{a}$ (24)
-$s.t.$ $\sum_{f\in F_{h}^{-}}X_{f}$ $=\sum_{a\in A_{h}^{+}}Y_{a}$ $h\in P_{N}$ (25)
-$\sum_{f\in F_{h}^{+}}X_{f}$ $=\sum_{a\in A_{h}^{-}}Y_{a}$ $h\in D_{N}$ (26)
-$\sum_{f\in F_{i}}X_{f}$ $=1$ $i\in P$ (27)
-$\sum_{a\in A_{O^{+}}^{+}}Y_{a}$ $\leq|V|$ (28)
-$x_{fv}$ $\leq X_{f}$ $f\in F,v\in V$ (29)
-$\sum_{v\in V}x_{fv}$ $=X_{f}$ $f\in F$ (30)
-$\sum_{f\in F^{-}_{i}}x_{fv}+Y_{a}$ $\leq\sum_{f\in F^{+}_{j}}x_{fv}+1$ $a=(i,j)\in A_{N}:i\neq O^{+},j\neq O^{-},v\in V$ (31)
-$\sum_{f\in F}L_{f}x_{fv}$ $\leq L$ $v\in V$ (32)
-$X_{f},Y_{a}$$\in\{0,1\}$ $f{\in}F,a{\in}A_{N}$ (33)
-$x_{fv}$$\in\{0,1\}$ $f{\in}F,v{\in}V$ (34)
-The objective function (24) minimizes the weighted cost of used fragments and node arcs.
-Constraints (25) and (26) describe the network flow for fragments and node arcs.
-Constraints (27) guarantee that each pickup location is visited exactly once.
-Constraints (28) state that at most $|V|$ vehicles are used.
-Constraints (29)-(31) assign the fragments to vehicles.
-Specifically, constraints (29) impose that if a fragment is traversed by a vehicle, the fragment is used; constraints (30) state that if a fragment is chosen, it must be assigned to a vehicle; and constraints (31) express that if a node arc is used, the fragments preceding the arc and the fragments succeeding the arc must be traversed by the same vehicle.
-This explains why it does not matter which vehicle traverses the node arc.
-Constraints (32) set the maximal number of pickups.
-Constraints (33) and constraints (34) specify the domains of $X$, $Y$, and $x$.
-Although the variables $X$ and $x$ in this formulation can be expressed solely in terms of $x$, keeping both $X$ and $x$ yields better computational results.
-3.3.5 Comparison of FFF and FAF
-In this subsection, we will establish that FFF is a stronger formulation than FAF.
-We will show this in two steps.
-Firstly, any feasible solution to the linear relaxation of FFF also qualifies as a feasible solution to the linear relaxation of FAF.
-Secondly, we will provide an example solution from the polytope of FAF that does not appear in the polytope of FFF.
-Regarding the first step, it is easy to observe that constraints (18)-(20) and constraints (21) in FFF can imply constraints (25)-(27) and constraints (32) in FAF.
-Furthermore, constraints (23) in FFF can indicate constraints (28) in FAF.
-Therefore, our only task is to demonstrate how a solution ($x^{\prime}$, $y^{\prime}$) that is feasible under FFF constraints (18) and (19) can meet the requirements of FAF constraints (31).
-Let ($x^{\prime}$, $y^{\prime}$) represent a solution from the linear relaxation of FFF.
-We introduce two auxiliary variables: $X^{\prime}$ and $Y^{\prime}$.
-Here, $X^{\prime}_{f}=\sum_{v\in V}x_{fv}$ for all $f\in F$ and $v\in V$; and $Y^{\prime}_{f}=\sum_{v\in V}y^{\prime}_{av}$ for all $a\in A_{N}$ and $v\in V$.
-Proposition 1.
-Constraints (18) and constraints (19) in FFF imply constraints (31) in FAF.
-Proof.
-Consider an arc $a=(i,j)$.
-Our objective is to demonstrate that any solution $(x^{\prime},y^{\prime},X^{\prime},Y^{\prime})$ satisfying the FFF constraints will also be feasible for the FAF network flow constraints (31).
-According to the FFF network flow constraints (18) and (19), we have:
-$\sum_{f\in F_{i}^{-}}x^{\prime}_{fv}=\sum_{a\in A_{i}^{+}}y^{\prime}_{av},\quad\text{and}\quad\sum_{f\in F_{j}^{+}}x^{\prime}_{fv}=\sum_{a\in A_{j}^{-}}y^{\prime}_{av}$
-Thus, we can transform the FAF constraints (31) as follows:
-$\sum_{f\in F^{-}_{i}}x^{\prime}_{fv}+Y^{\prime}_{a}\leq\sum_{f\in F^{+}_{j}}x^{\prime}_{fv}+1\iff\sum_{a^{\prime}\in A_{i}^{+}}y^{\prime}_{a^{\prime}v}+Y^{\prime}_{a}\leq\sum_{a^{\prime}\in A_{j}^{-}}y^{\prime}_{a^{\prime}v}+1\iff\sum_{a^{\prime}\in A_{i}^{+}\setminus a}y^{\prime}_{a^{\prime}v}+Y^{\prime}_{a}\leq\sum_{a^{\prime}\in A_{j}^{-}\setminus a}y^{\prime}_{a^{\prime}v}+1$
-We now examine the left side of the last inequality, $\sum_{a^{\prime}\in A_{i}^{+}\setminus a}y^{\prime}_{a^{\prime}v}+Y^{\prime}_{a}$, using FFF constraints (18). Then, we obtain:
-$\sum_{a^{\prime}\in A_{i}^{+}\setminus a}y^{\prime}_{a^{\prime}v}+Y^{\prime}_{a}\leq\sum_{a\in A_{i}^{+}}Y_{a}=\sum_{f\in F^{-}_{i}}X_{f}\leq 1\leq\sum_{a^{\prime}\in A_{j}^{-}\setminus a}y^{\prime}_{a^{\prime}v}+1$
-Thus, we see that constraints (31) are automatically fulfilled if constraints (18) and (19) hold. ∎
-The second step is to show that the polyhedra of the two formulations do not coincide.
-To illustrate this point, we provide an example solution of FAF that is cut away by FFF.
-For example: if $V=\{v1,v2\}$, $F^{-}_{i}=\{f\}$, $F^{+}_{j}=\{g\}$, and $a=(i,j)$, constraints (31) of FAF become: $x_{fv1}+x_{fv2}+Y_{a}\leq x_{gv1}+x_{gv2}+1$.
-A potential solution under the given constraints is $x_{fv1}=0.5$, $x_{fv2}=0.5$, $Y_{a}=1$, $x_{gv1}=1$, $x_{gv2}=0$.
-However, this solution does not satisfy constraints (18) and (19) of FFF, where $x_{fv2}=y_{av2}$ and $y_{av2}=x_{gv2}$.
-Based on the foregoing, we conclude that FFF dominates FAF.
-Despite this dominance result, we should still verify empirically how the two formulations behave in terms of runtimes.
-FAF, for instance, has less decision variables since FAF utilizes arc-based variables $Y$, whereas FFF uses arc-vehicle-based variables $y$.
+$X_{fl}$ = 1, if a vehicle traverses fragment $f\in F$ with $l$ remaining allowable pickups
+before entry; = 0, otherwise
+$Y_{al}$ = 1, if a vehicle traverses node arc $a\in A_{N}$ while having $l$ remaining allowable
+pickups available; = 0, otherwise
+$\min\sum_{l\in\mathcal{L_{P}}}\sum_{f\in F}c_{f}X_{fl}+\sum_{l\in\mathcal{L_{D}}}\sum_{a\in A_{N}}c_{a}Y_{al}$ (25)
+$s.t.\quad$ $\sum_{f\in F_{h}^{-}}X_{f,l+L_{f}}$ $=\sum_{a\in A_{h}^{+}}Y_{al}$ $h\in P_{N},\;l\in\mathcal{L_{P}}$ (26)
+$\sum_{f\in F_{h}^{+}}X_{fl}$ $=\sum_{a\in A_{h}^{-}}Y_{al}$ $h\in D_{N},\;l\in\mathcal{L_{D}}$ (27)
+$\sum_{l\in\mathcal{L_{P}}}\sum_{f\in F_{i}}X_{fl}$ $=1$ $i\in P_{N}$ (28)
+$\sum_{a\in A_{o}^{+}}Y_{aL}$ $\leq|V|$ (29)
+$X_{fl}$$\in\{0,1\}$ $f\in F\;l\in\mathcal{L_{P}}$ (30)
+$Y_{al}$$\in\{0,1\}$ $a\in A_{N}\;l\in\mathcal{L_{D}}$ (31)
+The objective 25 is to minimize the total weighted cost of fragments and arcs.
+Constraints (26) and (27) establish the network flow constraints on each pickup and delivery node at every remaining pickup level.
+These constraints help track the remaining allowable pickup number throughout the trip.
+Specifically, constraints (26) ensure that if a fragment $f$ initially has a remaining allowable pickup number of $l+L_{f}$ then it will end at a delivery node with level $l$, and then a node arc must depart from this delivery node at the same level.
+Constraints (27) specify that if a fragment departs from a pickup node at level $l$, then a node arc must end at this pickup node at level $l$ (and this node arc also starts at level $l$).
+Constraint (28) ensures that each pickup location is visited exactly once, thereby guaranteeing that all customer requests are fulfilled.
+Constraint (29) imposes an upper bound on the total number of vehicles used in the network.
+In particular, it restricts the number of pickup-space node arcs that depart from the origin depot with a remaining number of pickups equal to $L$.
+Constraints (30) and (31) define the domains of $X$ and $Y$.
 3.3.5 Inequalities for fragment formulations
-The above fragment-based formulations FFF and FAF may not be sufficient to lead to feasible solutions.
+The above fragment-based formulations FFF and PSFF may not be sufficient to lead to feasible solutions.
 Subtour elimination, time windows, and maximum ride time constraints will be implemented by the callbacks of infeasible path constraints, in line with Alyasiry et al.
-(2019); Rist & Forbes (2021).
+[2], Rist & Forbes [18].
 The details are as follows.
@@
 A cycle has the same connection requirement as a chain but includes an additional node arc $a_{c}$ that connects $f_{c}$ back to $f_{1}$.
-If we obtain a chain, we connect its initial and final fragments to the origin and destination depots, respectively, to construct a path.
+For FFF, if we obtain a chain, we connect its initial and final fragments to the origin and destination depots, respectively, to construct a path.
 After that, we establish a valid time schedule for the path, following the constraints of time windows and maximum ride time.
@@
 $\sum_{k=1}^{c}x_{f_{k}v}+\sum_{k=1}^{c}y_{a_{k}v}$ $\leq 2c-1$ $v\in V$ (33)
-In addition to the necessary constraints mentioned earlier, we incorporate three types of inequalities proposed by Rist & Forbes (2021), namely fragment-fragment, arc-fragment, and fragment-arc inequalities, and add the subset-row inequalities proposed by Jepsen et al.
-(2008).
-Fragment-fragment inequalities stipulate that when a fragment is traversed, at least one fragment from its preceding and one fragment from its succeeding fragment set must also be traversed.
-The preceding and succeeding fragment sets include all fragments except the ones that, when combined with the current fragment, would result in infeasible paths.
-Arc-fragment inequalities state that if a node arc is traversed, at least one fragment from its succeeding and one fragment from its preceding fragment set must be traversed to maintain path feasibility.
-Fragment-arc inequalities specify that if a fragment is traversed, at least one succeeding node arc and one preceding node arc in the corresponding sets must also be traversed.
-In addition, for subset-row inequalities, we follow the same approach as Rist & Forbes (2021) and Jepsen et al.
-(2008).
-Furthermore, to enhance the formulations (both FFF and FAF), we introduce additional constraints based on the minimum number of used vehicles, as shown in constraints (37).
-While these constraints may be redundant, they significantly speed up the computation process.
-Additionally, for constraints (23), we set $w_{v}$ equal to 1 for $v\in V_{f}$.
-$\sum_{f\in F}x_{vf}$ $\geq 1$ $v\in V_{f}$ (37)
+For PSFF the procedure remains the same, but for the callback constraints the variables $x$ and $y$ in constraints (32) and (33) are replaced by $\sum_{l\in\mathcal{L_{P}}}X_{f_{k}l}$ and $\sum_{l\in\mathcal{L_{D}}}Y_{a_{k}l}$, respectively, and the vehicle index and vehicle set are no longer required.
+3.4 Path-based formulation
+Since a fragment generally spans a longer sequence than a location arc, its enumeration requires more computational effort.
+To further understand the trade-off, we enumerate all feasible paths and solve the problem using a path-based formulation (PBF).
+Given the limited number of pickups per trip, such enumeration may become practically viable, especially when the maximum number of pickups $L$ is small or when the instance size is relatively small.
+The path-based formulation follows a set partitioning structure, as in Ropke et al.
+[20].
+Let $\omega$ denote the set of all enumerated feasible paths.
+For each path $p\in\omega$, we define $a_{p}^{i}$ as a binary parameter that takes value 1 if location $i\in N$ is visited in path $p$, and 0 otherwise.
+Let $c_{p}$ represent the cost associated with path $p$.
+The decision variable $z_{p}$ equals 1 if path $p$ is selected, and 0 otherwise.
+The formulation is given below.
+$\min\quad$ $\sum_{p\in\omega}c_{p}z_{p}$ (34)
+$s.t.\quad$ $\sum_{p\in\omega}a_{p}^{i}z_{p}=1$ $i\in N$ (35)
+$\sum_{p\in\omega}z_{p}\leq|V|$ (36)
+$z_{p}\in\{0,1\}$ $p\in\omega$ (37)
+The objective (34) minimizes the total cost of the selected paths.
+Constraints (35) ensure that each location is visited exactly once, while constraint (36) specifies the maximum number of vehicles.
+Finally, constraints (37) define the binary domain of the decision variables $z$.
 3.5 MDARP-LPT: problem description and formulations
@@
 In all other aspects, MDARP-LPT adheres to the same principles as DARP-LPT.
-The formulations developed supra (ABF, FFF, and FAF) for DARP-LPT can also be used for MDARP-LPT, with some modifications as explained below.
-For ABF, only the network flow constraints (4)-(6) need to be updated to reflect the unique origin and destination for each vehicle. The complete formulation can be found in A.
-For FFF, all constraints (18)-(22) remain applicable, but the network configuration assigns unique origin and destination nodes, fragment sets, node arc sets, and node sets to each vehicle.
-For FAF, in addition to the network adjustments, two further constraints, constraints (38) and (39), must be incorporated.
-These constraints stipulate that when a node arc starting from an origin is utilized, the subsequent fragments must also be employed, and similarly, when a node arc ending at a destination is employed, the preceding fragments must also be utilized.
-Specifically, for each vehicle, it possesses a unique set of node arcs that originate from its origin and a unique set of node arcs that end at its destination node.
-$Y_{a}$ $\leq\sum_{f\in F^{+}_{j}}x_{fv}$ $a=(O_{v},j)\in A_{N},v\in V$ (38)
-$Y_{a}$ $\leq\sum_{f\in F^{-}_{j}}x_{fv}$ $a=(j,D_{v})\in A_{N},v\in V$ (39)
-Additionally, for both FFF and FAF, the model for MDARP-LPT needs to incorporate infeasible path constraints (constraints (35)) through callback procedures to handle the constraints of subtour elimination, time windows, and maximum ride time.
-Given that MDARP-LPT involves distinct vehicles with unique origins and destinations, constraints such as (37) based on the minimum number of vehicles are not used for MDARP-LPT.
+The formulations developed supra (ABF, FFF, PSFF, and PBF) for DARP-LPT can also be used for MDARP-LPT, with some modifications as explained below.
+For ABF, only the network flow constraints (4)-(6) need to be updated to reflect the unique origin and destination for each vehicle.
+Specifically, the generic locations 00 and $12n+$ 1 are updated to $o_{v}$ and $d_{v}$, respectively.
+For the case of FFF and PSFF, all constraints remain applicable, but the network configuration assigns unique origin and destination nodes, fragment sets, node arc sets, and node sets to each vehicle.
+Finally, for PBF the only modification is to replace the constraint (36) by a separate inequality $\sum_{p\in\omega_{v}}z_{p}\leq 1$ for each vehicle $v$, where $\omega_{v}$ denotes the set of paths associated with vehicle $v$.
 4 Fragment generation
-In this section we describe how the fragments that constitute the basis for the formulations FFF and FAF can be acquired.
-Section 4.1 describes the FF generation procedure of Alyasiry et al.
-(2019) and the RF generation method of Rist & Forbes (2021), both of which involve enumeration.
-The resulting fragment sets for (M)DARP-LPT are denoted as $S_{FF}$ and $S_{RF}$, respectively.
-From these sets, all fragments whose length exceeds $L$ are already eliminated.
-The set $S_{FF}$ with longer fragments tends to be larger than $S_{RF}$ with shorter fragments, yet all combinations of RFs from $S_{RF}$ can generate more (redundant) FFs than those in $S_{FF}$.
-In DARP, solution paths can vary in length, potentially exceeding $L$.
-However, in (M)DARP-LPT, solution paths must not exceed this limit.
-Since $S_{FF}$ and $S_{RF}$ are primarily designed for DARP, some of their fragments (even if not exceeding $L$ in length) might only appear on solution paths that exceed $L$ and not on those that adhere to the limit (because some higher-cost paths are eliminated from consideration; see Section 4.2).
-Therefore, in Section 4.2, we introduce a novel path-enumerated full fragment set $S_{PFF}$ by enumerating all (M)DARP-LPT paths first and then decomposing them into full fragments.
-Consequently, $S_{PFF}$ must be a subset of $S_{FF}$.
-Another novel method in Section 4.3 is to create a mixed fragment set $S_{MF}$ by converting certain FFs from $S_{FF}$ into RFs while keeping others, aiming to leverage the strength of both FFs and RFs and potentially reduce the set size.
-All these fragment sets ($S_{FF}$, $S_{RF}$, $S_{PFF}$, and $S_{MF}$) are valid for both MDARP-LPT and DARP-LPT.
+This section describes how the fragments that constitute the basis for the formulations FFF and PSFF can be acquired.
+Section 4.1 outlines the enumeration-based fragment generation procedures of FFs [2] and RFs [18], resulting in fragment sets $S_{FF}$ and $S_{RF}$, respectively.
+Section 4.2 introduces extended fragments (EFF) and extended restricted fragments (ERF), along with their generation procedures for the corresponding sets $S_{EFF}$ and $S_{ERF}$.
+Finally, Section 4.3 describes a novel mixed fragment set $S_{MF}$, derived by converting FFs from $S_{FF}$ into RFs and ERFs.
+This approach aims to reduce the fragment count while maintaining more useful fragments.
+From these sets, all fragments whose length exceeds $L$ are eliminated.
+The set $S_{FF}$, containing longer fragments, tends to be larger than $S_{RF}$, which contains shorter fragments.
+However, all combinations of RFs from $S_{RF}$ can generate more (redundant) FFs than those present in $S_{FF}$.
+This observation also applies to the comparison between $S_{ERF}$ and $S_{EFF}$.
+All these fragment sets ($S_{FF}$, $S_{RF}$, $S_{EFF}$, $S_{ERF}$, and $S_{MF}$) are valid for both MDARP-LPT and DARP-LPT.
 However, in MDARP-LPT, where vehicles have different origins and destinations, each vehicle requires a unique fragment set (though many of the fragments will still be the same for each vehicle).
@@
 In contrast, DARP-LPT, with a fixed depot for all vehicles, only necessitates a single execution of the generation procedure.
-4.1 Traditional fragment generation
+4.1 RF and FF generation
 Alyasiry et al.
-(2019) introduce FF and $S_{FF}$, while Rist & Forbes (2021) propose RF and $S_{RF}$.
+[2] introduce FFs and $S_{FF}$, while Rist & Forbes [18] propose RFs and $S_{RF}$.
 $S_{FF}$ exclusively contains FFs, while $S_{RF}$ exclusively contains RFs.
-Their intersection includes all FFs with the format “$ppppdddd$”, which can also be regarded as RFs.
+Their intersection is all FFs with the format “$ppppdddd$”, which can also be regarded as RFs.
 As previously explained, “$ppppdddd$” indicates a sequence where all pickups precede all deliveries.
 […]
-4.2 Path-enumerated full fragment generation
-Generating $S_{PFF}$ involves three steps: enumerating all paths that adhere to the problem constraints, selecting enumerated paths, and generating fragments from these selected paths.
-The first step is to enumerate all possible paths.
-Typically, such a task can be challenging due to the exponential growth in the number of paths.
-With limited pickups, however, this number is polynomial for $L$.
-To be more precise, the number of paths is bounded by $\mathcal{O}((2n)^{2L})$.
-The second step involves path selection.
-For any pair of enumerated paths that visit an identical set of locations, only the one with lower cost is kept.
-The third step involves extracting full fragments from the selected paths.
-Within each path, nodes with empty loads are identified and marked.
-The route paths delimited by consecutive markers are then extracted to create fragments, which are subsequently recorded in $S_{PFF}$.
-Notably, $S_{PFF}$ should be a subset of $S_{FF}$, as $S_{PFF}$ is derived from the selected paths, and not all FFs from $S_{FF}$ will appear on these paths.
-Since $S_{PFF}$ includes all FFs obtained from all non-dominated (M)DARP-LPT paths (paths with higher cost are excluded), these FFs can also be re-combined to reconstitute all non-dominated paths.
-We illustrate the foregoing with an example.
-Consider an instance with only two customers, 1 and 2.
-The set $S_{FF}$ includes the following fragments: $(p1,p2,d1,d2)$, $(p1,p2,d2,\\ d1)$, $(p2,p1,d1,d2)$, $(p2,p1,d2,d1)$, $(p1,d1)$, and $(p2,d2)$.
-However, there are at most three useful paths: $(0,p1,d1,2n+1)$, $(0,p2,d2,2n+1)$, and then out of the four paths that include the four locations $p1$, $p2$, $d1$, and $d2$ and serve both customers, only one of the four (a shortest path) needs to be retained.
-As a result, $S_{PFF}$ can contain at most three fragments from the three paths and is thus clearly a subset of $S_{FF}$.
-4.3 Mixed fragment generation
-Path enumeration can become significantly more time-consuming as the limit on the number of pickups $L$ increases.
-Hence, we present an alternative procedure for creating fragments as $L$ grows.
-In this procedure, we initially create all FFs to get $S_{FF}$.
-Next, we employ a selection process to identify specific FFs to be decomposed into RFs.
-Finally, the remaining FFs together with the newly formed RFs constitute the set $S_{MF}$.
-The first step generates FFs adhering to the (M)DARP-LPT constraints, according to the procedure described in Section 4.1, resulting in the same fragment set $S_{FF}$.
-The second step decomposes all FFs into RFs and records the frequency of occurrence of each RF during this process.
-In a FF, pickup nodes whose preceding nodes are depots or delivery nodes, and delivery nodes whose subsequent nodes are depots or pickup nodes, are marked.
-The route paths between consecutive marks can be utilized to create RFs.
-During the process of obtaining RFs from FFs, we keep track of how frequently each RF appears in all FFs.
-This work is preparatory for the third step.
-An important aspect to note is that the resulting full RF set should be a subset of $S_{RF}$.
-This can be seen as follows: $S_{RF}$ encompasses all RFs obtained by partitioning FFs with the “ppppdddd” format in all possible ways.
-However, some RFs in $S_{RF}$ may never appear in any FF in $S_{FF}$.
-Specifically, these RFs might only be combined with specific other RFs to form FFs that are dominated.
-The third step involves selecting FFs for decomposition, aiming to create a mixed fragment set comprising both FFs and RFs.
-The main objective of working with such a mixed set is to reduce the fragment count.
-An initial assumption is that $S_{FF}$ typically contains more fragments than $S_{RF}$.
-Intuitively, decomposing all FFs into RFs might seem beneficial.
-However, we find that if an FF includes several RFs and these RFs do not appear in any other non-dominated FFs in $S_{FF}$, decomposing it can increase the total fragment count.
-Furthermore, our computational results in Section 5.2.1 indicate that when RFs are used in the formulations, they may lead to weaker lower bounds compared to FFs.
-Based on these two observations, we only decompose a portion of FFs in $S_{FF}$.
-Specifically, we maintain a record of how frequently each RF appears in the FFs from the previous step.
-Only the FFs for which all the constituent RFs appear multiple times (at least three times, in our implementation) within all FFs in $S_{FF}$ are candidates for decomposition.
-Other FFs that do not meet this criterion are retained in their original form.
-As a result, the undecomposed FFs together with the RFs obtained through this process constitute $S_{MF}$.
-Thus, it can be seen that $S_{MF}$ can serve the same role as $S_{FF}$ in enumerating all (M)DARP-LPT paths and act as the input for formulations without compromising optimality.
+4.2 EFF and ERF generation
+Rist & Forbes [17] introduce the concept of EFF, which is a combination of an FF with a node arc.
+Similarly, we define an ERF as a combination of an RF with a node arc.
+To construct these EFFs and ERFs, we first generate all FFs and RFs as in Section 4.1.
+Then, each fragment is extended by appending either a pickup node or the destination depot.
+The resulting extended fragments are retained if their corresponding route paths have a feasible time schedule.
+After the extension, the route path of the fragment becomes longer.
+The end node retains the same location information as the newly added node, while its load information remains unchanged, inheriting the load from the original fragment’s end node.
+Specifically, if the extended node is a pickup node rather than the destination depot, we construct an additional route path by further appending the corresponding delivery location of the extended pickup node.
+The extended fragment is retained only if this constructed path has a feasible schedule.
+Other processing steps, such as eliminating dominated fragments, remain unchanged.
+However, specific procedures such as combining with designated fragments for further feasibility checks are omitted, as the current enumeration process is already time-consuming and may largely exceed the actual solving time.
+An additional point of attention is that when we use EFFs or ERFs, the network flow constraints and the lazy constraints in formulations FFF and PSFF need to be updated, as most node arcs are eliminated and fragments are directly connected to each other.
+Specifically, in FFF we replace constraints (18) and (19) with the revised constraint (38), while in PSFF we update constraints (26) and (27) to (39).
+The lazy constraints (32) and (33) are also adapted accordingly.
+$\sum_{f\in F_{h}^{-}}x_{fv}$ $=\sum_{f\in F_{h}^{+}}x_{fv}$ $h\in P_{N},v\in V$ (38)
+$\sum_{f\in F^{-}_{h}}X_{f,l+L_{f}}$ $=\sum_{f\in F^{+}_{h}}X_{fl}$ $h\in P_{N},l\in\mathcal{L_{P}}$ (39)
+4.3 Mixed fragment set construction
+In this subsection we introduce a new fragment set denoted as $S_{MF}$, which consists of a mixture of RFs and ERFs.
+This set is derived from the original full fragment set $S_{FF}$ and can be a substitute for $S_{FF}$ and $S_{ERF}$.
+The primary motivation for constructing such a mixed set is to reduce the overall number of fragments while preserving high-quality fragments, thereby accelerating the computation of fragment-based formulations.
+This new fragment set is motivated by the observation that both fragment length and count affect the quality and efficiency of fragment-based formulations.
+While longer fragments (e.g., FFs vs.
+RFs, or ERFs vs.
+RFs) strengthen the formulation, their corresponding set is usually larger and potentially slows down the computation, as illustrated in Section 5.
+We observe that many ERFs in $S_{ERF}$ never appear in any EFF in $S_{EFF}$; in particular, some ERFs only combine with specific others to form EFFs that are ultimately dominated.
+Therefore, rather than exhaustively enumerating all ERFs, we propose deriving ERFs directly from $S_{EFF}$ or $S_{FF}$.
+Ultimately, we choose to derive ERFs from $S_{FF}$ for two main reasons.
+First, enumerating $S_{FF}$ requires less computational effort than $S_{EFF}$.
+Second, even with reduced redundancy, the number of ERFs generated from $S_{EFF}$ may still be large.
+Considering these points, we decide to generate fragments from $S_{FF}$.
+This, however, results in a mix of ERFs and RFs, rather than pure ERFs.
+Nevertheless, this approach significantly reduces the total number of fragments and thereby ensures computational efficiency, as demonstrated in Section 5.
+The generation process and the reason why mixed fragment types arise are further explained below.
+To construct $S_{MF}$, we decompose each FF by identifying pickup locations whose predecessors are delivery locations.
+If a FF contains a number $w$ of such pickup locations, it is split into $w+1$ pieces.
+The first $w$ pieces are used to generate ERFs, while the final piece forms an RF.
+For example, consider a fragment with route path $(p_{1},p_{2},d_{2},p_{3},d_{3},p_{4},d_{4},d_{1})$.
+This is divided into the following three parts: $(p_{1},p_{2},d_{2},p_{3})$, $(p_{3},d_{3},p_{4})$, $(p_{4},d_{4},d_{1})$.
+These segments are transformed into ERFs and RFs using the method described in Section 4.1.
+During this transformation, it is necessary to incorporate load information at the pickup locations.
+For instance, the locations $p_{3}$ and $p_{4}$ are transformed into nodes $(p_{3},\{l_{1}\})$ and $(p_{4},\{l_{1}\})$, respectively.
+In addition, some FFs cannot be divided and are thus retained (e.g., $(p_{1},d_{1})$).
+Therefore, $S_{MF}$ actually includes a mix of ERFs, RFs, and FFs.
+Since many fragments are RFs or FFs rather than strictly ERFs or EFFs, the fragment set $S_{MF}$ maintains a relatively low overall fragment count.
+In the resulting network, compared with the full fragment-based network, connections at pickup and delivery nodes with empty loads remain unchanged: fragments connect to node arcs, and node arcs connect to fragments.
+For pickup nodes with non-empty loads, fragments are directly connected to other fragments.
+Notably, there are no delivery nodes associated with non-empty loads.
+As a result, the network flow constraints for FFF and PSFF at nodes with empty loads remain the same as in Section 3.3.3 and Section 3.3.4, respectively, while the constraints at pickup nodes with non-empty loads are updated as described in Section 4.2.
+It is worth noting that if all FFs are decomposed into RFs (instead of ERFs) and node arcs, the resulting RF set is actually a subset of $S_{RF}$.
+This is because $S_{RF}$ includes all RFs that can be generated from FFs (with the “ppppdddd” structure) in all possible ways.
+However, not all RFs in $S_{RF}$ appear in any FF from $S_{FF}$; some can only form dominated FFs when combined with others.
 5 Computational results
 5.1 Data generation and implementation details
-To assess computational efficiency of different formulations and fragment sets, we utilize benchmark instances from Cordeau (2006), which are widely adopted in studies such as Ropke et al.
-(2007); Gschwind & Irnich (2015); Rist & Forbes (2021).
-These instances comprise two types, A and B, with type A having a tighter maximum ride time.
+To assess computational efficiency of different formulations and fragment sets, we utilize benchmark instances from Cordeau [6], which are widely adopted in studies such as Ropke et al.
+[20], Gschwind & Irnich [13], Rist & Forbes [18].
+These instances comprise two types, A and B, with type A having a tighter time window.
 In type A, customer demand is unit with a vehicle capacity of three, while type B involves demands from one to six customers with a vehicle capacity of six.
-For DARP-LPT, we generate both basic and extended instances.
-The basic instances adjust the vehicle numbers $|V|$ in Cordeau’s instances.
+In addition, we utilize benchmark instances from the SDARP problem [16, 17], which feature tighter time windows (typically three to six times the direct travel time for each customer) compared to those of Cordeau [6].
+There are two main reasons for using this second dataset.
+First, the tighter time window constraints are more representative of our application scenarios.
+Second, we aim to evaluate the performance of $S_{EFF}$ for DARP-LPT, as it was shown to perform well for SDARP in Rist & Forbes [17].
+For DARP-LPT, we adapt the aforementioned instances by modifying the number $|V|$ of vehicles in each instance.
 Given that DARP-LPT incorporates an additional parameter $L$ for the maximal number of pickups per trip, ensuring coverage for $n$ customers requires a minimum of $\lceil n/L\rceil$ vehicles.
 Here, we set $|V|$ as $\lceil n/(L-1)\rceil$, which is slightly higher than the minimum required.
-We also create an extra instance set to which we refer as “extended” instances, following Gschwind & Irnich, in which we increase time windows and capacity by 1/3 relative to the basic instances, thereby enhancing the diversity of the problem instances.
-For MDARP-LPT, which assumes different drivers have distinct origins and destinations, we generate driver information based on Cordeau’s instances.
-Specifically, the first $\lceil n/\alpha\rceil$ customers are designated as drivers, whose pickup and delivery nodes becoming the origins and destinations, while the remaining customers are passengers.
-The value $n$ represents the number of customers in each Cordeau instance, and $\alpha$ is an auxiliary parameter.
-The choice $L=4$ and $\alpha=4$ implies that one driver serves three customers on average, and $\alpha=3$ has one driver per two customers on average.
+For MDARP-LPT, which assumes different drivers have distinct origins and destinations, we generate driver information based on the aforementioned instances.
+Specifically, the first $\lceil n/3\rceil$ customers are designated as drivers, whose pickup and delivery nodes become the origins and destinations, while the remaining customers are passengers.
+The value $n$ represents the number of customers in each original instance.
 For drivers’ time windows, we set the earliest departure to 0, maintaining latest arrival times.
-Their maximum ride time is halved from the original driver’s maximum ride time (which is between six and 12 hours), as the latter is excessive for realistic scenarios such as carpooling.
 Unit customer demand and capacity 4 are chosen.
 Table 4 summarizes our experimental design.
-We use $S_{PFF}$ for MDARP-LPT and $S_{MF}$ for DARP-LPT.
-This decision is based on the appropriateness of $S_{PFF}$ for low $L$ values, which aligns well with real carpooling scenarios in MDARP-LPT.
-On the other hand, $S_{MF}$ is used for comparison with $S_{RF}$ and $S_{FF}$ across diverse $L$ values, a context more suitable for DARP-LPT.
-We do not apply $S_{MF}$ for MDARP-LPT because each vehicle has distinct fragment sets, and mixing the FFs and RFs per vehicle would become overly complex.
 Table 4: Experimental design
-[b] Problem Instance $L$ $|V|$ Customer number Formulation Fragment set DARP-LPT basic $4,6$ $\lceil n/(L-1)\rceil$ $n$ ABF, FAF, FFF $S_{FF}$, $S_{RF}$, $S_{MF}$ extended MDARP-LPT basic 4 $\lceil n/\alpha\rceil$ $n-\lceil n/\alpha\rceil$ FFF $S_{FF}$, $S_{RF}$, $S_{PFF}$
+[b] Problem $L$ $|V|$ Cus Formulation Fragment set DARP-LPT $4,6$ $\lceil n/(L-1)\rceil$ $n$ ABF, FFF, PSFF, PBF $S_{FF}$, $S_{RF}$, $S_{EFF}$, $S_{ERF}$, $S_{MF}$ MDARP-LPT $4$ $\lceil n/3\rceil$ $\lfloor 2n/3\rfloor$ PSFF, PBF $S_{FF}$, $S_{RF}$, $S_{MF}$
   • 1
-    Value $n$ is the number of customers in the original instance of Cordeau (2006).
+    Cus = customer number; Value $n$ is the number of customers in the original instance.
 All our algorithms are implemented using the Python programming language.
@@
 For DARP-LPT, fragment generation occurs in a single-processing environment, whereas for the MDARP-LPT, fragment generation is accomplished via multi-processing.
-Since this study involves two problems (DARP-LPT and MDARP-LPT), two different fragment-based formulations, four types of fragment sets, and two types of instances (basic and extended), we need to test multiple combinations.
-Therefore, from the existing benchmark instances, we only choose a subset, but we ensure a variety of sizes as indicated by the instance names in all following tables.
+Since this study involves two problems (DARP-LPT and MDARP-LPT), four different formulations, and five types of fragment sets, we need to test multiple combinations.
+Therefore, from the benchmark instances provided by Cordeau [6], we only choose a subset, but we ensure a variety of sizes as indicated by the instance names in all following tables.
+As the instances from Riedler & Raidl [16] have tighter time windows, which adheres more closely to the carpooling and crowdshipping scenarios for the problem addressed in this study, we retain all their instances.
 5.2 Discussion of the results
@@
 Time (s) indicates the total time, including the preprocessing time, network time, and CPU time for solver execution.
-Here, preprocessing time refers to the duration spent on time window tightening and arc elimination, as detailed by Cordeau (2006).
-Network time refers to the time required for fragment generation and the construction of the fragment-based network.
+Here, preprocessing time refers to the duration spent on time window tightening and arc elimination, as detailed by Cordeau [6].
+Network time refers to the time required for fragment generation and the construction of the fragment-based network (in FFF and PSFF), or for path generation (in PBF).
 OBJ and LB are the best-found integer solution value and the best lower bound within a 30-minute timeframe, respectively, and gap represents the relative percentage difference between OBJ and LB.
@@
 Fragment count is the total number of fragments in the fragment set.
-5.2.1 Results for DARP-LPT
-Table 5: Computational performance of different formulations for DARP-LPT within 30 minutes
-[b] ABF FAF ($S_{FF}$) FFF ($S_{FF}$) Name Time (s) OBJ LB Gap Time (s) OBJ LB Gap Time (s) OBJ LB Gap a3-18 24.3 302.5 302.5 0.00% 0.2 302.5 302.5 0.00% 0.2 302.5 302.5 0.00% a3-30 1800.0 496.7 451.2 9.16% 1800.0 494.8 491.6 0.64% 1800.0 494.8 491.1 0.74% a3-36 1800.0 574.4 541.2 5.78% 1800.0 574.4 565.9 1.49% 1800.0 574.4 565.0 1.63% a4-48 1800.0 712.0 578.3 18.78% 1800.0 698.6 685.4 1.90% 1800.0 697.9 687.4 1.50% a6-60 1800.0 920.2 635.4 30.95% 1800.0 879.2 843.2 4.09% 1800.0 864.9 843.4 2.48% a7-70 1800.0 1284.3 664.6 48.25% 1800.0 1004.0 958.2 4.56% 1800.0 986.6 956.5 3.05% a8-80 1800.0 1359.6 670.4 50.69% 1800.0 1100.7 1017.9 7.53% 1800.0 1071.1 1019.3 4.84% Avg-A 1546.3 23.37% 1542.9 2.89% 1542.9 2.03% b3-36 1800.0 576.5 548.4 4.88% 1800.0 576.5 566.8 1.69% 1800.0 576.5 567.7 1.53% b4-40 1800.0 637.5 612.5 3.92% 1800.0 635.2 630.2 0.79% 1800.0 635.2 631.1 0.64% b5-50 1800.0 796.5 659.3 17.22% 1800.0 798.4 765.7 4.10% 1800.0 789.1 768.3 2.63% b6-60 1800.0 937.9 749.9 20.05% 1800.0 940.6 896.5 4.68% 1800.0 931.7 900.7 3.33% b7-70 1800.0 1162.9 799.9 31.21% 1800.0 965.3 919.4 4.75% 1800.0 950.4 920.7 3.13% Avg-B 1800.0 15.46% 1800.0 3.20% 1800.0 2.25%
+Table 5: Computational performance of different formulations for DARP-LPT (30-minute limit)
+ABF FFF ($S_{FF}$) PSFF ($S_{FF}$) PBF
+Name Time OBJ LB Gap Time OBJ LB Gap Time OBJ LB Gap Net CPU Time OBJ LB Gap
+a3-18 24.3 302.5 302.5 0.00% 0.2 302.5 302.5 0.00% 0.1 302.5 302.5 0.00% 0.4 0.0 0.5 302.5 302.5 0.00%
+a3-30 1800.0 496.7 451.2 9.16% 1800.0 494.8 491.1 0.74% 1.2 494.8 494.8 0.00% 2.5 0.3 3.2 494.8 494.8 0.00%
+a3-36 1800.0 574.4 541.2 5.78% 1800.0 574.4 565.0 1.63% 0.4 574.4 574.4 0.00% 6.6 1.3 9.1 574.4 574.4 0.00%
+a4-48 1800.0 712.0 578.3 18.78% 1800.0 697.9 687.4 1.50% 0.7 697.9 697.9 0.00% 26.0 3.0 33.2 697.9 697.9 0.00%
+a6-60 1800.0 920.2 635.4 30.95% 1800.0 864.9 843.4 2.48% 2.7 864.9 864.9 0.00% 5.2 1.0 7.1 864.9 864.9 0.00%
+a7-70 1800.0 1284.3 664.6 48.25% 1800.0 986.6 956.5 3.05% 8.3 984.1 984.0 0.00% 9.9 1.8 13.4 984.1 984.0 0.00%
+a8-80 No 1800.0 1071.1 1019.3 4.84% 10.0 1058.7 1058.7 0.00% 25.9 5.6 35.7 1058.7 1058.7 0.00%
+a8-96 No 1800.0 1353.6 1302.7 3.76% 13.3 1342.2 1342.2 0.00% No
+Avg-A 1504.1 18.82% 1512.9 2.25% 4.59 0.00% 14.6 0.00%
+b3-36 1800.0 576.5 548.4 4.88% 1800.0 576.5 567.7 1.53% 0.4 576.5 576.5 0.00% 59.4 12.4 83.0 576.5 576.5 0.00%
+b4-40 1800.0 637.5 612.5 3.92% 1800.0 635.2 631.1 0.64% 1.0 635.2 635.2 0.00% 121.0 35.2 174.6 635.2 635.2 0.00%
+b5-50 1800.0 796.5 659.3 17.22% 1800.0 789.1 768.3 2.63% 3.1 789.1 789.1 0.00% 239.9 95.8 371.5 789.1 789.1 0.00%
+b6-60 1800.0 937.9 749.9 20.05% 1800.0 931.7 900.7 3.33% 4.9 927.1 927.1 0.00% 59.9 25.9 95.8 927.1 927.1 0.00%
+b7-70 1800.0 1162.9 799.9 31.21% 1800.0 950.4 920.7 3.13% 7.9 944.7 944.7 0.00% 105.1 33.9 158.0 944.7 944.7 0.00%
+b8-80 1800.0 1741.53 889.69 0.49% 1800.0 1104.8 1072.0 2.97% 3.6 1103.2 1103.2 0.00% No
+b8-96 No 1800.0 1322.6 1260.9 4.67% 7.4 1294.4 1294.4 0.00% No
+Avg-B 1800.0 12.96% 1800.0 2.70% 4.04 0.00% 176.6 0.00%
   • 1
+    Net: network time; CPU: CPU time for solver.
 The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
-Table 5 compares the three formulations ABF, FAF, and FFF for DARP-LPT for a runtime limit of 30 minutes, using the basic instances with the parameters $|V|=\lceil n/3\rceil$ and $L=4$.
-Both FAF and FFF utilize $S_{FF}$ as the input fragment set.
-The table indicates that, as observed from the time and gap indicators, both FFF and FAF largely outperform ABF, with FFF exhibiting superior performance compared to FAF.
-$S_{RF}$ $S_{FF}$ $S_{MF}$ Type F Net Time Gap F Net Time Gap F Net Time Gap Basic, $L=4$ A 475.7 0.4 1542.9 2.17% 567.3 1.4 1542.9 2.03% 452.0 1.7 1543.0 1.92% B 579.6 0.4 1800.0 2.42% 808.0 1.5 1800.0 2.25% 506.0 1.9 1800.0 2.32% Basic, $L=6$ A 475.7 0.4 1107.8 0.86% 937.3 5.8 1034.5 0.68% 454.6 5.8 1057.9 0.65% B 579.6 0.4 1444.6 0.57% 2449.2 10.2 1422.5 0.47% 511.8 11.1 1443.4 0.45% Extended, $L=4$ A 661.0 0.5 1542.9 2.44% 875.3 2.2 1542.9 2.15% 612.7 3.0 1543.0 2.17% B 1914.0 1.7 1800.0 3.67% 3070.0 8.7 1800.0 2.87% 1376.0 8.5 1800.0 2.94% Extended, $L=6$ A 661.0 0.5 1031.5 0.99% 1807.3 15.0 1031.7 0.91% 620.6 14.6 1030.8 0.94% B 1914.0 1.9 1228.9 1.08% 20382.6 204.3 1104.4 1.25% 1495.6 202.8 1191.7 0.96% 1 Type: instance type; F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver; basic: basic instances; extended: extended instances.
-Table 6: Average computational performance of FFF across all instances using various fragment sets for DARP-LPT within a 30-minute timeframe
-Table 6 presents a comparative analysis of three fragment set types $S_{RF}$, $S_{FF}$, and $S_{MF}$.
-These fragment sets are utilized as input for formulation FFF for solving DARP-LPT.
-The table covers both basic and extended instances, with $|V|=\lceil n/3\rceil$ when $L=4$ and $|V|=\lceil n/5\rceil$ when $L=6$.
-Due to the extensive amount of results, we present only the average metrics (fragment count, network time, total time, and gap) for all instances here.
-Detailed results for basic and extended instances with $L=4$ and $L=6$ are available in Tables 1, 2, 3, and 4 in the Appendix.
-We begin by comparing the two traditional fragment sets $S_{FF}$ and $S_{RF}$.
-The set $S_{FF}$ is better than $S_{RF}$ in most scenarios, particularly excelling in scenarios with a limited number of pickups per trip (e.g., $L=4$), flexible time windows (extended instances), and increased maximum ride time (instance type B), as reflected in the columns Gap.
-However, for extended type B instances with $L=6$, $S_{RF}$ may outperform $S_{FF}$.
-This variation can be attributed to fragment count and the quality of lower bound.
+“No” indicates that no feasible solution was obtained within the 30-minute time limit.
+Table 5 compares the four formulations ABF, FFF, PSFF, and PBF for DARP-LPT for a runtime limit of 30 minutes, using the instances of Cordeau [6] with the parameters $|V|=\lceil n/3\rceil$ and $L=4$.
+Both FFF and PSFF utilize $S_{FF}$ as the input fragment set.
+The table indicates that, as observed from the time and gap indicators, FFF and PSFF largely outperform ABF, with PSFF exhibiting superior performance compared to FFF.
+Even with $L=4$, PBF performs worse than PSFF in terms of the total time.
+For larger instances (e.g., a8-96 and b8-96), PBF does not have sufficient time to generate all the required paths.
+Table 6: Average computational performance of different fragment sets for DARP-LPT and MDARP-LPT (30-minute limit)
+$S_{RF}$ $S_{FF}$ $S_{MF}$ $S_{EFF}$ $S_{ERF}$
+Problem $L$ Type F CPU T F CPU T F CPU T F CPU T F CPU T
+DARP -LPT 4 A 599.4 0.6 1.5 713.0 1.0 4.5 700.3 1.1 3.6 15352.3 1.5 21.5 46594.0 6.6 27.9
+B 1231.9 1.9 3.2 1777.7 1.7 9.4 1287.6 1.8 7.9 24920.4 2.7 54.9 103744.7 7.2 55.6
+6 A 599.4 9.9 11.2 1153.0 10.8 21.4 707.6 11.1 20.4 29132.4 19.5 69.5 46594.0 207.7 229.9
+B 1231.9 9.9 12.0 9905.6 20.5 164.1 1439.0 16.3 170.3 6/7 solved 108323.7 46.7 104.1
+MDARP -LPT 4 A 252.5 19.0 35.0 252.5 20.1 39.1 264.9 19.2 36.5
+B 2079.6 82.4 124.4 2285.4 47.5 126.9 1708.4 71.4 154.8
+DARP -LPT 4 Ried -ler 1301.5 3.4 4.8 2328.4 2.1 6.3 1702.9 2.3 6.2 17471.0 2.6 18.0 60836.1 19.5 46.0
+6 1301.5 43.8 45.6 11845.3 49.6 83.2 1782.6 31.0 63.0 73336.5 64.4 172.8 26/30 solved
+MDARP -LPT 4 Ried -ler 454.7 17.2 31.6 633.87 14.6 31.9 547.0 14.3 32.0
+  • 1
+    Type denotes the instance type: A and B refer to type A and type B instances from Ropke & Cordeau [19], respectively; Riedler refers to instances from Riedler & Raidl [16].
+F: fragment count; CPU: CPU time for solver; T: Total time (s); “–” denotes that no further testing is performed for this scenario.
+Table 6 presents a comparative analysis of the five fragment sets $S_{RF}$, $S_{FF}$, $S_{MF}$, $S_{EFF}$, and $S_{ERF}$ for both DARP-LPT and MDARP-LPT.
+All these fragment sets are utilized as input for formulation PSFF.
+The table covers the instances from Cordeau [6] and Riedler & Raidl [16], with varying $L$.
+Due to the extensive amount of results, we present only the average metrics (fragment count, CPU time, and total time) for all instances here.
+The total time mainly comprises network time and CPU time, with all other components contributing less than 1%.
+Since both CPU time and total time are reported, we do not report network time separately.
+Detailed results for all instances are provided in B.
+The table provides several key insights.
+Our analysis focuses on three main aspects: determining the best fragment set for the current (M)DARP-LPT problem, evaluated by total time; assessing fragment set quality, measured by CPU time alone; and explaining why different fragment sets exhibit varying performance.
+Table 6 leads to several noteworthy observations.
+First, when fed into the formulation PSFF, the fragment set $S_{RF}$ consistently outperforms alternative sets in terms of the total time to reach optimality across all problem instances.
+Second, when CPU time is considered independently (that is, excluding network time) $S_{FF}$ achieves performance comparable to or even surpassing that of $S_{RF}$ for instances with tight time windows and low values of $L$ (e.g., in DARP-LPT and MDARP-LPT, for instances from Riedler & Raidl [16] with $L=4$).
+By contrast, in DARP-LPT, for type-B instances with $L=6$, $S_{RF}$ demonstrates a clear advantage over $S_{FF}$, primarily due to the substantially higher fragment count of $S_{FF}$.
+Third, $S_{MF}$ proves to be an effective alternative of $S_{FF}$ when the fragment count of $S_{FF}$ becomes prohibitively large.
+For example, for Riedler & Raidl’s instances in DARP-LPT with $L=6$, $S_{MF}$ is the most effective fragment set in terms of CPU time performance.
+Fourth, both $S_{EFF}$ and $S_{ERF}$ demonstrate inferior performance in terms of the total time to reach optimality, even when considering CPU time alone.
+This is primarily due to their excessive fragment counts.
+Analogously to the comparison between $S_{FF}$ and $S_{RF}$, $S_{EFF}$ achieves better performance than $S_{ERF}$ when $L$ is small and the instances have tight time windows and ride times.
+Given that $S_{ERF}$ and $S_{EFF}$ demonstrate significantly inferior performance compared to other fragment sets in DARP-LPT, we do not further evaluate them for MDARP-LPT.
+The performance differences observed across fragment sets can be largely attributed to two factors: fragment count and the quality of the resulting lower bound.
+The comparison between $S_{FF}$ and $S_{RF}$ is particularly illustrative of this underlying logic.
 When $L$ is lower, or when time windows or maximum ride time are tight, the fragment counts for $S_{FF}$ and $S_{RF}$ are similar or at least of the same magnitude.
@@
 Additionally, the fragments from $S_{FF}$ are longer than those from $S_{RF}$, in a similar way as paths are longer than individual arcs.
-This similarity helps to explain why $S_{FF}$ generally achieves better lower bounds than $S_{RF}$, because path-based (extended) formulations for routing problems typically yield superior lower bounds compared to arc-based formulations (Desrosiers et al., 2024).
-The mixed fragment set $S_{MF}$ demonstrates superior solution quality in scenarios with more flexible time windows and maximum ride time and higher values of $L$ (extended type B instances with $L=6$), as indicated by the gap metrics.
-Notably, the fragment count of $S_{MF}$ remains relatively low, especially when compared to $S_{FF}$.
-Since a larger number of fragments can slow down fragment-based formulations, $S_{MF}$ presents itself as an efficient alternative to $S_{FF}$ in this regard.
-5.2.2 Results for MDARP-LPT
-$S_{RF}$ $S_{FF}$ $S_{MF}$ $|V|$ Type F Net Time Gap F Net Time Gap F Net Time Gap $\lceil n/4\rceil$ A 269.1 8.2 813.9 1.50% 274.4 9.5 806.4 1.19% 162.7 69.7 777.4 0.82% B 1489.2 14.8 1471.1 3.02% 1650.6 20.5 1210.7 1.91% 674.4 96.6 797.5 0.78% $\lceil n/3\rceil$ A 201.9 10.0 636.5 0.56% 206.4 10.9 464.3 0.59% 124.3 52.7 423.5 0.31% B 1195.0 16.7 1107.3 1.80% 1227.6 20.0 872.0 0.55% 518.8 90.9 552.6 0.36% 1 Type: instance type; F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
-Table 7: Average computational performance of FFF across all instances using various fragment sets for MDARP-LPT within a 30-minute timeframe (basic instances with $L=4$)
-Table 7 compares three types of fragment sets for MDARP-LPT using FFF under different parameters.
-Formulation FAF for DARP-LPT can be extended with additional constraints to handle also MDARP-LPT, but its performance significantly lags behind that of FFF, which is why we only present results for FFF and not FAF.
-The reason for not altering $L$ is that $L=4$ represents a commonly used parameter in carpooling scenarios (the realistic background of MDARP-LPT).
-Here, we present only the average metrics for all instances.
-Detailed results can be found in Table 1 and 2 in the Appendix.
-Table 7 demonstrates the superiority of $S_{PFF}$ over $S_{RF}$ and $S_{FF}$ in computational performance, particularly for handling large instances.
-This is evident from the consistently lower time and gap values achieved with $S_{PFF}$ compared to $S_{FF}$.
-Notably, $S_{PFF}$’s superiority can likely be attributed to its lower fragment count.
-Our findings also indicate that across most instances, $S_{FF}$ generally outperforms $S_{RF}$ in terms of both time efficiency and solution quality.
+This similarity helps to explain why $S_{FF}$ generally achieves better lower bounds than $S_{RF}$, because path-based formulations for routing problems typically yield superior lower bounds compared to arc-based formulations [8].
+Additionally, when transitioning to $S_{MF}$, the set effectively shifts fragments from $S_{FF}$ to ERFs and RFs.
+As a result, the average fragment length in $S_{MF}$ lies between that of $S_{FF}$ and $S_{RF}$, while the fragment count remains of the same order of magnitude as that of $S_{RF}$.
+This balance explains why $S_{MF}$ achieves superior CPU time performance in certain scenarios.
+Given the relative simplicity of the current problem and the efficiency of the PSFF formulation, the benefits of higher-quality fragments are outweighed by the overhead of fragment enumeration.
+However, under the slower FFF formulation, the advantage of $S_{FF}$ over $S_{RF}$ becomes more pronounced.
+For example, in DARP-LPT with $L=4$, using the FFF formulation with a 30-minute time limit (but excluding the instance b8-96, where $S_{RF}$ fails to produce a feasible solution), $S_{FF}$ achieves average optimality gaps of 2.25% (type A) and 2.37% (type B), compared to 2.43% and 2.53% for $S_{RF}$, respectively (see C).
+Moreover, when the problem becomes more realistic by introducing additional constraints, the fragment enumeration time tends to remain stable (as the number of feasible paths becomes even smaller), whereas the solution time grows exponentially.
+This underscores the increasing value of employing higher-quality fragment sets as problem difficulty increases.
+Additionally, even without considering more complex settings but simply increasing the instance size, $S_{FF}$ outperforms $S_{RF}$ in terms of total time; this can be seen, for example, in DARP-LPT with Riedler & Raidl’s largest instance $60N\_5K\_C$ and $L=6$ (where both $S_{MF}$ and $S_{FF}$ even incur higher network construction time).
+Although the average metrics may not fully capture the importance of fragment set quality, the results on larger instances indicate that more difficult cases favor good-quality fragment sets.
+One further thing to mention is that some shared mobility studies [26, 23] use EFFs or FFs for DARP or similar problems.
+While the enumeration time is quite high in our study, it is significantly lower in theirs.
+This is partly because their instances feature much tighter time windows, and partly because we invest additional time in fragment elimination to ensure the best possible fragment set.
 6 Summary and conclusions
@@
 Our main contributions entail the extensions of a fragment-based formulation, the creation of improved fragment sets, and the insights gained from comparing fragment sets.
-More specifically, we have modeled (M)DARP-LPT using two fragment-based formulations, FFF and FAF, both of which are extended from the DARP fragment-based formulation of Rist & Forbes (2021), transitioning from fragment-based decision variables to fragment-vehicle-based variables.
-We find that FFF achieves better results than FAF in terms of solving time and optimality gap, which confirms our theoretical finding that FFF will also yield better lower bounds.
-We have also introduced improved fragment sets as formulation inputs.
-The two benchmark fragment sets, the full fragments set $S_{FF}$ and the restricted fragments set $S_{RF}$, are generated using the methods described by Alyasiry et al.
-(2019) and Rist & Forbes (2021), respectively.
-Our two novel sets are the path-enumerated full fragment set $S_{PFF}$ and the mixed fragment set $S_{MF}$.
-The generation of $S_{PFF}$ involves exhaustively enumerating paths, selectively choosing those meeting specific criteria, and then decomposing them into full fragments.
-This approach draws inspiration from the observation that carpooling and crowdshipping DARP routes typically involve a limited number of customers.
-We construct the set $S_{MF}$ by decomposing some of the full fragments in $S_{FF}$ into restricted fragments while preserving the remainder.
-This approach is designed for scenarios when $L$ is higher.
-All these fragment sets can be employed as input for fragment-based formulations to reach optimality.
-Our computational results demonstrate that for MDARP-LPT with $L=4$, a realistic carpooling setting, $S_{PFF}$ outperforms $S_{FF}$, and $S_{FF}$ performs better than $S_{RF}$ in terms of computational performance.
-For DARP-LPT, $S_{FF}$ generally outperforms $S_{RF}$ except when the instances have more flexible time windows, increased maximum ride times, and higher $L$ values.
-In such cases, $S_{MF}$ can replace $S_{FF}$ to achieve the best performance.
-We conclude with two supplementary observations.
-First, in DARP-LPT, FAF may outperform FFF for a number of small instances, as shown in Table 5.
-This advantage becomes more prevalent when a fixed number of vehicles is utilized.
-The underlying reason for this, intuitively, is that FAF, having fewer variables than FFF, solves its linear relaxations more quickly, but produces worse lower bounds.
-Thus, in small instances with fairly simple branch-and-bound trees, FAF tends to perform better.
-Conversely, in large instances, FAF’s performance tends to be worse than FFF.
-Additionally, when the number of vehicles is fixed, the complexity of the problem is reduced, allowing FAF to perform better than FFF in more scenarios.
-Second, the technique of variable fixing could be a viable option to reduce the number of fragments by removing those whose reduced cost is higher than the gap between a feasible solution’s objective value and the linear relaxation bound (Desaulniers et al., 2020).
-In the instances that we have tested, however, even if we set the objective value as the optimal value, this method typically removes only a small number of fragments, which hardly affects the final computational results.
-Appendix A Arc-based formulation of MDARP-LPT
-For MDARP-LPT, the full arc-based formulation (following Cordeau (2006)) is as follows:
-Table 1: Sets, parameters, and decision variables of arc-based formulation for MDARP-LPT
-Sets Definition
-$N_{v}$ Set of locations, $N=P_{v}\cup D_{v}\cup\{o_{v},d_{v}\}$, where $o_{v}$ and $d_{v}$ are respectively
-origin and destination of vehicle $v\in V$
-$P_{v}$, $D_{v}$ Pickup and delivery locations of vehicle $v\in V$
-$A_{v}$ Set of location arcs of vehicle $v\in V$
-Parameters Definition
-$e_{iv}$, $l_{iv}$ The earliest and latest time of vehicle $v$ at location $i\in N$
-Decision variables Definition
-$f_{ijv}$ =1, if the arc $(i,j)\in A_{v}$ is traversed by vehicle $v\in V$; =0, otherwise
-$t_{iv}$ the departure time of vehicle $v$ on location $i\in N_{v}$
-$Q_{iv}$ the load of vehicle $v$ after visiting location $i\in N_{v}$
-$\min\sum_{v\in V}\sum_{(i,j)\in A_{v}}C_{ij}f_{ijv}$ (40)
-$s.t.$ $\sum_{v\in V}\sum_{j\in N_{v}}f_{ijv}$ $=1$ $i\in P$ (41)
-$\sum_{j\in N_{v}}f_{ijv}$ $=\sum_{j\in N_{v}}f_{(i+n)jv}$ $i\in P_{v},v\in V$ (42)
-$\sum_{j\in N_{v}}f_{o_{v}jv}$ $=1$ $v\in V$ (43)
-$\sum_{j\in N_{v}}f_{jiv}$ $=\sum_{j\in N_{v}}f_{ijv}$ $i\in P_{v}\cup D_{v},v\in V$ (44)
-$\sum_{i\in N_{v}}f_{id_{v}v}$ $=1$ $v\in V$ (45)
-$t_{iv}+T_{ij}$ $\leq t_{jv}+M_{ijv}(1-f_{ijv})$ $(i,j)\in A_{v},v\in V$ (46)
-$e_{iv}\leq t_{iv}$ $\leq l_{iv}$ $i\in N_{v},v\in V$ (47)
-$Q_{iv}+q_{i}$ $\leq Q_{jv}+W_{ij}(1-f_{ijv})$ $(i,j)\in A_{v},v\in V$ (48)
-$\max(0,q_{i})\leq Q_{iv}$ $\leq\min(Q,Q+q_{i})$ $i\in N_{v},v\in V$ (49)
-$T_{i(n+i)}\leq t_{(n+i)v}-t_{iv}$ $\leq R_{i}$ $i\in P_{v},v\in V$ (50)
-$t_{(2n+1)v}-t_{0v}$ $\leq R_{0}$ $v\in V$ (51)
-$\sum_{(i,j)\in A_{v}}f_{ijv}$ $\leq 2L+1$ $v\in V$ (52)
-$f_{ijv}$$\in\{0,1\}$ $(i,j)\in A_{v},v\in V$ (53)
-To be concise, we outline the differences between the formulations of MDARP-LPT and DARP-LPT.
-Constraints (4) to (6) in DARP-LPT are modified to constraints (43) to (45) in MDARP-LPT, accommodating the network flow constraints for distinct vehicles with unique origins and destinations.
-Constraints (52) can be modified to $\sum_{(i,j)\in A_{v}}\max(q_{i},0)f_{ijv}\\ \leq 2\lambda+1$ if the constraint on the maximum number of pickups per trip is replaced by the constraint on the maximum quantity of items (or passengers) $\lambda$ per trip, where $q_{i}$ represents the customer load quantity.
-Importantly, the constraints on $\lambda$ remain applicable within our fragment-based approach.
-The necessary modification involves updating constraint (21) in FFF and constraint (32) in FAF from $\sum_{f\in F}L_{f}x_{fv}\leq L$ to $\sum_{f\in F}\lambda_{f}x_{fv}\leq\lambda$, where $\lambda_{f}$ represents the total quantity of items in fragment $f$.
-Appendix B Full computational results for DARP-LPT
-All four tables below serve as supplements to Table 6.
-$S_{RF}$ $S_{FF}$ $S_{MF}$ Name F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 73 0.3 0.0 0.4 302.5 302.5 0.00% 64 0.1 0.0 0.2 302.5 302.5 0.00% 64 0.3 0.2 1.0 302.5 302.5 0.00% a3-30 93 0.1 1799.9 1800.0 494.8 491.0 0.75% 89 0.1 1799.9 1800.0 494.8 491.1 0.74% 89 0.5 1798.8 1800.0 494.8 489.7 1.01% a3-36 121 0.1 1800.5 1800.0 574.4 560.6 2.40% 112 0.2 1797.6 1800.0 574.4 565.0 1.63% 113 0.1 1799.8 1800.0 574.4 566.7 1.35% a4-48 318 0.2 1798.9 1800.0 697.9 686.1 1.69% 312 0.5 1799.4 1800.0 697.9 687.4 1.50% 303 0.5 1798.8 1800.0 697.9 687.9 1.43% a6-60 601 0.4 1793.4 1800.0 869.0 843.2 2.97% 648 1.4 1795.5 1800.0 864.9 843.4 2.48% 579 1.8 1795.7 1800.0 864.9 844.8 2.32% a7-70 1036 0.8 1796.3 1800.0 986.3 958.5 2.82% 1273 2.6 1792.3 1800.0 986.6 956.5 3.05% 932 2.9 1795.1 1800.0 984.9 956.5 2.88% a8-80 1088 0.8 1789.1 1800.0 1064.6 1016.4 4.53% 1473 5.1 1785.7 1800.0 1071.1 1019.3 4.84% 1084 6.2 1782.4 1800.0 1063.4 1016.3 4.43% Avg-A 475.7 0.4 1542.9 2.17% 567.3 1.4 1542.9 2.03% 452.0 1.7 1543.0 1.92% b3-36 169 0.1 1799.9 1800.0 576.5 566.2 1.79% 161 0.1 1799.7 1800.0 576.5 567.7 1.53% 146 0.1 1799.8 1800.0 576.5 566.1 1.81% b4-40 311 0.2 1799.5 1800.0 635.2 630.4 0.76% 203 0.4 1799.1 1800.0 635.2 631.1 0.64% 202 0.4 1799.5 1800.0 635.2 631.8 0.54% b5-50 518 0.4 1798.6 1800.0 789.1 766.2 2.91% 803 1.2 1797.6 1800.0 789.1 768.3 2.63% 496 2.2 1796.2 1800.0 789.5 766.7 2.90% b6-60 940 0.6 1793.3 1800.0 931.4 895.5 3.86% 1218 2.6 1792.3 1800.0 931.7 900.7 3.33% 799 2.0 1794.2 1800.0 931.7 900.4 3.36% b7-70 960 0.7 1796.4 1800.0 948.0 921.6 2.79% 1655 3.2 1789.5 1800.0 950.4 920.7 3.13% 887 4.6 1789.1 1800.0 950.2 921.7 3.00% Avg-B 579.6 0.4 1800.0 2.42% 808.0 1.5 1800.0 2.25% 506.0 1.9 1800.0 2.32% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
+More specifically, we model (M)DARP-LPT using two fragment-based formulations, FFF and PSFF, both of which are extended from the DARP fragment-based framework proposed by Rist & Forbes [18].
+For FFF the extension resides in the transition from fragment-based to fragment–vehicle–based decision variables, whereas PSFF retains the original variables but augments the network with a pickup-count dimension, resulting in a pickup-space network.
+We find that PSFF achieves better results than FFF.
+We also compare several fragment sets used as inputs to the formulations.
+The full fragment set $S_{FF}$ and the restricted fragment set $S_{RF}$ are generated following the methods proposed by Alyasiry et al.
+[2] and Rist & Forbes [18], respectively.
+The extended full fragment set $S_{EFF}$ and the extended restricted fragment set $S_{ERF}$ are constructed based on the approach of Rist & Forbes [17].
+In addition, we develop the new mixed fragment set $S_{MF}$ by decomposing full fragments from $S_{FF}$ into (extended) restricted fragments.
+All these fragment sets can be used as inputs to fragment-based formulations to reach optimality.
+Our computational results show that, for both DARP-LPT and MDARP-LPT, $S_{RF}$ outperforms $S_{FF}$ in total solution time when used with the PSFF formulation.
+However, in terms of fragment set quality, measured solely by CPU time, $S_{FF}$ remains superior in instances with tight time windows and a limited number of pickups (which is also close to realistic application scenarios).
+In DARP-LPT, $S_{MF}$ serves as a high-quality alternative to $S_{FF}$, especially when the size of $S_{FF}$ becomes large.
+The incorporation of a time dimension into the formulations FFF or PSFF may be a valuable direction for future work.
+For example, PSFF can naturally be extended to a pickup-time-space formulation to incorporate both time and pickup dimensions.
+However, since vehicle indices and pickup count are already used as dimensions, and PSFF performs well under the current setup, we have not pursued this extension in this paper.
+Appendix A Additional explanation for fragment-based network
+Decomposing feasible DARP routes into fragments yields a significantly smaller number of fragments compared to DARP routes.
+Conversely, if we combine these fragments, we can obtain all feasible DARP routes, albeit with the possibility of some redundant routes.
+Fig.
+A.1 illustrates how feasible DARP routes can be gathered in a transitional network (as an intermediate step towards the fragment-based network).
+Fig.
+1 presents several DARP routes and their corresponding fragments.
+The network in Fig.
+A.1 will use these fragments and allow to enumerate all the DARP routes shown in Fig.
+1.
+Identical fragments across different routes in Fig.
+1 can be consolidated into a single fragment while maintaining the connections to other fragments.
+Subsequently, the fragment-based network of Fig.
+2 is derived from the network of Fig.
+A.1.
+When different fragments share the same start and/or end nodes, then those nodes can also be merged.
+For instance, two fragments with the same start and end nodes but different route paths $(p1,d1)$ and $(p1,p4,d4,d1)$ become two route paths originating from node $p1$ and terminating at node $d1$.
+Such consolidations help reduce connections (node arcs).
+Node arcs with the same start node and end node can also be merged; the three connections from $(p1,d1)$ to $(p2,d2)$, from $(p1,d1)$ to $(p2,p3,d2,d3)$, and from $(p1,p4,d4,d1)$ to $(p2,d2)$ in Fig.
+A.1, for instance, can be merged into one node arc $(d1,p2)$ in Fig.
+2.
+$O^{+}$$p3$($d2$,{$l_{3}$})($p1$,{$l_{3}$})$d3$$O^{-}$($p3$,$p2$,$d2$)($p1$,$d1$,$d3$)$p1$$d1$$p2$$d2$$p3$$d3$($p1$,$d1$)($p2$,$d2$)($p3$,$d3$)$p1$$d1$$p2$$d3$($p1$,$p4$,$d4$,$d1$)($p2$,$p3$,$d2$,$d3$)
+Figure A.1: A transitional network for DARP by merging all identical fragments from Fig. 1
+Appendix B Full computational results of PSFF for DARP-LPT and MDARP-LPT
+The tables below provide supplementary information an the results for PSFF, accompanying Table 6.
+As nearly all instances achieve optimality across different fragment set inputs, we report only one objective value (OBJ) in the first column and keep four indicators: fragment count, network time, CPU time, and the total time.
+The test instances comprises: (1) eight Type A and seven Type B instances from Cordeau [2006], and (2) 30 instances from Riedler & Raidl [2018].
+Tables 13 present the computational performance of PSFF with different fragment sets on the instances from Cordeau [2006].
+Specifically, Tables 1 and 2 correspond to DARP-LPT, while Table 3 focuses on MDARP-LPT.
+Similarly, Tables 46 evaluate PSFF’s performance on the instances from Riedler & Raidl [2018], with Tables 4 and 5 dedicated to DARP-LPT and Table 6 to MDARP-LPT.
+$S_{RF}$ $S_{FF}$ $S_{MF}$ $S_{EFF}$ $S_{ERF}$ Name OBJ F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time a3-18 302.5 73 0.1 0.0 0.1 64 0.1 0.1 0.2 71 0.1 0.0 0.2 455 0.2 0.0 0.3 1310 0.4 0.0 0.5 a3-30 494.8 93 0.1 0.1 0.2 89 0.2 0.1 0.4 95 0.1 0.1 0.3 1202 0.8 0.1 0.9 2902 0.8 0.1 1.1 a3-36 574.4 121 0.1 0.1 0.3 112 0.1 0.1 0.4 122 0.2 0.2 0.4 1730 1.1 0.1 1.5 4261 1.1 0.2 1.7 a4-48 697.9 318 0.2 0.3 0.7 314 0.9 0.1 1.2 338 0.8 0.1 1.3 4943 5.6 0.2 6.1 14468 5.5 0.4 7.2 a6-60 864.9 601 0.3 0.4 1.1 655 3.4 0.6 4.6 670 1.7 0.4 2.5 14542 12.5 1.5 15.0 48423 17.2 1.9 23.8 a7-70 984.1 1036 0.7 0.8 2.1 1306 4.0 1.9 6.8 1211 2.4 1.0 3.8 19146 22.5 2.1 26.5 57082 21.1 14.4 41.4 a8-80 1058.7 1088 0.6 1.1 2.4 1480 7.1 0.9 8.7 1386 4.5 0.9 5.8 33590 47.3 2.6 52.6 89453 30.3 15.2 56.0 a8-96 1342.2 1465 1.37 2.2 5.0 1684 7.9 4.3 13.3 1709 7.2 6.3 14.3 47210 59.5 5.5 68.8 154853 55.0 20.3 91.6 Avg-A 599.4 0.6 1.5 713.0 1.0 4.5 700.3 1.1 3.6 15352.3 1.5 21.5 46594.0 6.6 27.9 b3-36 576.5 169 0.1 0.1 0.3 166 0.2 0.2 0.5 168 0.2 0.1 0.4 1473 1.4 0.1 1.6 3834 1.2 0.1 1.6 b4-40 635.2 311 0.2 0.1 0.4 212 1.3 0.1 1.6 233 0.5 0.1 0.8 3572 3.9 0.2 4.3 14683 4.8 0.3 6.2 b5-50 789.1 518 0.3 0.7 1.2 804 1.6 0.4 2.5 649 1.4 0.6 2.3 9808 7.4 0.7 8.7 19161 6.3 1.0 9.3 b6-60 927.1 940 0.8 2.3 3.7 1231 2.8 1.6 4.8 1030 1.9 1.1 3.4 12290 15.1 2.1 18.6 28817 9.1 2.1 13.4 b7-70 944.7 960 0.8 4.6 5.9 1678 4.5 3.5 8.7 1188 3.3 2.4 6.0 18758 24.0 2.6 28.5 52404 21.2 3.7 30.0 b8-80 1103.2 1195 0.8 2.3 3.6 1813 5.2 1.5 7.2 1429 4.4 3.0 7.9 26484 43.4 3.0 49.3 71693 27.0 3.4 37.0 b8-96 1294.4 4530 3.1 3.1 7.4 6540 34.9 4.4 40.8 4316 28.5 5.3 34.9 102058 256.3 10.0 273.5 535621 204.4 39.5 291.9 Avg-B 1231.9 1.9 3.2 1777.7 1.7 9.4 1287.6 1.8 7.9 24920.4 2.7 54.9 103744.7 7.2 55.6 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; CPU: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
 The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
-Table 1: Computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes (Basic instances, $L=4$)
-$S_{RF}$ $S_{FF}$ $S_{MF}$ Name F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 73 0.2 0.3 1.0 295.8 295.8 0.00% 64 0.1 0.0 0.2 295.8 295.8 0.00% 64 0.1 0.0 0.2 295.8 295.8 0.00% a3-30 93 0.3 17.5 18.9 477.0 477.0 0.00% 89 0.1 4.5 4.7 477.0 477.0 0.00% 89 0.2 4.4 4.7 477.0 477.0 0.00% a3-36 121 0.1 534.4 535.0 549.8 549.8 0.00% 115 0.2 36.0 36.4 549.8 549.8 0.00% 120 0.2 200.2 200.6 549.8 549.8 0.00% a4-48 318 0.3 1798.6 1800.0 669.4 662.9 0.97% 342 0.7 1798.9 1800.0 669.4 664.4 0.74% 305 0.7 1798.3 1800.0 669.4 665.5 0.58% a6-60 601 0.4 1798.0 1800.0 825.0 819.0 0.72% 817 4.5 1796.3 1800.0 825.0 820.3 0.57% 589 3.6 1794.6 1800.0 825.0 820.6 0.54% a7-70 1036 0.7 1793.6 1800.0 926.7 912.4 1.54% 2114 10.8 1786.0 1800.0 927.6 912.8 1.60% 944 10.7 1786.6 1800.0 927.6 913.1 1.56% a8-80 1088 0.7 1790.5 1800.0 1000.9 973.1 2.77% 3020 24.2 1770.1 1800.0 991.3 973.1 1.83% 1071 25.3 1767.5 1800.0 992.8 974.0 1.89% Avg-A 475.7 0.4 1107.8 0.86% 937.3 5.8 1034.5 0.68% 454.6 5.8 1057.9 0.65% b3-36 169 0.1 22.7 23.1 557.9 557.9 0.00% 210 0.3 13.5 14.0 557.9 557.9 0.00% 154 0.3 16.2 16.8 557.9 557.9 0.00% b4-40 311 0.2 1799.3 1800.0 627.4 623.9 0.56% 228 1.5 1696.5 1698.5 627.4 627.4 0.00% 216 1.7 1797.9 1800.0 627.4 626.2 0.20% b5-50 518 0.3 1798.7 1800.0 751.9 745.3 0.87% 1868 5.9 1791.6 1800.0 751.9 744.2 1.03% 492 7.1 1793.4 1800.0 751.9 745.7 0.82% b6-60 940 0.6 1795.9 1800.0 877.0 868.0 1.03% 3769 14.7 1782.2 1800.0 877.0 867.1 1.13% 806 14.7 1783.4 1800.0 877.0 867.8 1.05% b7-70 960 0.7 1798.2 1800.0 896.4 892.9 0.39% 6171 28.5 1766.2 1800.0 896.4 894.6 0.21% 891 31.7 1765.5 1800.0 896.4 894.8 0.19% Avg-B 579.6 0.4 1444.6 0.57% 2449.2 10.2 1422.5 0.47% 511.8 11.1 1443.4 0.45% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
+Table 1: Computational performance of PSFF with different fragment sets for DARP-LPT within 30 minutes (Cordeau’s instances, $L=4$)
+$S_{RF}$ $S_{FF}$ $S_{MF}$ $S_{EFF}$ $S_{ERF}$ Name OBJ F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time a3-18 295.8 73 0.0 0.0 0.1 64 0.5 0.2 1.0 71 0.1 0.0 0.1 470 0.2 0.0 0.3 1310 0.4 0.1 0.6 a3-30 477.0 93 0.1 0.2 0.3 89 0.6 0.9 1.8 95 0.1 0.2 0.7 1250 0.7 0.3 3.1 2902 0.9 0.5 1.8 a3-36 549.8 121 0.1 0.1 0.3 115 0.4 0.1 0.9 122 0.2 0.1 0.4 1951 1.5 0.2 1.9 4261 1.4 0.2 2.3 a4-48 669.4 318 0.2 0.6 1.1 344 1.8 0.5 3.4 339 0.9 0.5 1.6 5818 7.4 1.3 9.5 14468 4.9 4.1 11.3 a6-60 825.0 601 0.3 0.9 1.7 827 5.2 1.1 6.8 673 3.5 0.8 4.7 23568 34.0 4.4 44.4 48423 14.1 19.5 41.3 a7-70 925.7 1036 0.7 6.6 8.3 2157 15.5 3.1 19.5 1233 13.3 2.6 16.5 38512 66.8 10.8 84.7 57082 17.2 126.5 153.1 a8-80 990.8 1088 0.9 8.9 11.1 3028 28.6 7.7 37.3 1390 25.5 6.7 33.0 81310 123.8 24.5 157.7 89453 26.7 317.5 359.6 a8-96 1262.3 1465 1.0 61.7 66.7 2600 25.8 73.2 100.1 1738 26.5 78.2 105.9 80180 128.0 114.4 254.6 154853 49.2 1192.9 1269.1 Avg-A 599.4 9.9 11.2 1153.0 10.8 21.4 707.6 11.1 20.4 29132.4 19.5 69.5 46594.0 207.7 229.9 b3-36 557.9 169 0.1 0.1 0.3 219 1.1 0.2 1.7 172 0.3 0.1 0.5 1723 2.0 0.1 2.3 3857 1.2 0.2 2.1 b4-40 627.4 311 0.2 0.7 1.1 238 5.0 1.0 6.3 244 2.6 1.0 3.8 5033 12.6 1.2 14.5 14683 4.8 1.9 8.8 b5-50 751.9 518 0.4 3.4 4.3 1869 8.6 1.3 10.4 653 5.5 0.9 6.6 31667 48.9 4.6 57.3 19197 5.8 1.6 10.4 b6-60 876.8 940 0.5 3.7 5.0 3818 16.5 2.1 19.3 1079 23.6 2.1 26.2 30029 73.5 10.7 89.8 28817 10.5 5.3 19.7 b7-70 896.4 960 1.0 2.2 4.5 6251 28.7 1.8 31.4 1252 38.3 2.1 41.1 67029 149.1 7.3 163.2 52603 23.0 6.8 38.4 b8-80 1053.8 1195 1.1 12.2 14.5 5896 34.9 14.2 50.1 1481 39.7 10.8 51.4 78389 173.2 14.8 196.8 71693 24.9 14.5 52.4 b8-96 1222.0 4530 3.8 47.1 54.1 51048 902.7 122.7 1029.7 5192 963.1 97.5 1062.4 No 567416 212.9 296.4 597.0 Avg-B 1231.9 9.9 12.0 9905.6 20.5 164.1 1439.0 16.3 170.3 6/7 solved 108323.7 46.7 104.1 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; CPU: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
+“No” indicates that no optimal solution is obtained within the 30-minute time limit.
 The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
-Table 2: Computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes (Basic instances, $L=6$)
-$S_{RF}$ $S_{FF}$ $S_{MF}$ Name F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 96 0.1 0.3 0.4 288.5 288.5 0.00% 81 0.1 0.5 0.6 288.5 288.5 0.00% 81 0.1 0.5 0.7 288.5 288.5 0.00% a3-30 106 0.1 1799.8 1800.0 491.4 486.6 0.97% 101 0.1 1799.7 1800.0 491.4 485.9 1.12% 101 0.1 1799.7 1800.0 491.4 486.7 0.96% a3-36 161 0.1 1799.5 1800.0 541.9 532.2 1.79% 159 0.2 1798.7 1800.0 541.9 535.4 1.20% 155 0.2 1799.7 1800.0 541.9 536.4 1.01% a4-48 420 0.2 1798.2 1800.0 697.8 680.2 2.52% 431 0.6 1798.0 1800.0 697.8 680.9 2.42% 407 0.7 1798.6 1800.0 697.9 679.9 2.58% a6-60 859 0.6 1793.8 1800.0 846.1 828.9 2.04% 1026 2.3 1794.7 1800.0 847.5 830.6 1.99% 830 2.2 1793.9 1800.0 846.1 826.6 2.31% a7-70 1502 1.2 1792.9 1800.0 971.5 939.8 3.26% 2055 5.0 1789.1 1800.0 968.3 941.4 2.78% 1285 6.8 1786.7 1800.0 968.9 940.5 2.93% a8-80 1483 1.1 1788.7 1800.0 1073.8 1004.4 6.46% 2274 7.0 1782.7 1800.0 1065.2 1006.0 5.56% 1430 10.7 1771.7 1800.0 1060.3 1003.0 5.41% Avg-A 661.0 0.5 1542.9 2.44% 875.3 2.2 1542.9 2.15% 612.7 3.0 1543.0 2.17% b3-36 460 0.3 1799.4 1800.0 556.5 540.9 2.80% 567 0.6 1799.1 1800.0 556.5 542.9 2.45% 435 0.7 1797.7 1800.0 556.5 544.0 2.25% b4-40 910 0.6 1798.1 1800.0 582.8 571.0 2.02% 906 2.1 1797.4 1800.0 582.8 573.4 1.61% 614 1.8 1797.2 1800.0 582.8 573.8 1.55% b5-50 1686 1.4 1792.0 1800.0 755.1 730.2 3.30% 3196 12.4 1785.3 1800.0 754.7 732.3 2.97% 1348 7.7 1789.4 1800.0 754.7 731.5 3.07% b6-60 3036 2.9 1790.3 1800.0 873.8 837.0 4.21% 3890 9.6 1781.1 1800.0 868.9 841.7 3.13% 1929 10.5 1782.0 1800.0 874.6 842.0 3.73% b7-70 3478 3.4 1753.1 1800.0 911.6 856.7 6.02% 6791 18.8 1762.9 1800.0 900.4 862.7 4.19% 2554 21.8 1755.4 1800.0 897.1 860.5 4.09% Avg-B 1914.0 1.7 1800.0 3.67% 3070.0 8.7 1800.0 2.87% 1376.0 8.5 1800.0 2.94% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
+Table 2: Computational performance of PSFF with different fragment sets for DARP-LPT within 30 minutes (Cordeau’s instances, $L=6$)
+$S_{RF}$ $S_{FF}$ $S_{MF}$ Name OBJ F Net CPU Time F Net CPU Time F Net CPU Time a3-18 206.7 37 8.1 0.1 8.3 33 7.2 0.1 7.4 35 14.0 0.1 14.3 a3-30 315.7 36 8.3 0.3 8.8 36 7.1 0.3 7.6 36 8.9 0.6 9.8 a3-36 379.9 47 7.4 0.5 10.6 44 16.9 0.6 20.2 47 7.8 0.4 8.7 a4-48 428.7 116 10.0 1.7 13.2 109 11.7 1.8 14.8 119 8.2 1.7 11.0 a6-60 532.2 175 10.2 5.0 17.5 156 10.4 7.4 20.0 176 9.2 5.1 16.5 a7-70 535.4 507 16.3 19.4 40.0 505 17.2 24.3 47.0 528 14.6 21.8 41.2 a8-80 637.3 495 21.2 37.5 65.6 586 20.8 39.5 68.9 576 21.1 42.3 70.4 a8-96 775.6 607 16.0 87.3 116.0 551 29.8 86.5 127.3 602 27.3 81.4 119.9 Avg-A 252.5 19.0 35.0 252.5 20.1 39.1 264.9 19.2 36.5 b3-36 364.1 146 7.5 0.5 8.6 148 8.9 0.5 10.0 147 9.4 0.7 12.1 b4-40 345.1 370 8.4 1.4 11.0 320 9.9 1.3 12.5 322 18.9 1.9 22.0 b5-50 417.0 1649 20.0 20.8 46.7 1682 32.2 14.7 51.4 1457 30.7 18.4 53.7 b6-60 517.9 2269 26.5 30.9 67.3 2084 31.6 21.6 58.8 1788 35.0 23.5 65.2 b7-70 512.4 1541 26.0 62.0 98.6 2049 48.9 27.6 86.4 1493 53.1 48.9 110.6 b8-80 601.0 2062 29.8 116.0 160.7 3133 88.4 96.0 200.7 2267 93.7 101.2 209.1 b8-96 714.8 6520 90.3 345.2 477.6 6582 266.8 170.8 468.5 4485 275.2 304.9 611.1 Avg-B 2079.6 82.4 124.4 2285.4 47.5 126.9 1708.4 71.4 154.8 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; CPU: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
 The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
-Table 3: Computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes (Extended instances, $L=4$)
-$S_{RF}$ $S_{FF}$ $S_{MF}$ Name F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 96 0.0 0.9 1.0 282.4 282.4 0.00% 88 0.1 0.6 0.8 282.4 282.4 0.00% 90 0.3 1.9 2.5 282.4 282.4 0.00% a3-30 106 0.1 11.9 12.2 472.5 472.5 0.00% 101 0.1 9.4 9.7 472.5 472.5 0.00% 101 0.3 8.2 8.7 472.5 472.5 0.00% a3-36 161 0.1 6.9 7.5 510.6 510.6 0.00% 179 0.4 10.6 11.6 510.6 510.6 0.00% 158 0.3 3.6 4.2 510.6 510.6 0.00% a4-48 420 0.3 1798.6 1800.0 661.2 656.5 0.70% 501 1.3 1797.6 1800.0 661.2 657.7 0.52% 404 1.3 1797.6 1800.0 661.2 657.4 0.57% a6-60 859 0.6 1797.8 1800.0 807.1 796.7 1.29% 1623 9.6 1790.0 1800.0 807.1 797.7 1.17% 833 9.8 1788.7 1800.0 807.1 797.6 1.18% a7-70 1502 1.1 1794.6 1800.0 908.5 891.3 1.89% 4266 26.8 1766.3 1800.0 907.9 891.7 1.79% 1341 28.1 1767.3 1800.0 908.5 892.6 1.75% a8-80 1483 1.1 1782.0 1800.0 988.3 958.3 3.03% 5893 66.6 1720.0 1800.0 986.9 958.2 2.90% 1417 62.1 1727.7 1800.0 989.3 959.1 3.06% Avg-A 661.0 0.5 1031.5 0.99% 1807.3 15.0 1031.7 0.91% 620.6 14.6 1030.8 0.94% b3-36 460 0.2 9.0 9.7 530.6 530.6 0.00% 1199 2.5 18.6 22.1 530.6 530.6 0.00% 436 2.2 12.0 14.9 530.6 530.6 0.00% b4-40 910 0.7 733.4 734.9 564.6 564.6 0.00% 2247 18.0 80.7 99.9 564.6 564.6 0.00% 676 18.5 524.5 543.8 564.6 564.6 0.00% b5-50 1686 1.4 1795.4 1800.0 710.8 701.4 1.33% 17112 90.7 1699.1 1800.0 710.8 703.0 1.10% 1420 94.8 1701.4 1800.0 710.8 702.8 1.14% b6-60 3036 3.1 1793.6 1800.0 817.8 799.3 2.26% 22909 177.8 1600.9 1800.0 818.7 799.3 2.38% 2135 185.2 1609.2 1800.0 818.3 804.5 1.69% b7-70 3478 4.1 1778.4 1800.0 837.7 822.7 1.79% 58446 732.6 957.9 1800.0 844.8 821.3 2.78% 2811 713.3 1075.2 1800.0 841.0 824.5 1.97% Avg-B 1914.0 1.9 1228.9 1.08% 20382.6 204.3 1104.4 1.25% 1495.6 202.8 1191.7 0.96% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
+Table 3: Computational performance of PSFF with different fragment sets for MDARP-LPT within 30 minutes (Cordeau’s instances, $L=4$)
+$S_{RF}$ $S_{FF}$ $S_{MF}$ $S_{EFF}$ $S_{ERF}$ Name OBJ F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time 30N_4K_A 409.8 376 0.2 0.1 0.4 372 0.4 0.1 0.5 410 0.4 0.1 0.5 3499 2.5 0.1 2.9 12876 3.7 0.5 5.8 30N_4K_B 482.4 179 0.1 0.1 0.3 171 0.1 0.1 0.4 180 0.2 0.2 0.4 1771 1.5 0.1 1.7 6404 2.2 0.3 3.0 30N_4K_C 436.5 603 0.4 1.2 1.9 899 1.0 1.1 2.2 764 1.2 2.1 3.4 5250 4.8 0.9 6.1 17664 5.1 3.1 9.6 30N_5K_A 425.9 525 0.3 0.5 1.0 759 0.7 0.4 1.4 688 0.9 0.5 1.6 4823 3.5 0.5 4.3 13595 4.0 0.8 6.1 30N_5K_B 450.9 327 0.2 0.5 0.8 353 0.3 0.5 0.9 365 0.5 0.5 1.1 2912 2.1 0.3 2.6 8955 2.6 0.9 4.4 30N_5K_C 460.0 466 0.3 0.7 1.1 547 0.5 0.4 1.0 547 0.7 0.9 1.7 3251 2.9 0.4 3.6 11851 4.0 1.5 6.5 40N_4K_A 554.0 759 0.4 1.0 1.7 965 1.4 0.5 2.1 887 1.4 0.7 2.3 7629 10.3 1.5 12.7 27638 8.1 3.0 13.4 40N_4K_B 552.7 998 0.9 2.0 3.3 1375 1.9 1.3 3.5 1273 1.9 1.5 3.6 14541 13.3 3.8 18.0 52456 15.5 5.3 25.6 40N_4K_C 591.6 629 0.3 0.7 1.2 657 1.3 0.7 2.2 689 0.8 0.5 1.5 5536 4.3 1.1 5.9 18807 5.1 1.8 9.0 40N_5K_A 515.0 799 0.5 0.5 1.2 1276 2.0 1.5 3.8 1055 1.5 1.3 3.0 8613 6.9 0.6 8.0 25871 8.1 3.1 13.5 40N_5K_B 556.8 1023 0.6 1.4 2.3 1648 2.2 0.9 3.3 1370 1.7 1.3 3.3 8184 7.2 2.2 10.0 27449 8.6 8.9 19.9 40N_5K_C 565.7 626 0.3 0.3 0.9 832 1.2 0.3 1.6 760 1.1 0.2 1.5 6608 5.2 1.0 6.7 22313 6.8 1.9 10.7 44N_4K_A 580.5 1242 0.9 1.1 2.2 2015 2.6 0.8 3.6 1584 2.6 0.9 3.8 16698 12.8 2.6 16.5 65445 19.4 7.0 32.1 44N_4K_B 597.3 940 0.5 0.5 1.3 1700 1.8 0.2 2.3 1312 1.9 0.7 2.8 11465 9.7 1.5 12.3 32793 9.5 5.2 17.6 44N_4K_C 588.9 1596 1.0 4.4 5.8 2653 3.2 2.0 5.4 1978 3.5 2.6 6.4 15657 16.5 3.0 20.8 61526 19.3 11.9 36.7 44N_5K_A 582.1 1162 0.7 0.5 1.6 1771 2.0 0.7 2.9 1460 2.0 0.4 2.6 10167 7.5 0.8 8.9 32297 10.2 1.9 14.9 44N_5K_B 624.9 945 0.6 1.0 1.9 1426 1.8 0.7 2.7 1143 1.8 1.5 3.5 10108 7.6 0.6 8.9 39795 12.8 4.0 20.2 44N_5K_C 592.1 1056 0.6 2.0 2.8 1313 2.4 1.0 3.7 1237 1.7 1.2 3.1 10907 7.5 1.8 10.1 38224 11.8 5.1 20.2 50N_4K_A 665.0 1570 1.1 2.3 3.9 2997 4.7 2.8 7.9 2273 4.7 6.9 12.1 12507 11.7 3.8 16.5 41095 13.4 15.7 34.6 50N_4K_B 685.3 1322 0.7 0.7 1.9 2094 3.2 0.4 3.9 1732 3.3 0.4 4.0 18800 13.8 0.8 15.9 59798 29.0 6.0 41.5 50N_4K_C 641.7 1146 0.8 2.2 3.4 1888 2.8 1.3 4.3 1525 2.8 1.3 4.5 14709 10.8 2.2 14.0 46530 18.4 17.5 41.0 50N_5K_A 645.5 2245 2.3 10.2 13.2 5473 7.2 4.5 12.1 3267 9.1 5.3 14.9 32723 26.7 5.8 34.5 124110 52.0 87.2 152.0 50N_5K_B 666.2 1393 0.8 3.6 5.0 2623 4.5 2.1 7.0 1896 3.9 1.2 5.5 21982 16.3 3.3 21.5 71971 25.8 12.9 46.2 50N_5K_C 667.8 1163 0.8 4.7 5.9 1438 3.1 6.8 10.3 1347 2.6 4.4 7.4 18851 13.5 4.6 19.5 77905 24.8 38.4 72.9 60N_4K_A 814.5 1830 1.2 7.2 8.9 3313 4.8 4.5 9.7 2514 4.8 4.7 9.9 30150 23.0 4.0 29.3 88197 34.2 23.1 66.1 60N_4K_B 727.3 4533 4.5 21.2 26.9 10314 18.3 16.3 35.5 5961 18.0 15.0 33.7 68432 58.2 10.5 72.8 257386 86.3 142.3 252.1 60N_4K_C 747.1 2588 1.7 16.2 18.7 5205 12.6 4.8 18.2 3409 9.2 5.8 15.6 40287 32.3 6.2 41.3 141157 43.3 58.6 117.5 60N_5K_A 743.2 2618 1.7 6.4 9.1 5543 12.9 0.9 14.5 3726 8.8 2.1 11.4 44007 35.9 4.9 43.8 146211 45.4 33.5 96.0 60N_5K_B 800.8 2260 1.7 2.8 5.4 4585 9.2 3.6 13.4 3103 8.1 1.8 10.5 39063 30.2 3.8 37.0 126061 41.1 62.5 115.6 60N_5K_C 822.7 2125 2.0 6.4 9.1 3647 6.7 1.7 8.9 2633 6.6 2.8 9.8 34999 27.5 5.3 35.4 128702 42.4 21.8 75.8 Avg 1301.5 3.4 4.8 2328.4 2.1 6.3 1702.9 2.3 6.2 17471.0 2.6 18.0 60836.1 19.5 46.0 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; CPU: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
+Avg represents the average metric values calculated across all instances.
+Table 4: Computational performance of PSFF with different fragment sets for DARP-LPT within 30 minutes (Riedler & Raidl’s instances, $L=4$)
+$S_{RF}$ $S_{FF}$ $S_{MF}$ $S_{EFF}$ $S_{ERF}$ Name OBJ F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time 30N_4K_A 392.2 376 0.2 0.5 1.3 386 1.2 1.1 2.4 411 0.7 0.3 1.2 4188 3.6 0.8 5.0 12876 5.1 1.3 12.0 30N_4K_B 459.7 179 0.1 0.1 0.3 174 0.6 0.4 1.4 180 0.2 0.2 0.4 1775 1.4 0.5 2.1 6404 1.8 0.4 3.1 30N_4K_C 419.5 603 0.3 20.9 21.8 2565 11.4 15.8 27.5 777 5.4 4.9 10.4 13862 18.4 11.2 31.3 17664 8.3 58.8 69.8 30N_5K_A 408.5 525 0.3 4.3 4.9 1396 2.7 10.9 13.8 694 2.3 5.9 8.5 9187 9.1 2.5 12.9 13595 4.8 17.0 24.8 30N_5K_B 425.2 327 0.3 1.2 1.7 499 1.0 1.4 2.6 366 1.1 3.6 5.0 4269 7.5 1.9 10.8 8955 2.6 4.4 9.3 30N_5K_C 445.1 466 0.3 1.7 2.3 857 1.6 1.7 3.5 564 2.2 1.7 4.1 5210 9.4 1.8 12.3 11851 4.5 11.2 17.5 40N_4K_A 525.3 759 0.5 3.8 4.8 2277 7.0 3.7 11.1 918 5.8 2.4 8.5 15364 25.8 7.9 36.0 27638 8.8 80.9 95.1 40N_4K_B 522.2 998 0.7 3.8 5.0 1752 8.7 3.1 12.2 1288 6.8 3.3 10.5 25987 29.0 10.3 42.4 52456 18.0 26.4 52.7 40N_4K_C 557.5 629 0.5 7.5 8.6 822 2.1 4.2 6.5 696 1.9 9.0 11.2 6472 7.2 1.6 10.1 18807 6.1 19.9 30.1 40N_5K_A 477.5 799 0.5 1.6 2.7 3252 10.0 1.8 12.3 1069 12.4 2.3 15.0 20806 27.0 2.2 31.8 25871 12.7 19.6 38.1 40N_5K_B 528.9 1023 0.7 12.0 13.2 3705 7.6 23.3 31.5 1397 7.7 17.6 25.7 16061 19.9 13.6 35.8 27449 8.9 79.8 94.7 40N_5K_C 536.8 626 0.3 0.4 1.1 1768 4.0 0.3 4.7 769 3.9 0.4 4.6 12393 14.8 0.8 17.4 22313 6.7 10.6 21.6 44N_4K_A 541.6 1242 0.9 4.5 5.9 6294 14.4 4.2 19.4 1623 13.8 4.3 18.5 50962 54.8 10.6 70.6 65445 29.0 185.5 225.2 44N_4K_B 562.4 940 0.6 8.1 9.4 6362 9.6 10.2 20.5 1341 13.2 11.1 24.7 32137 33.7 6.1 43.0 32793 12.3 155.9 174.3 44N_4K_C 551.6 1596 1.1 32.2 33.9 12758 29.5 19.6 50.3 2125 33.3 25.5 59.4 61590 78.3 43.1 127.7 61526 23.0 247.0 279.7 44N_5K_A 554.9 1162 0.7 13.3 14.7 4183 7.2 15.7 23.4 1491 8.6 8.6 17.7 23984 25.7 5.0 33.4 32297 13.0 142.7 161.3 44N_5K_B 589.3 945 0.6 10.7 11.9 3677 8.7 9.8 19.3 1167 8.4 9.9 18.7 22135 28.0 4.4 34.7 39795 15.4 19.2 40.8 44N_5K_C 568.4 1056 0.7 26.6 27.9 2218 6.6 22.7 29.8 1277 6.3 24.7 31.4 23095 25.7 16.4 44.7 38224 12.3 105.6 125.6 50N_4K_A 628.4 1570 1.3 31.1 33.3 7105 18.3 45.2 64.2 2310 16.3 28.3 45.1 24443 39.2 12.1 54.2 41095 15.3 355.3 378.9 50N_4K_B 639.0 1322 0.8 6.0 7.6 7106 20.3 5.0 26.2 1755 18.2 4.0 22.7 51665 53.1 7.5 65.8 59798 21.5 172.1 203.3 50N_4K_C 606.8 1146 0.7 13.7 15.1 5169 13.5 15.3 29.6 1530 11.3 14.6 26.4 30064 32.4 15.7 51.6 46530 15.3 500.0 524.0 50N_5K_A 601.8 2245 1.7 177.2 179.9 45134 120.3 294.2 418.0 3506 108.9 127.1 236.8 221956 322.3 211.4 553.3 124110 53.7 1723.0 1800.9 50N_5K_B 624.4 1393 1.0 55.3 57.0 11381 29.3 32.0 62.9 1935 29.0 26.7 56.2 81754 99.2 102.5 212.1 71971 26.2 399.7 439.6 50N_5K_C 619.7 1163 0.7 46.2 47.6 2141 11.6 20.5 32.5 1362 11.1 22.4 34.4 33933 35.0 32.6 71.1 77905 32.9 289.3 338.5 60N_4K_A 762.8 1830 1.6 16.3 19.7 10331 35.1 15.1 51.3 2592 33.0 5.9 39.6 93547 133.0 57.8 200.7 88197 29.0 315.9 364.2 60N_4K_B 680.1 4533 5.9 205.6 214.8 103503 273.4 337.3 619.1 6569 273.1 113.5 388.1 553669 797.0 586.0 1432.7 No 60N_4K_C 698.1 2588 2.1 270.2 273.7 32289 91.7 234.5 329.3 3642 88.7 198.2 288.0 212470 303.8 147.1 472.8 No 60N_5K_A 698.8 2618 2.3 37.1 41.2 36234 107.6 78.8 189.2 4011 105.0 27.2 133.1 229204 327.9 244.7 594.6 No 60N_5K_B 748.1 2260 1.7 133.2 136.6 25308 65.7 152.4 220.3 3234 63.8 128.8 193.4 166511 245.4 203.4 467.3 No 60N_5K_C 768.6 2125 1.6 167.4 170.7 14714 51.2 107.4 160.1 2878 52.3 96.6 149.6 151402 222.1 169.3 407.0 128702 44.6 505.3 572.4 Avg 1301.5 43.8 45.6 11845.3 49.6 83.2 1782.6 31.0 63.0 73336.5 64.4 172.8 26/30 solved 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; CPU: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
+Avg represents the average metric values calculated across all solvable instances.
+“No” indicates that no optimal solution is obtained within the 30-minute time limit.
+Table 5: Computational performance of PSFF with different fragment sets for DARP-LPT within 30 minutes (Riedler & Raidl’s instances, $L=6$)
+$S_{RF}$ $S_{FF}$ $S_{MF}$ Name OBJ F Net CPU Time F Net CPU Time F Net CPU Time 30N_4K_A 267.1 146 10.7 1.4 13.3 134 8.1 0.9 9.5 156 8.1 0.7 9.2 30N_4K_B 313.1 72 8.9 1.0 10.4 66 7.2 0.4 7.8 72 7.2 0.4 9.9 30N_4K_C 280.4 146 9.0 3.3 12.8 132 7.3 4.0 12.1 149 7.0 2.1 9.4 30N_5K_A 251.1 194 11.8 2.2 15.7 205 12.4 1.2 14.1 219 12.0 1.6 14.1 30N_5K_B 295.5 150 10.4 1.3 12.2 140 8.2 0.5 9.2 155 7.2 0.8 8.5 30N_5K_C 299.4 180 8.4 1.2 10.3 195 7.9 1.6 9.9 206 7.2 4.8 12.4 40N_4K_A 371.2 259 12.0 2.5 16.0 266 7.6 1.7 10.1 273 11.0 1.5 13.6 40N_4K_B 311.3 441 10.6 9.1 21.7 445 12.0 2.7 15.7 479 17.8 5.2 24.1 40N_4K_C 374.1 197 10.6 0.9 12.8 182 14.3 0.7 15.7 192 8.5 0.9 10.2 40N_5K_A 294.6 271 9.1 2.7 13.1 327 10.0 2.0 13.1 321 8.3 1.5 10.8 40N_5K_B 340.8 327 11.6 2.6 15.8 395 16.8 2.2 19.9 373 7.9 1.9 10.7 40N_5K_C 366.7 229 9.3 6.1 16.6 226 9.5 4.4 14.8 239 7.6 2.5 10.9 44N_4K_A 369.2 361 13.7 4.8 20.4 437 9.6 3.5 14.3 423 11.3 4.9 17.5 44N_4K_B 381.4 383 10.2 10.6 23.0 547 13.2 3.1 17.5 498 15.0 7.0 23.4 44N_4K_C 390.5 503 11.5 20.3 34.0 668 11.6 11.5 24.6 601 12.9 21.3 36.4 44N_5K_A 340.6 533 13.2 4.9 19.9 719 13.1 1.7 16.3 614 12.8 2.7 17.5 44N_5K_B 368.0 326 9.1 6.9 17.4 378 10.5 6.2 18.1 375 10.8 10.4 22.9 44N_5K_C 369.3 334 14.6 9.7 25.6 319 16.0 7.3 24.5 352 11.8 8.1 21.6 50N_4K_A 396.8 475 10.3 13.2 26.0 535 13.5 6.3 21.6 542 23.0 10.9 36.2 50N_4K_B 423.9 539 12.8 11.7 27.5 686 17.9 6.9 26.8 632 21.8 13.3 37.6 50N_4K_C 386.5 519 12.7 9.2 24.3 802 13.5 7.9 23.7 681 19.0 15.4 37.5 50N_5K_A 388.7 962 18.7 28.1 50.6 1928 22.7 36.7 64.4 1283 24.2 31.8 59.9 50N_5K_B 407.2 612 17.5 25.3 45.7 965 17.4 28.7 49.3 769 21.5 48.8 73.2 50N_5K_C 401.2 415 16.1 16.2 34.2 458 12.8 6.1 21.0 473 13.1 9.6 24.7 60N_4K_A 509.6 622 11.2 12.5 27.4 741 18.1 7.6 28.9 728 22.2 8.8 34.7 60N_4K_B 437.0 1502 18.3 88.7 114.9 2657 38.9 58.1 104.5 1879 35.1 26.8 68.5 60N_4K_C 475.4 891 12.9 116.9 135.6 1432 28.1 161.1 194.2 1162 22.8 40.1 67.8 60N_5K_A 476.8 736 10.1 64.9 80.7 1203 21.9 32.5 58.4 985 27.1 92.6 124.5 60N_5K_B 469.5 781 12.5 25.6 43.0 1194 42.8 15.7 63.3 987 29.1 34.6 70.0 60N_5K_C 531.9 535 10.3 12.0 26.3 634 16.4 14.2 33.9 593 20.4 18.2 41.8 Avg 454.7 17.2 31.6 633.9 14.6 31.9 547.0 14.3 32.0 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; CPU: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
+Avg represents the average metric values calculated across all instances.
+Table 6: Computational performance of PSFF with different fragment sets for MDARP-LPT within 30 minutes (Riedler & Raidl’s instances, $L=4$)
+Appendix C Full computational results of FFF for DARP-LPT
+Table 1 describe the computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes.
+$S_{RF}$ $S_{FF}$ Name F Net CPU Time OBJ LB Gap F Net CPU Time OBJ LB Gap a3-18 73 0.3 0.0 0.4 302.5 302.5 0.00% 64 0.1 0.0 0.2 302.5 302.5 0.00% a3-30 93 0.1 1799.9 1800.0 494.8 491.0 0.75% 89 0.1 1799.9 1800.0 494.8 491.1 0.74% a3-36 121 0.1 1800.5 1800.0 574.4 560.6 2.40% 112 0.2 1797.6 1800.0 574.4 565.0 1.63% a4-48 318 0.2 1798.9 1800.0 697.9 686.1 1.69% 312 0.5 1799.4 1800.0 697.9 687.4 1.50% a6-60 601 0.4 1793.4 1800.0 869.0 843.2 2.97% 648 1.4 1795.5 1800.0 864.9 843.4 2.48% a7-70 1036 0.8 1796.3 1800.0 986.3 958.5 2.82% 1273 2.6 1792.3 1800.0 986.6 956.5 3.05% a8-80 1088 0.8 1789.1 1800.0 1064.6 1016.4 4.53% 1473 5.1 1785.7 1800.0 1071.1 1019.3 4.84% a8-96 1465 1.0 1756.7 1800.0 1354.8 1296.8 4.28% 1684 7.3 1759.3 1800.0 1353.6 1302.7 3.76% Avg-A 599.4 1566.9 1575.1 2.43% 706.9 1566.2 1575.0 2.25% b3-36 169 0.1 1799.9 1800.0 576.5 566.2 1.79% 161 0.1 1799.7 1800.0 576.5 567.7 1.53% b4-40 311 0.2 1799.5 1800.0 635.2 630.4 0.76% 203 0.4 1799.1 1800.0 635.2 631.1 0.64% b5-50 518 0.4 1798.6 1800.0 789.1 766.2 2.91% 803 1.2 1797.6 1800.0 789.1 768.3 2.63% b6-60 940 0.6 1793.3 1800.0 931.4 895.5 3.86% 1218 2.6 1792.3 1800.0 931.7 900.7 3.33% b7-70 960 0.7 1796.4 1800.0 948.0 921.6 2.79% 1655 3.2 1789.5 1800.0 950.4 920.7 3.13% b8-80 1195 2.0 1790.3 1800.0 1105.8 1071.8 3.08% 1813 5.1 1788.6 1800.0 1104.8 1072.0 2.97% b8-96 No 6540 33.2 1734.9 1800.0 1322.6 1260.9 4.67% Avg-B 682.2 1796.3 1800.0 2.53% 975.5 1794.5 1800.0 2.37% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
 The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
-Table 4: Computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes (Extended instances, $L=6$)
-Appendix C Full computational results for MDARP-LPT
-Both tables below serve as supplements to Table 7.
-One point worth noting is that while the instances used in both tables are derived from Cordeau, they are not identical since we modify the number of vehicles and the number of customers (see Table 4 for details).
-$S_{RF}$ $S_{FF}$ $S_{PFF}$ F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 45 8.9 0.5 9.4 239.6 239.6 0.00% 40 6.2 0.4 6.6 239.6 239.6 0.00% 35 17.9 0.7 18.7 239.6 239.6 0.00% a3-30 42 7.9 1.4 9.4 357.9 357.9 0.00% 42 9.6 2.1 11.9 357.9 357.9 0.00% 40 10.2 1.2 11.4 357.9 357.9 0.00% a3-36 51 6.5 47.1 53.7 474.8 474.8 0.00% 48 8.0 31.3 39.5 474.8 474.8 0.00% 37 11.1 38.1 49.2 474.8 474.8 0.00% a4-48 182 6.8 217.5 224.7 499.1 499.1 0.00% 166 6.9 179.6 186.8 499.1 499.1 0.00% 124 21.0 131.0 152.2 499.1 499.1 0.00% a6-60 236 7.2 1791.4 1800.0 632.1 624.9 1.14% 210 8.3 1791.2 1800.0 632.1 625.7 1.00% 150 52.3 1557.4 1610.2 632.1 632.1 0.00% a7-70 715 10.3 1788.5 1800.0 646.9 619.6 4.23% 687 13.0 1786.1 1800.0 644.6 622.7 3.40% 364 183.5 1615.6 1800.0 639.9 625.9 2.19% a8-80 613 10.1 1788.7 1800.0 753.6 714.7 5.16% 728 14.2 1784.6 1800.0 747.4 718.0 3.93% 389 192.2 1606.8 1800.0 746.7 720.1 3.56% Avg-A 269.1 8.2 813.9 1.50% 274.4 9.5 806.4 1.19% 162.7 69.7 777.4 0.82% b3-36 203 6.6 148.7 155.5 425.6 425.6 0.00% 227 8.3 479.7 488.2 425.6 425.6 0.00% 133 12.7 181.1 193.8 425.6 425.6 0.00% b4-40 518 9.7 1789.8 1800.0 405.9 402.4 0.85% 502 8.8 156.0 165.1 405.9 405.9 0.00% 260 24.2 132.4 156.8 405.9 405.9 0.00% b5-50 1878 20.9 1777.6 1800.0 497.5 480.1 3.51% 1903 15.6 1783.3 1800.0 494.3 488.4 1.20% 865 44.8 432.3 477.6 494.3 494.3 0.00% b6-60 2537 17.4 1779.9 1800.0 603.1 574.4 4.77% 2334 23.5 1774.6 1800.0 597.8 589.9 1.33% 969 127.6 1230.6 1359.3 596.6 596.5 0.00% b7-70 2310 19.5 1777.3 1800.0 610.8 574.4 5.96% 3287 46.4 1750.2 1800.0 625.7 581.8 7.02% 1145 273.7 1525.2 1800.0 609.0 585.3 3.89% Avg-B 1489.2 14.8 1471.1 3.02% 1650.6 20.5 1210.7 1.91% 674.4 96.6 797.5 0.78% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
-The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
-Table 1: Computational performance of FFF with different fragment sets for MDARP-LPT within 30 minutes (basic instances: $|V|=\lceil n/4\rceil$, $L=4$)
-$S_{RF}$ $S_{FF}$ $S_{PFF}$ F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 37 7.5 0.7 8.2 229.2 229.2 0.00% 33 15.1 1.3 16.5 229.2 229.2 0.00% 29 8.0 0.3 8.4 229.2 229.2 0.00% a3-30 36 6.9 0.6 7.6 325.8 325.8 0.00% 36 7.5 0.6 8.2 325.8 325.8 0.00% 34 7.0 2.0 9.2 325.8 325.8 0.00% a3-36 47 9.1 5.9 15.1 397.2 397.2 0.00% 44 6.3 6.5 12.9 397.2 397.2 0.00% 34 15.2 4.0 19.3 397.2 397.2 0.00% a4-48 116 8.3 40.0 48.6 450.8 450.8 0.01% 109 10.0 27.5 37.9 450.8 450.8 0.00% 92 21.1 23.7 45.1 450.8 450.8 0.00% a6-60 175 9.3 766.0 775.9 556.1 556.1 0.00% 155 7.5 490.1 498.0 556.1 556.1 0.00% 113 40.3 396.6 437.4 556.1 556.1 0.00% a7-70 507 12.0 1786.8 1800.0 552.5 549.8 0.48% 482 14.2 861.7 876.9 552.5 552.5 0.00% 250 139.2 504.8 645.1 552.5 552.5 0.00% a8-80 495 16.9 1781.1 1800.0 663.9 641.2 3.41% 586 15.9 1783.0 1800.0 670.7 642.9 4.15% 318 137.9 1661.4 1800.0 661.8 647.6 2.14% Aver-A 201.9 10.0 636.5 0.56% 206.4 10.9 464.3 0.59% 124.3 52.7 423.5 0.31% b3-36 146 8.3 72.9 81.9 389.3 389.3 0.00% 148 9.5 36.9 46.9 389.3 389.3 0.00% 100 15.2 27.6 42.9 389.3 389.3 0.00% b4-40 370 8.7 45.6 54.7 355.7 355.7 0.00% 314 8.6 23.5 32.3 355.7 355.7 0.00% 178 21.0 16.3 37.5 355.7 355.7 0.00% b5-50 1649 16.8 1781.2 1800.0 436.3 427.0 2.12% 1635 20.4 659.2 680.9 436.3 436.3 0.00% 753 45.4 298.9 344.9 436.3 436.3 0.00% b6-60 2269 31.3 1763.9 1800.0 527.9 512.3 2.95% 2041 28.8 1768.7 1800.0 527.7 522.6 0.98% 816 139.1 397.4 537.9 527.7 527.7 0.00% b7-70 1541 18.6 1778.4 1800.0 530.5 509.6 3.93% 2000 32.5 1764.2 1800.0 528.4 519.1 1.76% 747 233.9 1565.0 1800.0 528.4 518.8 1.81% Aver-B 1195.0 16.7 1107.3 1.80% 1227.6 20.0 872.0 0.55% 518.8 90.9 552.6 0.36% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver.
-The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
-Table 2: Computational performance of FFF with different fragment sets for MDARP-LPT within 30 minutes (basic instances: $|V|=\lceil n/3\rceil$, $L=4$)
+Avg-B does not involve the information of “b8-96” for both $S_{RF}$ and $S_{FF}$ since the former cannot lead to any feasible solution within the 30 minutes.
+Table 1: Computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes (Cordeau’s instances, $L=4$)